A Complete Bibliography of Publications in Theoria

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

28 September 2023 Version 1.18

Title word cross-reference

$150.00 [1533]. $50.00 [1558]. ^{,} [553]. ^{1} [574]. _{ω1ω0} [563]. E = K [1498]. H [296]. K [748]. T_{1} [686].

-Clauses [1502].

. [170].

0 [1175, 1192, 1168, 1172, 1166, 1167, 1181, 1182]. 0-19-823754-5 [1181]. 0-19-924382-4 [1166]. 0-19-926975-0 [1176]. 0-19-927002-3 [1175]. 0-19-927040-6 [1176]. 0-521-63056-8 [1182]. 0-521-81047-7 [1172]. 0-521-81800-1 [1192]. 0-674-00616-X [1173]. 0-674-01043-4 [1168]. 0-7486-1638-1 [1167]. 0-7486-1684-5 [1183].

1 [1184, 1605, 1192, 1167, 1191]. 1-84064-964-x [1184]. 1-904987-00-1 [1191]. 10 [1262]. 12th [1735]. 12th- [1735]. 14th [1735]. 14th-Century [1735]. 18th [904]. 1927 [972]. 1938 [120]. 1939 [87]. 1940 [113]. 1945 [184, 193]. 1946 [194]. 1947 [216]. 1948 [234, 235]. 1949 [275]. 1950 [276, 282]. 1952 [315, 320]. 1953 [327]. 1958 [423]. 1959 [465]. 1984 [855]. 1985 [846]. 1986 [860]. 1993 [966, 965]. 1st [170, 184, 193, 194, 215, 216, 234, 235, 275, 276, 282, 315, 320, 327, 423, 465, 513, 525].

2003 [1174]. 23rd [120]. 25.00 [1650]. 2nd [1174].

3 [1175]. 300-Jahresfeler [46]. 300-year-old [46]. 30th [184, 194, 215, 234, 276, 315, 327]. 31st [170, 193, 216, 235, 275, 282, 320, 423, 513, 525].

4 [1168, 1166]. 407 [1715]. 40th [1741]. 40th-Year [1741].

5 [1181].

6/9 [87].

7 [1172]. 7/7 [87].

8 [1182].

9 [1624, 1551]. 978 [1624, 1551, 1605]. 978-0-19-061214-6 [1590]. 978-0-19-881282-1 [1605]. 978-0-19-966877-9 [1533]. 978-0-262-03560-6 [1558]. 978-1-138-64378-9 [1551]. 978-3-11-067740-9 [1624]. 978-3-89665-694-0 [1532].

A. [646, 41, 410]. Åke [380]. Åqvist [1574, 634]. Åsa [1135]. Abailard [782]. Abduction [1210, 1018]. Abductive [978]. Abfassungszeit [68]. Ability [1075, 854]. Abnormalism [637]. Aboensis [841]. Aboutness [1633]. Aboutness-Fixing [1633]. above [175]. Absences [1403]. Absolutist [1709]. Abstract [1307, 1372, 1642, 749]. Abstraction [1838, 1837, 609, 942]. Abstractionism [1840]. Abstractionists [1833]. absurdities [802]. absurdity [1788]. Abu [1731]. Abus [418]. abuse [418]. Academia [1532]. Academiae [841]. Academic [1572, 1566, 1672]. Acceptance [1010, 1845]. Access [1431, 1526, 1458]. accidents [535]. According [293, 527, 175, 944]. Account [1709, 1222, 1602, 1487, 1633, 1412, 1744, 1353, 1420, 1845]. Accountability [1094]. Accuracy [1622]. Acknowledgments [1672]. Acquaintance [906]. Acquisition [1641]. Act [626, 913, 883]. Act-utilitarian [883]. Acta [841]. Action [1348, 1702, 1231, 1375, 436, 296, 639, 1853, 915, 546, 855, 1797]. Action-Individuation [1348]. actions [852, 817]. active [1824]. Acts [1478, 433, 1017]. Actual [1088]. Actualism [732, 845]. Addendum [672, 1218]. Addition [969]. additional [699]. Additivity [1195]. Adequacy [1033]. Adequate [1420]. Adequately [1353]. Adoption [1226]. adults [260]. Advantage [1024]. Aesthetic [693, 692, 695, 696, 697]. Aesthetic/Non [695]. Aesthetic/Non-aesthetic [695]. Aesthetics [389]. af [4]. Affairs [1519]. affectifs [311]. Afraid [1064]. African [1482, 1174, 1766]. after [1240, 1540, 1573, 959]. Again [1215, 858]. Against [1082, 1072, 1202, 1093, 1497, 1332, 1813, 1770, 1201, 1503]. Age [1363, 1461, 1631, 1355, 1742]. Ageing [1624]. Agency [1664, 1673, 1285, 1348, 1806]. agent [987]. agent-relative [987]. Agential [1375]. agentives [880]. agents [807]. Ages [1049]. Aggregating [1055]. agreement [954]. Agreements [1585]. agriculture [199]. Aharonov [1787]. Aigle [333]. Aims [1037, 960]. Akbarian [1736]. aktuella [26]. al [1731, 1731]. al-Balkhi [1731]. al-Malahimi [1731]. al-Qasim [1731]. Alain [1184]. Alanen [1168]. Albert [1508, 3]. Alcohol [418]. alcoolique [418]. Alethic [1817]. Alex [1590]. Alexander [219, 1183]. Alf [263]. algebra [394]. algebraic [950]. Algebras [1007]. Algorithm [599, 577]. Alive [1180]. Allan [1441, 1209]. alleged [240]. Allen [196]. Alternate [636]. Alternative [1637, 1837, 654]. alternatives [842, 639]. Always [1478, 995, 919, 610]. Am [1275, 120, 335]. Ambiguity [522, 1679, 662, 768, 434]. Amoralism [897, 896]. Amsterdam [217]. Analogy [1211, 1018]. Analyse [142, 147]. analyser [248]. Analyses [553, 1853]. Analysis [1700, 1294, 91, 586, 1473, 298, 830, 1851, 488, 1788, 740, 787, 716, 633, 343, 167, 303, 142, 147]. Analytic [271, 1264, 1287, 1436, 824, 353]. analytic-synthetic [353]. Analytic/synthetic [824]. analytical [15, 294, 272, 278]. Analyticity [397, 743, 1815]. analytischen [294, 1325]. analytiska [15]. analyze [248]. Anatoli [1734]. Anatomy [1154]. anchored [1845]. Ancient [1628, 1532]. Anders [175, 757, 758, 113, 175]. Anderson [1518]. Andersonian [925]. Andries [775, 774]. Anil [1212]. animae [252]. Animal [1310]. Animals [1482]. anknytning [26]. Anmälningar [5, 9, 14, 20, 27, 30]. Anniversary [1741]. Announcement [735]. Anonymous [1527]. anorexia [1072]. Ansätze [141, 519]. Anschaulichkeit [386]. Anschauung [57]. Answer [1733, 1752, 1444, 175, 280]. Answerable [1299]. Answers [1716]. Antecedents [859]. Anthony [1406]. Anthropological [1763]. Anti [1341, 1147, 1361, 1485, 1805, 759]. anti-exceptionalism [1805]. anti-moralistic [759]. Anti-Psychologism [1485]. Anti-Realism [1361, 1147]. Anti-Scepticism [1341]. Antinomien [326]. antinomies [326, 641]. Antiquity [1731]. Antiquiw [1192]. Antireductionism [1277]. Antonio [361]. Antwort [175, 280]. Anything [1630]. apodictic [332]. Applicability [608]. Application [428]. applications [902, 478, 394]. Applied [1052, 1333, 916]. Apply [1584]. Applying [1097, 566]. apprehension [671]. Approach [1250, 1701, 1259, 1505, 1677, 629, 1261, 1805, 870, 997, 653, 770, 1814, 1757, 825, 1756, 834, 700, 1435]. Approaches [141, 519]. April [184, 194, 215, 234, 120]. Apriorism [74]. Apriorisme [74].

Aprioristic [1540]. Apriority [1530]. Apropos [1042]. Arabic [1732]. arbete [15]. archetypical [800]. Archimedean [1195, 1092]. area [2, 7]. Arendt [1461]. Argument [1569, 1368, 1187, 1082, 1330, 1332, 1477, 1245, 1514, 1698, 1779, 1147, 1676, 585, 398, 1093, 1291, 1381, 1398, 1211, 1813, 1847, 1086, 986]. Argumentation [1735]. Arguments [1736, 1731, 1476, 1611, 1498, 530, 1160]. Aristotelianism [1834]. Aristotle [666, 932, 543, 1234]. Arithmetic [1004]. Arithmetical [752, 1548]. arithmetization [1005]. Armchair [1575]. Armstrong [1139]. Arne [422, 473]. Art [1595, 1689, 1635, 47, 928, 255, 697, 690, 612, 1479]. Artefacts [1354]. Arthur [645, 674, 586]. article [1573]. Articles [1672, 290]. artifacts [928]. Artistic [1688, 1685]. arts [694, 1183]. Artworks [689, 1292, 690]. artworld [690]. ascending [706]. Aseity [414]. aspect [378]. Aspects [285, 219, 252, 767]. Assassination [1355]. Assault [1560]. Asserted [1647]. Assertion [1714, 755, 212]. assessment [1848]. associated [801]. Assuming [963]. Assumption [1030]. assumptions [470, 301]. Asveld [1311]. Asymmetric [1355]. Asymmetry [1349]. Atheism [1163]. atomic [746]. atomique [242]. Attack [1503]. Attitude [1700, 204, 696, 301]. Attitudes [1700, 1607, 1648]. Attribution [1761]. Aufbau [167]. Auffassung [302]. August [216, 235, 170]. Augustin [1532]. Ausdruck [280]. Auseinandersetzung [388, 422, 1325]. Auselm [325]. Austin [627]. Author [1496]. Authorial [1618]. Autobiographical [939, 938, 822]. Autoepistemology [1080]. Autonomy [1797, 1798, 1713, 1796]. Averroistes [372, 372]. Avoid [1055, 1471]. Axel [1326, 55, 1323, 87, 1489]. axiom [756]. axiomatic [770]. axiomatizability [680]. axiomatizable [748]. Axiomatization [634]. Axioms [508, 156]. Ayer [582, 410]. Ayten [1461].

B [1184, 1287, 970, 718]. Bachelors [1396]. Back [1742, 1764]. Background [323, 1050]. Bad [1840, 1424]. Badness [1104, 1102, 1103]. Bailhache [860]. Balkhi [1731]. banana [977]. Bänden [871]. Barhebraeus [1737]. Based [1474, 1459, 1659, 1389, 311, 583, 872, 418, 998, 575, 1726]. Bases [1073]. Basic [1357, 157, 4, 151, 409, 473, 26, 137]. Basing [1556]. basis [1803, 843, 254]. Baumann [1520]. Bayes [1486]. Bayesian [1008, 1176]. Bayesianism [1008]. Be [1378, 1418, 1332, 1312, 1609, 1503, 1696, 322, 851, 1496, 1661, 1843, 409, 473, 1118]. Bearers [1502]. Beatrice [1232]. Bedeutung [68, 141]. Been [1724]. Before [552]. beginning [515]. Begins [1774]. Begreppet [2, 7]. Begriff [237, 279, 111, 268, 221, 1325, 444]. Begriffe [444]. Begriffsanalyse [91]. Begriffsbildung [62]. Begründung [137]. Behavior [304, 211]. behaviour [1848]. beim [211]. being [982, 909, 970, 1340]. Beings [1429]. Beitrag [161, 69]. Beiträge [290]. Belief [1529, 1613, 1426, 1505, 1039, 1339, 1321, 1493, 1074, 1466, 1094, 1306, 1000, 737, 927, 998]. Beliefs [1495, 1663, 857, 1338, 1581, 1600, 1646, 1849]. Believe [1434, 601]. believing [857]. Belnap [808]. Belysning [4].

Bemærkninger [4]. Bemerkung [280]. Bemerkungen [41, 280, 279, 268, 57]. Ben [1199]. Ben-Menahem [1199]. benefit [969]. Benevolence [884]. Bentham [1271]. Bergstrom [464, 686, 714]. Berkeley [501, 329, 427, 541, 433, 377, 398, 349]. Berlin [1624]. Bermúdez [1060]. Berrocal [1488]. Bertrand [324]. Besires [1230]. Besprechung [175]. best [1824, 995]. betingelserna [15]. betrachtet [306]. Betrachtung [68]. Betrachtungen [46]. betreffend [252]. Better [1751]. Between [1624, 1429, 967, 1595, 1763, 564, 1585, 736, 395, 998, 1287]. Bewegung [422]. Beyond [867, 1407]. Béziau [1149, 1154]. Bias [1629]. Biased [617]. Biases [1493]. Bibliographical [118, 124, 130, 135, 140, 154, 145, 150, 170, 160, 165, 188, 184, 193, 194, 202, 207, 215, 216, 224, 234, 235, 270, 275, 276, 282, 309, 315, 320, 327, 338, 339, 358]. Bibliographie [263, 49]. Bibliography [369, 379, 396, 406, 423, 440, 465, 500, 513, 525, 833, 1137, 1217, 1216, 646, 1061, 1077, 1113, 1156, 1185, 1246, 1272, 1308, 1342, 1364, 1399, 1427, 1454, 1481, 1510, 1509, 1096, 263, 49, 263]. Biblographical [177]. bifurcationism [1745]. binary [911, 479]. Bio [110]. Bio-Psychology [110]. Biology [810, 632]. Biosemantics [1538, 1536]. Biscuits [1671]. Bivalence [976, 1016]. Black [1616]. Blackburn [1082]. Blackmail [1355]. blameworthiness [956]. blank [867]. Blendungsphänomene [247]. Blind [1136]. Blowing [1566]. blue [977]. Bob [1429]. Bobro [1198]. Body [1378, 1301, 1315, 40, 41, 660]. body-mind [660]. Boèce [372]. Boethius [1048, 372]. Bohm [1787]. Bolzano [271]. Book [105, 385, 570, 600, 905, 818, 1533, 1624, 307, 1590, 1532, 1551, 1605, 1650, 1558, 101, 175, 64]. Books [171, 178, 195, 206, 310, 328, 340, 359, 365, 375, 424, 441, 451, 458, 466, 475, 485, 493, 514, 526, 532, 557, 591, 625, 670, 681, 688, 698]. Boolean [394]. Bootstrap [1150]. Bound [1632]. Bounded [1004]. Bourdeau [1490]. Bourdieu [1301, 1315]. Bourne [1239]. Bovens [1176]. Brain [1116, 1632, 537, 16, 120]. Brain-Zap [1116]. branching [719]. Brandom [1299, 1287]. Brazilian [1560]. breadth [237]. break [794]. Breite [237]. Brian [1254]. bridge [1827, 1808]. Bridges [1191]. British [904]. Broad [146, 348]. Broadbent [1590]. Broadie [1183]. Bronaugh [687]. Broncano [1488]. Broncano-Berrocal [1488]. Broome [1421]. Broome’s [969]. Brueckner [1406, 1401]. Brussels [1193]. Buches [175]. Bullshit [1719, 1620]. Burden [1514, 1011]. Burden-Shifting [1514]. Byzantium [1735].

C [1207, 416, 146, 288]. C. [146]. Calculi [607]. Calculus [508, 552, 453, 233, 413, 394]. called [220, 198, 318]. Callergård [1225]. Cambridge [1223, 1090, 1248, 1173, 1192, 1172, 1182, 1558]. Campbell [1408]. Can [1378, 851, 1418, 1116, 1619, 766, 1455, 1599, 1587, 1503, 348, 1471, 1738, 1555, 1353, 1118, 1190, 1661, 248]. Cancellation [577, 599]. Cannibals [1027]. Cannot [1420, 1190, 409, 473]. canonical [880]. Cans [431]. Cappelen [1533]. caractères [277]. Careful [1378]. Carens [1462].

Caring [1281]. Carnap [167, 422, 422]. Carolyn [1483]. Cartesian [722]. Case [1668, 1703, 1797, 1269, 1585, 1715, 1786, 819]. Cases [1269]. Cassier [179]. Cassirer [313]. Cassirers [34]. Castañeda [463]. Casting [1824]. categorial [787]. Categoricity [1691, 1692, 413]. Categories [628, 555]. Cats [1396]. Causal [1777, 581, 1313, 312, 1307, 1438, 838, 1744, 535, 1274, 1658, 772, 1743]. Causal-Informational [1744]. causalité [61]. Causality [61]. Causation [1673, 503, 1222, 1696, 1806]. Causation^{≫} [324]. Causes [636, 742, 814, 764, 477]. Ceilings [1669]. Centennial [1469]. central [1814]. Centred [1647, 1780]. Century [1742, 1735, 904, 1732]. chain [706]. chains [1743]. Challenge [1760, 1374, 1789]. Chalmers [1704]. Change [146, 1844, 801, 765, 888, 597, 927]. changes [999]. Changing [1253, 784, 956, 1362, 1790]. characteristics [277]. characterizability [563]. characterization [387]. characterological [4]. Charge [1479]. Charity [1278, 886, 1789]. Charles [357]. Charlotte [1226]. charts [260]. Cheating [1579]. Check [1760]. Cheryl [1650]. chez [350]. Children [1044]. Choice [231, 988, 1757, 1756, 948, 1442]. choices [817]. Christia [1200]. Christian [1382]. Christine [1285]. Christoph [136]. Chronological [573]. Chronology [82]. Cicero [1668]. circle [722, 19]. Claim [1724]. Clarendon [1166, 1181]. Clash [1670]. Class [283, 1797]. classes [461]. Classical [658, 1192, 1191, 684, 736]. Classificatory [1289]. Clauses [1502]. Cleaning [1721]. Climate [1421]. clock [492]. Close [1701]. Closed [1269]. Closure [1087, 1816, 1497]. Codes [126, 132]. codices [803]. Coetzee [1174]. Cogitare [1628]. Cogito [335, 540]. Cognitive [1835, 1094]. Cognitivism [1204, 1529]. Cognitivists [1027]. Coherence [1201, 1486, 73, 1001, 1051]. collapse [762]. Collapsing [1544]. Collection [1042]. Collective [1702, 229, 852]. Collectivism [190]. College [1191]. Collingwood [841]. Collisions [1222]. Colocationist [1716]. color [977]. Commandment [121]. Commandments [1390]. comme [75]. Comment [685, 714, 248, 1296]. Commentaire [333]. Commentary [249, 1408, 333]. commentators [951]. Comments [685, 463, 362, 1020, 832, 280, 473, 4]. commersio [252]. Commitments [1322]. Common [1498, 58, 1818, 970]. Common-sense [58]. Communicating [1570]. Communication [1638, 1563, 1527, 1647, 1646]. Communists [1027]. Communitarian [1300, 1804]. communities [1765]. Companion [1223, 1184, 1248, 1182]. Company [1840]. Comparability [1475]. Comparative [1290]. comparing [747]. comparisons [754]. Competing [1010]. Complete [1101, 709, 724]. Completeness [664, 523, 587, 675, 619, 738, 949]. Complexes [1404]. Complexity [1004, 1688, 1713]. compliance [816]. Compositionality [1158]. Computability [1490]. Computation [1607, 1648]. Computer [1363]. computers [770]. Computing [1191]. conceive [863]. Concept [1209, 1220, 198, 1641, 442, 1168, 351, 381, 1713, 56, 997, 265, 813, 46, 63, 695, 175, 2, 7, 245, 4, 142, 147, 656, 671, 34, 111, 272, 278, 221, 261, 579, 444, 1791]. Conception [1714, 110, 1362, 1492, 793]. Conceptions [1329, 1050, 1631].

Concepts [1252, 1640, 1638, 1244, 1670, 367, 538, 1835, 533, 1570, 1639, 426, 321, 1832, 662, 834, 505, 497, 26]. Conceptual [1531, 1851, 91]. Conceptualism [1641]. Conceptualization [62]. Concern [1101]. Concerning [346, 323, 1248, 426, 233, 245, 252, 1745, 479]. Concerns [1187, 1105]. Conciliationism [1760]. Conclusion [1475, 1471, 1171, 1088]. Conclusive [1081]. Condensed [1169]. Condillac [350]. Condition [1516]. Conditional [1853, 1067, 682, 633]. Conditionals [1623, 747, 664]. Conditions [1316, 15, 772, 911, 477]. Condorcet [1754]. Confessions [1388]. Confidence [1517]. Confirmation [547]. Confirmed [1388]. Conflict [1678, 1715, 736, 1355]. conflicting [750]. Conflicts [1322]. Confrontation [422]. Confronts [1331]. Confusion [1351, 1110]. Congrés [208, 213]. Congress [208, 217, 213]. Conjunct [1269]. Connection [494, 872, 26]. Connections [1450]. connective [247]. Conquer [1653]. conscience [248, 333]. Conscious [1306]. Consciousnes [119]. Consciousness [119, 248, 716, 1166, 1060]. Consequence [293, 1847, 955, 1791]. Consequences [1586, 240, 15]. Consequential [1492]. Consequentialism [1036, 1053, 1197, 985]. Consequentialist [1053, 773]. consequents [859]. Conservatism [1504]. Conservativeness [844]. consideration [68]. Considerations [1268]. Consistency [1367]. Consonance [1189]. Constancy [1675, 1030]. Constants [1433]. Constituents [778, 1404]. Constitution [1285, 1675, 1642]. Constitutive [1408]. Constraints [1644]. Constructing [914]. construction [167]. constructions [777]. Constructive [1014]. Constructivism [1331]. Constructivity [1490]. construites [241]. Consumer [1797]. containing [701]. Contemporary [1199, 1461, 249]. Content [1538, 1596, 1284, 352, 1679, 970]. Context [1253, 1147, 1643]. Contexts [1638, 1644, 1623, 737, 762]. Contextualism [1520]. Contingency [1612]. Contingent [1543, 1695, 1518, 1750, 1041]. Contractarianism [1024]. Contraction [801]. Contractualism [1053]. Contractualist [1054, 1142]. Contradiction [1109]. Contrariety [1154]. Contrasting [782, 1792]. Contrastive [1658, 1705]. Contribution [161, 90, 1718, 263, 69, 544]. Contributions [1395, 1470, 1654, 1301, 1315]. Control [1664, 849]. Controversy [392, 346, 197]. Convention [341]. Conventionalism [876]. conventions [790]. converging [713]. Conversation [1692]. conversations [1782]. Conviction [162]. Cooperation [1022, 1803]. Copenhagen [120]. Copernican [1251]. Corazón [1685]. Cornell [1419, 1353, 1420]. corporis [252, 252]. correct [1843]. Corrected [599]. Correction [1827]. Corrections [590]. correctness [1784]. Correspondence [73]. corresponding [731]. Corruption [1563]. could [1811]. Counter [502]. Counter-Induction [502]. Counterexample [1549, 1500, 1582]. Counterfactual [1696, 664]. Counterfactuals [531, 1283]. Counterpart [1187, 1696, 1802]. Craig [1239]. Credence [1613]. creeds [366]. crime [311]. Criminalizing [1586]. criminogène [291]. criminogenic [291]. Crisis [69]. Criteria [596, 533]. Critical [1136, 68, 1547, 91, 973, 298, 1009, 691, 1256, 739, 272, 278, 348, 974, 167, 823, 318, 473, 17, 1057]. criticism [17]. Critics [898]. Critique [1499, 1701, 41, 50, 816, 695, 1813, 305]. critiques [318]. Crito [954]. Cronus [789]. cross [1807]. cross-linguistic [1807]. Crossroads [1595]. Cryptography [1004]. CS [634]. Cultural [1495]. Culture [1632, 304]. Culture-Bound [1632]. Cultures [1279]. current [26, 25]. Customs [126, 132]. cyclic [989]. Cyclical [988].

D [1343, 146, 288]. D. [1223, 1069, 1139]. Dacia [1048, 372]. Dacie [372]. Dag [1021, 1017]. Dancy [1036, 1034, 1175, 1228, 1503]. dangereux [277, 331]. dangerous [277, 331]. Danielle [1131]. Danish [6, 33, 4]. Dans [860, 74, 329, 331]. Dark [1445, 772]. Darwin [357, 1763, 740]. Dasgupta [1009]. Data [119]. datum [268]. David [1213, 1256, 1107, 1166, 1191]. Davidson [1430, 967, 966, 1181]. Davis [1184]. d’axiomes [156]. deal [1821]. Debate [1299, 1723, 1230]. Debates [1008]. December [193, 275, 282, 320, 423, 465, 513, 525]. Deception [1525]. Decidability [594, 593, 664]. Decision [1549, 1115, 498]. declaration [2, 7]. Deducibility [1446]. deducing [1812]. Deduction [1070, 840, 1018, 1726]. Deductive [1497]. deeds [588]. Deep [865]. defeasability [804]. Defective [1306]. Defence [1529, 1575, 1368, 1319, 1400, 1656, 1419, 1851, 1853, 812]. defences [782]. Defense [1101, 1570, 1520]. Deference [1643, 1172]. Defining [1740, 1765, 745]. Definite [306, 585, 1196, 768]. definiten [306]. Definition [271, 1732, 494, 1159, 1479, 330, 337, 1617, 555, 1778, 272, 278, 478]. Definitions [1208, 529, 298, 351, 665, 554]. deflationism [1789]. Degrees [1318]. Deliberation [1526, 944]. Deliberative [1664, 1526]. délinquance [418]. delinquency [418]. delusion [1852]. Demand [1599, 1600]. Demarcation [1768]. demise [979]. Democracy [1526]. Democratic [1508]. Demon [1190, 958]. Demotivating [1512]. Demystified [1298]. Denken [141, 211, 334, 181]. Denkens [141]. Denkprozessen [247]. Dennett [1847, 977]. Denotation [912, 430, 778]. Denotationless [620]. Denoting [731]. Denying [1030]. Deontic [1043, 1140, 902, 453, 925, 955, 498, 1433, 1670, 1180, 650, 1580, 1518, 648, 1417]. deontically [648]. deontics [491]. Deontological [1504]. deontologies [987]. Department [853, 854]. Dependence [414, 1800, 1261]. Derek [1470, 1474, 1475, 1471, 1473, 1470, 1472]. derogatory [1850]. Descarte [68, 335, 46]. Descartes [46, 68, 1011, 335, 1168, 1340, 1190]. Description [640, 912]. Descriptions [306, 585, 1280, 1466, 1196, 805, 768]. descriptive [1784, 272, 278]. descriptives [241]. Descriptivist [1368]. deserves [438]. Design [1205]. designation [804]. Desire [1569, 1034]. Desires [1385, 893, 892]. Deskriptionen [306]. Determination [1798, 1679]. Determining [1101]. Determinism [1656, 348, 934, 885, 700]. Detour [1605]. Deutschland [69]. Development [1393, 697, 120]. developments [1573]. Deviant [1706]. d’Ezéchiel [333]. Diagonalizable [1007]. Diakonistyrelses [113]. Dialectic [1734]. Dialetheism [1439, 1825]. Dialogue [1208, 1186, 1365, 1383, 1053, 668, 1059]. Dialogues [82]. Dickie [1633]. dicto} [314]. did [1425]. Differ [1678]. difference [815]. differences [814]. Different [1213, 1474, 650]. Different-Number-Based [1474]. Differs [1719]. difficulty [902]. Digby [427]. Dilemma [1502, 371, 1541]. Dilemmas [1355, 865, 883]. dilemme [371]. Diodorus [789]. Direct [1134]. Directly [602]. Disability [1545, 1765]. Disabiliy [1207]. Disagreeing [1517, 1570]. Disagreement [1736, 1779, 1709, 1384, 1645, 1678, 1528, 1531, 1822, 1722, 1224]. disagreements [1789]. disant [318]. Discipline [1068]. Disciplines [1740]. Discours [46]. Discourse [1432, 343, 463, 464, 46]. Discrimination [1314, 1659, 1629]. Discussion [273, 344, 354, 374, 384, 445, 499, 524, 569, 576, 606, 631, 463, 464, 415, 624, 410, 314, 395, 421, 411, 15, 268, 388, 348, 598, 15, 21, 28, 31]. Discussions [42, 52, 59, 70, 108, 112, 116, 122, 138, 143, 148, 152, 158, 182, 191, 200, 205, 222, 269, 390, 400, 405, 439, 450, 457, 507, 519, 548, 549, 556, 562, 580, 589, 613, 618, 103, 511, 536, 98, 133, 128, 163, 168]. Disease [1192]. Disguised [1626]. disjunctive [825]. Disjunctivism [1458]. Diskussion [279, 268]. diskussionen [15]. Diskussionsinlagg [10, 15, 21, 28, 31]. disorders [311]. Dispensable [1330]. Displacement [1507]. disposition [426]. Dispositional [1035, 1849, 834]. Dispositions [1037, 1282]. Disproved [1724]. Disputations [1051]. Dispute [1039]. disputes [1818]. Dissent [1766]. Dissertation [1455]. Dissertations [780]. dissolved [869]. Distance [1468, 1488, 505, 1488]. distances [998, 673]. Distant [1410]. Distinct [1570]. distinction [395, 353]. distinctions [762]. Distributive [1803]. disvalue [1100]. divalent [261]. Diverse [1277]. Diversity [284, 1465]. Divide [1653]. Divided [1664]. Dividing [1655, 1653]. division [393]. Do [1495, 1043, 1609, 250, 409, 1835, 1678, 119, 1190, 422, 473]. Docent [15]. Docenterna [15]. Doctors [1192]. Doctrine [284, 543, 294, 115]. Does [1386, 1360, 1584, 471, 1103, 1304, 40, 41, 949]. Doesn’t [1541]. Dogmas [1768]. Doing [1394, 1443, 325, 1287, 854]. Domain [1448, 1762]. Don [6]. Donald [1181, 967, 966]. Don’t [601, 1536, 1686]. Dordrecht [855]. Double [1760, 372]. Double-Check [1760]. double-vérité [372]. Doubt [1341, 335]. Dougherty [1573]. Douglas [1214, 898]. Doxastic [1348, 1492, 1373, 1297]. Dream [1190, 733]. Dreams [403, 539]. Dress [1412]. Dretske [1515]. Driven [1418]. Droit [226, 228]. Drugs [1214]. Dualism [1093]. Dubucs [1490]. Duhem [1042]. Dummett [1147, 769, 756]. Duncan [1206]. d’une [311, 263]. Duns [1707]. Durations [673]. Duties [1300, 244, 792]. Duty [121, 296]. dyadic [955]. Dzur [1508].

each [970]. Eagle [333]. Eagleton [1286]. Early [1223, 1323, 697, 1200]. Earth [1394, 1460, 1409, 1744, 1381]. earthquake [239]. Ease [902]. Economics [1184]. economy [13]. écrits [263]. ed [1223, 1199, 1174, 1183, 1182]. Edinburgh [1183, 1167]. Edited [1255, 1508, 1241, 871, 1248, 1310, 1489, 1311, 1490, 1309]. edition [1256]. Editor [937, 1123, 1334, 1002, 1063, 918, 995, 1336, 1029, 1335, 1046, 1296].

Editorial [1028, 1084, 1078, 1091, 1114, 1122, 1138, 1146, 1208, 1219, 1683, 918, 645, 1022, 1038, 1052, 1072, 1062, 1068, 1097, 1106, 1151, 1157, 1169, 1186, 1177, 1162, 1202, 1227, 1242, 1249, 1235, 1264, 1271, 1279, 1257, 1302, 1312, 1288, 1294, 1346, 1319, 1333, 1327, 1351, 1358, 1365, 1371, 1400, 1383, 1390, 1377, 1407, 1416, 1423, 1431, 1443, 1448, 1437, 1455, 1484, 1463, 1469, 1476, 1491, 1501, 1496, 1522, 1045, 1194, 1511]. eds [1200, 1213, 1226, 1533, 1624, 1184, 1174, 898, 1193]. Educate [1084]. Educated [1558]. Edward [1184]. effect [651, 1787]. Effective [1397]. Efraim [125]. Egalitarian [1101, 1105]. Egalitarianism [1101]. Egalitarianism. [1241]. ego [350]. egoism [843, 1812]. Eijk [1192]. Eileen [1200]. Einige [386, 252]. einiger [286]. Eino [1296, 1376, 401]. Einstein [41, 189]. Einsteins [189, 3]. Einsteinschen [41]. Einteilung [393]. Ekelöf [899]. Element [256]. Elemental [326]. Elementare [326]. Elementary [614, 760, 720, 577, 272, 278]. Elements [1737, 769, 773]. Elena [1325]. Elgar [1184]. Eliaeson [1489]. Eliminativism [1439]. Elin [1193]. Elizabeth [1419]. Elliott [1198]. else [1703]. elsewhere [786]. Eltroxin [1797]. embeddings [720]. Emergency [1631]. Emergentism [1543]. Emerson [357]. Emotion [542, 1483]. emotional [311]. emotions [1194]. Emotive [942, 693]. Empirical [447, 1530, 1807, 422]. Empirically [1712]. Empiricism [1673, 1376, 257, 422, 1405, 19, 23, 1212]. Empiricist [1222, 1554]. empirische [422]. empirismen [19]. Empirismens [23]. Empirismus [422]. Empty [1437]. enactment [841]. Encores [1412]. endeavour [1851]. endpoints [15]. Endurance [1159]. Endurantism [1259]. Energy [1396]. Enflo [1442]. Engineering [1165]. English [1219, 1351]. Enhancement [1798]. enligt [175]. Entail [1360, 1304]. Entailment [1269, 233]. Entities [1307, 1839, 749]. entrenchment [927]. Environment [110, 740, 916]. Enzyklopädie [871]. Epicurus [934]. Epigenetic [1632]. Epistemic [1433, 1538, 1047, 1331, 1019, 1515, 1410, 1480, 1273, 1270, 1720, 1715, 663, 550, 1206, 1520]. Epistemolgy [76]. Epistemological [1767, 1458, 1087]. Epistemology [84, 722, 92, 1705, 1165, 56, 811, 862, 945, 946, 812, 1176]. Equal [1634, 1795, 678]. equal-interval [678]. Equality [1085, 1526, 1803, 1785, 1241]. equanimity [18]. equivalence [708, 723]. Equivalent [592]. Era [1227, 68]. Erfahren [211]. ergo [540, 335]. Eric [1363]. Erik [833, 1201, 1532, 1293]. Erkenntnisphilosophie [56]. Erkenntnistheoretische [189]. Ernst [34, 34, 179]. Errata [486, 856, 486]. Erratum [355, 449, 558, 1263, 1345, 435, 1345]. Error [1701, 1717, 1323]. Erwachsenen [260]. Escape [1699]. Eschatological [236]. Eschatologische [236]. Essay [1225, 101, 1415, 1248, 1429, 1605, 623, 677, 348]. Essays [1285, 1241, 1309, 1226, 1388, 1406, 1042]. Esse [412, 415, 377]. Essence [666, 1747, 721]. essentially [418]. essentiellement [418]. est [377]. Establishment [107, 263]. état [331]. Etchemendy [1791]. Ethical [1586, 1484, 1681, 292, 481, 1603, 1812]. Ethics [1462, 65, 1213, 1474, 1255, 93, 1052, 1333, 1443, 1584, 1766, 78, 1311, 1382, 392, 1604, 1507, 1551, 1009, 1172, 1167, 1340, 1179, 1012, 1193, 1421, 1143, 795, 121, 151, 916, 137, 1175, 1482, 1058, 1286, 1357, 1369]. Ethik [137, 121, 151]. étude [504, 418]. EU [1586, 1585]. Eubulides [326, 326]. EUR39.00 [1532]. Evaluating [1635]. Evaluation [1283]. evaluations [989]. event [478]. Events [1696, 560]. ever [471]. every [1811]. everyday [940, 952]. Everything [1359]. Evidence [1562, 1663, 1613, 1389, 1498, 1715, 386, 838, 332]. Evidence-Based [1389]. Evident [602]. Evidential [1163]. Evidentialism [1767, 1221]. Evidentiary [947]. Evidenz [386]. Evolution [1495, 1763, 1343]. Evolutionary [945, 551, 811, 944, 812]. exact [1661]. Examination [531, 392, 452]. Example [1154]. Examples [1169]. exceptionalism [1805]. Exceptions [1032]. Excerpt [1692]. Excess [1650]. Exchange [967]. excluded [1019]. Excluding [538]. Exclusion [1305, 1314]. Excuses [528]. Exegetical [1250, 1243]. Exist [1835, 119, 250]. Existe [438]. Existe-t-il [438]. Existence [571, 1457, 279, 621, 1447, 232, 265, 459, 666]. Existentials [788]. existents [788]. Existenzbegriffes [265]. existiert^{≪} [280]. exists [280, 471]. expansion [947]. Expansionism [1594]. Expectation [913]. Expectations [1260]. Expected [817, 1086]. experéences [418]. Experience [446, 1126, 1299, 1826, 336, 421, 1212]. experiencing [211]. Experimental [1407, 1367]. experiments [418]. Expertise [1575, 1484, 1592, 1094]. Explaining [1445, 814, 836]. Explanation [1658, 1313, 1305, 551, 230, 1179, 476, 1274, 1824, 742, 635, 699, 477, 661]. Explanations [1231, 1600, 1164, 798]. Explanatory [1260, 1305, 1477, 1633, 1274, 1374]. Explicit [521]. Exploring [1460, 1105]. Expression [232, 858, 280]. Expressions [605, 628]. expressive [620]. Expressivism [1601]. Expressivist [1056]. extended [744]. Extension [1347, 748]. Extensional [708, 723]. Extensionalist [1388]. Extensions [614, 1089, 720, 760]. External [1692]. Externalism [1394, 1603, 1087, 1523]. Externalist [1393]. Externalized [1585]. Extra [1676]. Extra-Value-of-Knowledge [1676]. Ezekiel [333].

F [1182]. Facere [1628]. facie [649]. facteur [291]. Factive [1724]. factor [291]. Factors [1486]. Facts [299, 560, 731, 746]. factual [970]. Failed [1728]. Failure [1568, 1466]. Fain [415]. Fairness [1545, 754]. False [1663, 1466, 1646, 409, 473]. Falsemakers [1494]. Family [283, 1612, 597, 579]. Fanaticsm [895]. färdert [422]. Father [1580]. Faultless [1709, 1645]. Favour [1498]. Feit [1823]. Feldtheorie [141]. Feminist [1226]. Fiction [1628]. Fictionalism [388]. Fictions [1292, 388, 777]. field [25, 46, 199, 141]. Figurative [1065]. Fiktionalismus [388]. Fiktionen [388]. Fill [1352]. Filosofi [113, 33]. filosofiskt [13]. filtration [655]. Fine [1183]. Finetti [924]. Finite [678]. finitely [748]. Finks [1677]. Finland [853, 854]. First [565, 1742, 609, 1762, 101, 941, 873, 413, 429]. First-Order [609, 1762, 413, 429]. Fission [1541]. Fitting [1700]. Five [739, 1257]. fixed [1762]. Fixing [1633, 1750]. Flat [1773]. Flexibility [1593]. Flikschuh [1090]. Floridi [1356]. Flow [1561]. ^{≪} [307]. Focus [1199]. Fodor [1158]. Folke [1224]. Following [1268, 967]. follows [756]. fonction [331]. fondée [311, 418]. Force [1581]. Forced [1507]. Forces [1595]. foreign [24]. Forgiveness [1143]. Förlag [113]. Form [1425, 551, 1153, 1325, 1211, 880, 349]. Formal [1400, 674, 1754, 700]. Formalized [407]. Formation [1731, 13]. formed [641, 803]. Forms [313]. Formula [324]. formulas [419]. Forty [1746]. Fotion [898]. Foundation [1835]. Foundations [972, 663, 460, 467]. Four [1047, 642, 594]. Fourteen [482]. Frage [102]. Fragmentalism [1561]. Frame [1433]. frames [709, 724]. främmande [24]. Francaise [213]. Frank [1650]. fransk [33]. FRE [1715]. free [912, 794, 700]. Freedom [1442, 1463, 1572, 1801, 322]. freethinking [727]. Frege [1325, 1457, 662, 367, 1714, 1362, 1752, 1243, 1485, 1605, 428, 1057]. Fregean [1833]. Freiheit [322]. Freitag [1468]. French [213, 311, 208, 83, 321, 61, 218, 226, 74, 329, 51, 89, 64, 48, 33, 75, 228, 277, 291, 318, 331, 418, 292, 248, 371, 504, 372, 96, 156, 361, 438, 333, 213, 255, 350, 263]. Freud [302]. Freudian [1331]. Freuds [302]. Friendly [1261]. Friendship [1513]. frivolous [1770]. Frontiers [1207]. Full [1073]. Function [447, 1153, 203, 363, 16, 331]. Functional [1189, 935, 246, 531, 394]. Functions [1347, 801, 662, 863]. Fundamental [1767, 1666, 1164, 662, 17, 1164, 386]. Fundamentalsätzen [386]. funktion [16]. Funktionszusammenhang [246]. Further [1573]. Fusions [1707]. Future [1024, 1339, 1239, 1338, 741, 761, 719]. Fuzziness [836].

G [841]. G. [45]. G.3 [1089]. Galen [237, 1814]. Galois [872]. Game [1434, 1048, 1221, 908]. game-theoretical [908]. Gap [1260]. Gaps [1081, 652]. Garson [1233]. Gary [1430]. Gaynesford [1125]. Gebot [121]. Gedanken [286]. Gegenwartsphilosophie [345]. Gehirnentwicklung [120]. Geist [252]. Gendler [1533]. General [318, 1724, 411, 436, 729, 682, 259]. générale [318]. generalisation [656]. Generality [459, 1556, 1516, 1596, 650]. Generalization [530, 708, 723]. Generalized [565, 1799]. Generation [1009]. Generation-Relative [1009]. Generations [1024]. Generic [720, 1792]. Genetic [110, 1798]. Genius [1448]. Genuine [1458, 1645, 516]. Geometrical [1011]. Geometrie [57]. Geometry [1764, 829, 57]. Georg [136]. Gerhard [1435]. German [236, 237, 49, 40, 62, 326, 322, 519, 581, 265, 46, 68, 114, 81, 63, 386, 41, 175, 286, 189, 305, 91, 162, 36, 50, 306, 131, 161, 180, 246, 247, 290, 107, 142, 147, 293, 34, 111, 146, 280, 268, 294, 335, 173, 252, 102, 121, 151, 296, 334, 388, 57, 181, 115, 185, 221, 393, 256, 260, 422, 141, 211, 302, 261, 264, 383, 120, 69, 137, 444, 295, 56, 136]. Germany [69]. Gerson [1234]. Gesamtheit [393]. Geschichte [256]. Gesichtspunkte [252]. Gestalt [421]. Gesundheit^{≪} [237]. Get [831, 1359]. Getting [1806]. Ghazali [1733]. Gibbard [1209, 1441]. Gibson [1182]. Give [518]. Given [1786]. Gjelsvik [1127]. Gladness [300]. Glare [247]. Glüer [1129]. God [305, 468, 1422]. Gödel [1684]. Gödel’s [1557]. Gogh [47]. Golden [357, 1742]. Goldman [1676]. Good [1428, 1098, 392, 1424, 1801, 111, 221]. goodness [487]. Göran [1506]. Gottesbeweises [305]. Governance [1585]. Gradation [1525]. Grammar [994, 729, 659]. grand [64]. gränser [23]. Gratitude [1672]. Gravity [1712]. Great [1044, 239, 64]. Greek [1532]. Grelling [326, 326]. Grelling-Nelson [326, 326]. grip [1806]. Gross [1355]. Grösse [40, 41]. Ground [1038, 1835, 1818]. Grounded [1666]. Grounding [1568, 1716, 1703]. grounds [1745]. group [1845]. Grundbegreb [4]. grundbegrepp [26]. Grundmotiv [151]. grundståndpunkt [17]. Grundvoraussetzungen [137]. Gruyter [1624]. Guarded [1400]. Guest [1123, 1063, 1029, 1046]. Guidance [1453, 1681, 1398]. Guiding [1775]. guilt [90]. Gündogdu [1461]. Gunky [1597]. Gunnar [88, 293, 15, 293]. Gupta [1212]. Gurmund [368]. Guten [111, 221].

H [1462, 1174, 1069]. Habermas [1763, 959]. Hacking [1173]. Haecceitism [1802]. Hagerström [175, 175, 87, 286, 1489, 1323, 175, 1326]. Hägerströms [286]. haine [51]. Hale [1429]. Hallden [943, 307, 944, 949]. Halldén-completeness [949]. Haltaufderheide [1624]. Hampshire [511]. Handbook [1533]. Handlung [296]. hands [975]. Hannah [1461]. Hanno [1558]. Hans [386, 386, 155, 272, 278]. Hansen [1215]. Hansson [1193, 1296]. happens [588]. Happiness [1244, 1340, 1428]. Harald [176]. Hardback [1650, 1533, 1183, 1176]. hardcover [1173]. Hare [1058, 890, 392, 986, 898]. Harm [1659, 1823]. Harm-Based [1659]. Harmony [1582, 1785, 1778]. Harms [1055]. Harry [72]. Hart [1343]. Hartmann [1176]. Hartnack [280]. Harvard [1173, 1168]. Haskell [415]. Haslanger [1226]. hatred [51]. Hauptsatz [587]. Hauptströmungen [345]. having [970]. Hawthorne [1533, 1269]. hbk [1558]. Headaches [1055]. Health [1059, 1192, 237]. Heart [1685]. Heavyweight [1269]. Hedenius [799]. Hedonism [1366]. Heikki [841]. Helsingfors [326]. Helsinki [326]. Hemmung [247]. Herder [114]. Herman [1533]. Herrlin [55]. Hesperus [819]. heterologischen [37]. heute [422]. Heytesbury [1316]. Heyting [595]. high [839]. Higher [1195]. Hilary [1199, 1394, 1395, 1391, 1392, 1393, 1396]. Hinge [1705]. Hintikka [717, 1456]. Hirsch [1789]. Histoire [74, 89, 61]. Historical [1433, 36, 1688, 1490, 477, 476, 742, 661, 635, 699, 764, 581, 1173]. Historische [581]. History [74, 77, 1062, 1258, 1327, 1045, 1296, 1376, 313, 69, 61, 89, 256, 841]. Hjärnfunktion [16]. Hobbes [430]. Hoist [1237]. Holism [1711]. Holtug [1241, 1344]. hook [826]. hopes [885]. Horseshoe [826]. Horsthemke [1482]. hos [26]. hostile [916]. hours [199]. Hovemann [1624]. Howard [1303]. Howick [1389]. Human [1256, 1698, 1248, 1232, 443, 120, 1532]. Humaniora [841]. humanism [378]. Hume [1256, 462, 443, 506]. Humean [1222]. Husak [1214]. Husserl [332, 818]. hypertheories [999].

Hypotheses [1397]. Hypothesis [1075].

Ian [1508, 1173]. Ibn [1731]. idea [578, 853]. Ideal [1723, 332]. Idealism [1350, 1413, 243, 91, 398, 1401, 501, 33]. Idealisme [33]. Idealismus [91]. idealistic [151]. idealistischen [151]. Idealizations [1424]. Ideas [1437, 490, 621, 797, 904, 115]. Ideenlehre [115]. Identicals [1446]. Identifying [1385, 517]. Identities [1658, 1354, 1707]. Identity [284, 1259, 571, 1473, 1577, 1780]. Ideology [1415]. If [1630, 1696, 1103, 1703, 986, 180, 1103]. if-then-else [1703]. Ifs [431]. Ignorance [1202, 1321]. Ignoring [1394]. ihre [393]. II [147, 132, 1290, 147, 710, 725]. IIIe [213]. illumination [4]. Illusion [398, 1717, 119, 8]. im [69]. Images [1687, 534, 490]. Imagination [1576, 1780]. Immateriality [1698]. Immediate [582, 734]. Immigration [1462]. immortality [487]. immutable [956]. Impact [1557, 793]. Impending [1631]. Imperative [872, 172, 663, 121]. imperative-epistemic [663]. imperativen [121]. Imperatives [667, 437, 568, 567, 472, 886]. imperfect [807]. Imperiled [1776]. Impersonal [1099]. Implementations [1277]. Implicata [1720]. Implication [407, 210, 322, 233, 879, 373, 675, 293]. Implicational [508]. Implications [1640, 1603, 616]. Implicit [1629]. implied [322]. Implikation [293]. impliziert [322]. import [866]. importance [1823, 1473]. important [378]. Imprecise [1549, 1475]. Imprecision [1474]. impression [645]. imprimés [263]. improper [805, 946]. improved [432]. Improving [1581]. Inadequacy [1170]. Inappropriate [1600]. Incarceration [1508]. Include [1672]. Incoherent [1712]. Incommensurability [1514]. incomplete [701, 751]. Incompleteness [665, 702, 767]. Inconsistant [38]. Inconsistent [1597]. Independence [1261, 642, 684]. independent [970]. Indeterminacy [1568, 1386, 873]. Indeterminism [348]. Indeterminist [652]. Index [1217, 1510]. Indexical [1688, 805]. Indexicals [1643]. Indicative [1623]. Indices [907]. Indirect [1440, 671]. Indispensability [1418, 1330]. Indispensability-Driven [1418]. Individual [1702, 1300, 254]. Individualist [1411]. Individuals [852]. Individuation [1348, 1049]. Induction [502, 317, 460, 467, 106]. Inductive [960]. Inequality [1104, 1102, 1103, 1100, 26]. Infallibilist [1708]. Infelicity [1708]. Inference [312, 1757, 1756, 567, 1196, 1824, 978, 417]. Inferences [1011]. inferential [1792]. Inferentialism [1852]. infinitary [738]. influence [289]. Informal [1178, 1825]. Information [1521, 1546, 1552, 1559, 1565, 1571, 1578, 1583, 1591, 1598, 1608, 1615, 1625, 1636, 1651, 1660, 1680, 1693, 1710, 1727, 1739, 1753, 1769, 1781, 1794, 1809, 1820, 1828, 1842, 1346, 946, 352, 1356]. Informational [1744]. Ingemar [799]. Ingmar [893, 974, 846]. Inherently [1231]. inhibition [247]. Injustice [1702]. inner [211]. inneren [211]. inotn [25]. Inquiry [121, 1369]. Instructions [1542]. instrumentalism [781]. Integrated [1684]. Intellect [1011]. Intellectual [1667]. Intellectualism [1324]. Intellectuals [1151]. Intelligent [1205]. intension [847]. intensional [658, 907]. Intensionality [1305]. Intent [1618, 2, 7]. Intention [849, 790]. Intentional [1512]. Intentionality [1299, 300, 818, 935]. Intentions [1646]. Intercultural [1383]. Interdisciplinary [1028]. Interest [1472, 233]. Interests [1344, 1715]. Interference [1026]. Interferers [1677]. Internal [716, 1691, 1692]. Internalism [1603, 1034, 1270, 1570]. International [217, 208]. interpersonal [754, 969]. Interpretability [1006]. Interpretation [360, 407, 1670, 286, 1513, 1405, 1238, 694, 1306, 1772, 253, 1001, 419]. Interpretational [1838]. interpretations [776, 815]. interrogative [663]. Intertemporal [1322]. interval [678]. Interview [966, 965, 1320, 1158, 1228, 1209, 1069, 1139]. Intramural [1299]. Intrinsic [881]. Introduction [1741, 1638, 1683, 995, 1107, 1535, 1627, 1771, 1392, 1756, 1167, 1542, 983, 1551]. Introspection [1434, 1189, 942]. Intuition [1524, 317, 57, 1]. intuitionen [1]. Intuitionism [921, 769]. Intuitionistic [1070, 752, 736]. Intuitions [1460, 1092, 1381, 894, 1558]. invalidity [978]. Invariance [1376, 1203]. Invariantism [1589]. inversion [1726]. investigation [272, 278]. investigations [326]. Invitation [1080]. IQ [289]. Irrational [1495]. Irrelevant [1467]. Is/Ought [462]. ISBN [1533, 1624, 1184, 1590, 1532, 1183, 1173, 1551, 1605, 1175, 1192, 1168, 1176, 1172, 1166, 1167, 1181, 1191, 1182, 1558]. Islam [1736]. Islamic [1731]. isn’t [995]. isolation [740]. Isomorphisms [1007]. Israeli [1797]. Issue [1521, 1546, 1552, 1559, 1565, 1571, 1578, 1583, 1591, 1598, 1608, 1615, 1625, 1636, 1651, 1660, 1680, 1693, 1710, 1727, 1739, 1753, 1769, 1781, 1794, 1809, 1820, 1828, 1842]. Issues [1158]. iterated [649]. iteration [968]. ix [1176].

J [1318, 1174, 1192]. J. [1154, 854]. J.-Y. [1154]. Jaakko [1456]. Jackson [783, 977]. Jacob [1734]. Jaederholm [45]. Jäger [1338]. Jahre [69, 268]. Jahre^{≪} [279]. Jahresfeler [46]. James [1233, 64]. January [215, 513, 525, 184, 194, 234, 275, 276, 315, 327, 423, 465]. Jaques [1490]. Jérôme [333]. Jeremy [1389]. Jerome [333]. Jerry [1158]. Jesper [1255]. Jevons [444]. Jevonsschen [444]. Jochen [1624, 1184]. Joe [1309]. Johanna [1624]. Johansson [595, 1506]. John [1533, 1447, 1421, 1184, 969, 1133, 1605]. Johnston [1619]. joint [1845]. Jonathan [1369, 1247, 1228, 1175]. Jönsson [1339]. Jordan [1303]. Joschka [1624]. José [1060]. Joseph [1462]. Journal [1390]. Journals [1672]. JP’ [644]. Jr [1182]. Juan [6]. Judgement [853, 1425, 351, 1117, 979]. Judgements [1678]. judging [612]. Judgment [1130, 162, 294]. Judgments [509, 172, 1558, 691]. July [282, 320]. June [276, 315, 327, 972]. Juokko [1261]. Jürgen [871, 1763]. Juridical [334]. juridique [248]. juridische [334]. Jurisprudence [1254]. Just [1437, 1681, 851]. Justice [1115, 1207, 1669, 1803, 259, 815, 773, 1344]. Justification [1010, 1504, 1845, 1201]. Justified [1306, 851]. Justus [280, 280].

Kai [1482]. Kaila [1296, 1376, 401]. Kalam [1730]. Kant [1090, 1763, 305, 759, 1130, 294, 1717, 115, 185, 793, 823, 1232]. Kantian [773]. Kantianism [1132]. Kants [115, 305, 294]. karakterologisk [4]. Karin [1442]. Karl [1250]. Kasper [1241]. Kathrin [1129]. Katrin [1090]. Kausalsätze [581]. Kelsen [272, 278]. Kemp [1430]. Kevin [1551]. Kicking [1180]. Kierkegaard [6, 6, 239]. Kill [1055, 1405]. Kind [1396]. Kinds [1313, 1338, 1094, 1396]. King [1191]. Know [1075, 521, 1738, 1324, 1359, 409, 473]. Know-How [1075, 1324]. Knowability [1332, 1247, 1309]. Knowable [1367, 1359]. Knower [1433]. Knowing [1394, 958, 1444, 1304, 550]. Knowledge [316, 1634, 1701, 1332, 582, 571, 1360, 1665, 1676, 1321, 1316, 1530, 1269, 1093, 1111, 1276, 1576, 1353, 1420, 1134, 1375, 1724, 443, 527, 870, 940, 945, 952, 934, 734, 909, 810, 410]. Known [1269]. Kollektivs [615]. konnektive [247]. können^{≫} [322]. Konrad [279, 930]. konsekvensbegreppets [15]. Konsequenz [293]. Kopenhagen [120]. Körpers [40, 41]. Korsgaard [1285]. Kotarbiński [516, 545]. Kripke [1742, 1368, 887, 1612, 1747, 709, 724, 1752, 1447, 868]. Kritik [41, 305, 17]. kritische [68]. kritischer [91]. Kritisk [17]. kunskapsläran [25]. kunskapsteoretiska [24]. Kvanvig [1247]. Kyrkans [113].

L [1355, 563]. Labour [1702]. l’action [860]. Ladder [1555]. Lagerborg [425]. Lang [1193]. Langage [38]. Language [1351, 50, 1158, 428, 991, 561, 411, 436, 729, 993, 448, 955, 213, 818]. Langue [213]. Lars [765, 1506, 368, 464]. Larsson [155]. Last [1411, 921, 197, 268]. Late [1403, 1731]. Later [1746, 579]. Latin [1351]. Laureate [1474, 1394, 1267, 1395, 1475, 1265, 1391, 1535, 1392, 1471, 1473, 1393, 1470, 1472, 1396]. Law [1250, 1569, 683, 230, 455, 254, 1172, 528, 765, 861]. Laws [1775, 1554, 174, 584, 794, 332, 647, 535]. leading [919]. Learning [1754, 1062, 1662, 1338, 25, 211]. lecture [967, 120]. Lectures [220, 1447]. Legacy [1749]. Legal [1081, 199, 248, 272, 278, 836, 303]. Legalization [1214]. Lehre [294]. Lehrer [766, 1010]. Leibniz [197, 240, 323, 1667, 414, 1198]. Leiter [1254]. length [505]. Lennart [1059, 1574]. Leod [775, 774]. Leon [455, 39]. Lernen [211]. Leśniewskian [643]. less [1164]. Let [1760]. l‘établissement [263]. l‘état [277]. Letter [1296]. Letters [1336]. L‘étude [331]. letzten [279, 268]. Levels [847, 650]. Levi [1115]. Lewis [233, 1107, 1283, 1380]. Lex [1248]. Lexical [1593]. Liar [1237]. Libertarianism [348]. Library [846]. Libya [1585]. Lichtenberg [136, 136]. lies [1846]. Life [239, 1244, 1326, 1436, 1428, 885]. life-hopes [885]. Light [190, 1115, 68]. Likelihood [1338]. Likhet [26]. Liljeqvist [125]. Lilli [1168]. Limitations [266, 311]. limites [311]. Limits [1585, 1324, 443, 23, 896]. Lindahl [765]. Lindström [1755, 1756, 1295]. Line [1655]. linear [654]. Lines [1653]. Linguistic [1351, 1423, 306, 1695, 336, 1324, 1851, 1848, 1807]. links [1743]. Linnebo [1840]. Lippert [1241]. Lippert-Rasmussen [1241]. Literal [1065].

Literary [1292, 1713, 1685, 820]. Literature [615]. Litigation [1797, 1770]. live [809]. Lives [1616, 1718, 1101]. livre [64]. Lloyd [1234]. Loaded [1444]. Loader [1508]. local [1003]. Locke [1248, 1447, 527, 101, 238, 1411, 284, 797, 578, 621, 622, 1276, 623]. Lockean [904]. Locutionary [627]. Logic [469, 1433, 1464, 565, 1266, 1047, 1043, 1233, 627, 1180, 1186, 1580, 1714, 1689, 1183, 614, 1485, 1717, 1601, 572, 608, 962, 573, 1261, 574, 575, 568, 567, 609, 456, 1654, 1417, 801, 519, 1827, 663, 1805, 577, 776, 653, 999, 63, 1018, 619, 583, 1140, 902, 960, 366, 657, 701, 453, 655, 709, 724, 752, 755, 650, 550, 713, 872, 347, 683, 912, 983, 760, 656, 661, 925, 102, 994, 635, 699, 1844, 1808, 587, 736, 786, 633, 647, 738, 955, 702, 710, 725, 373, 648, 261, 498, 161, 1191, 1343, 1261, 765, 787]. logic-as-models [1805]. Logical [1021, 1805, 481, 1425, 1033, 1703, 1153, 1690, 1376, 186, 257, 1581, 436, 743, 233, 705, 1819, 419, 349, 382, 167, 303, 1051, 444, 19]. logically [970]. Logico [1250]. Logico-Semantical [1250]. Logics [1612, 630, 1367, 1493, 634, 575, 595, 747, 658, 706, 950, 684, 707, 664, 544, 594, 654, 1762, 751]. Logik [63, 102, 161, 261]. logische [167]. logischen [1325, 444]. logiska [19]. London [1174, 1191]. Longuenesse [1232]. Loon [1445]. Loose [538]. Losing [1038]. lot [1824]. Lotte [1311]. Lower [1260, 394]. Luc [1176]. Luciano [1356]. Luck [1468, 1410, 1488, 1206]. Luis [1060]. lumière [68]. Lund [846]. Lurz [1310]. Lying [1679]. lyricism [388]. Lyrik [388].

M [1058, 1285, 1139]. M. [263]. MA [1173, 1558]. Månillusionens [8]. Mac [775, 774]. Macbeth [1131]. MacFarlane [1133]. Macnish [1551]. Macpherson [1411]. made [539]. Magnitudes [1764]. maintenance [15]. Make [1434, 1044]. Make-Believe [1434]. Makers [1744]. Makes [1659]. Making [1341, 809, 915]. Makinson [1191]. maladie [47]. Malahimi [1731]. Male [1448]. Malfatti [1614]. Malfunction [1329]. Malpass [1813]. mammal [1847]. Man [297, 422]. Mandeville [166]. Manfred [926]. manipulation [1806]. Manuscript [1293]. many [395]. Mapping [1683]. Marc [930, 1198, 279]. Marc-Wogau [930, 279]. Marciano [1184]. Margaret [1240]. Marneffe [1214]. Martha [1207]. Martin [1428]. Marxism [236, 877]. Marxismus [236]. Masks [1677]. Mass [1508]. Material [210]. Mathematical [668, 1530, 1111, 1654, 1685]. Mathematics [1687, 1178, 972]. matrices [763]. matrix [944]. Matter [1200, 1386, 1541, 1103, 932, 949, 1616]. Matters [1226, 1709, 1421]. Mauss [1301, 1315]. maximal [915]. Maximilian [1125]. maximizing [1086]. May [170, 322, 216, 235]. McDougall [262]. McDowell [1299, 1130, 1132, 1405]. McGee [1500]. Me [1760, 1455]. Mean [1707, 543]. Meaning [1732, 1209, 1537, 1595, 811, 862, 1067, 494, 1850, 616, 1605, 736, 1718, 518, 1065, 818, 399, 402, 346, 68, 693, 827, 597, 873, 1844, 1016, 141, 363, 106, 1806, 1430, 1441, 1436]. meanings [408]. Measurement [677, 1376, 510, 679, 678, 680]. Measures [1588, 1442]. Measuring [969, 1100, 1103]. mechanics [382]. Mechanism [1367]. mechanisms [947].

Meddelanden [22, 29]. Media [1525]. medical [1765]. Medicine [1590, 1192, 1389]. Medieval [1047, 1580, 1050, 1729]. Meet [1073]. meets [1805]. Meinong [1325]. Meixner [1317]. Mellor [1069]. Membership [1804]. Memoriam [1574, 1295, 774, 799, 757, 1303, 1755, 1456]. memory [674]. Menahem [1199]. Menschenleiden [120]. Mental [433, 119, 1412, 1724, 745]. Mentalism [1512]. Mentalist [1767]. Mercer [1200]. Mere [1249]. Merely [1818]. Mereological [1594, 1838, 1707]. mérite [438]. Messaging [22, 29]. Meta [481, 1603]. Meta-ethical [481, 1603]. Metalinguistic [1822]. Metaphern [388]. Metaphor [1067, 1066]. Metaphorology [1576]. metaphors [388]. Metaphysical [1432, 1833, 1734]. Metaphysics [1550, 1069, 1139, 414, 443, 727, 728, 1016, 991, 1200]. Method [36, 209, 127, 655, 46]. Methode [36, 46]. Methodological [1302]. Methodology [1655, 1533]. Metric [572]. Metz [1436]. Meurling [72]. Michael [769, 1355, 973]. Michel [1490]. Middle [1049, 538, 1019]. Migrants [1461]. Migration [1702, 1585]. Mike [1428]. Mill [638]. Millians [1502]. Millikan [1535, 1537]. Mimamsa [1670]. Mimesis [697]. Mimic [1495]. Mimickers [1677]. Mind [1664, 1200, 1378, 1158, 1326, 1405, 818, 245, 584, 660, 885, 943, 1168]. Mindful [1094]. Minds [316, 1394, 1134, 245, 734, 1310]. Mindus [1489, 1326]. Miners [1487]. Minimal [1519, 429]. Minimalism [1537, 1834]. Minorities [1288]. miracle [945]. Miracles [1236]. Misak [1650]. Mises [615]. Misleading [1663]. Mistakes [1511, 741]. Misuses [1501]. Misusing [605]. Mittelstrass [871]. Mixed [1380]. Modal [1368, 1187, 1701, 1082, 1253, 583, 931, 630, 1177, 1050, 1291, 608, 575, 654, 609, 1284, 456, 1066, 747, 776, 619, 902, 739, 657, 655, 656, 594, 762, 866, 702, 710, 725, 751, 767, 1233]. Modalities [1838, 1518, 796, 649, 387, 1317]. Modality [1538, 454, 1429, 314, 789]. Model [1655, 523, 1654, 656, 1791]. Model-Completeness [523]. model-theoretic [1791]. Modellings [927]. Models [447, 1723, 1805, 668, 998, 420, 429]. Moderate [1300]. Modern [1200, 1223, 329, 1050, 1489, 1355, 56, 1090, 241, 267, 255]. moderne [329, 255]. modernen [56]. Modified [1366]. Moi [350]. moments [515]. Monde [242]. Monism [846]. monogénésique [311]. monogenic [311]. Monotonic [1493]. moon [8]. moon-illusion [8]. Moore [1278, 671, 1238, 802, 1306, 1439, 1799]. Moore-Paradoxicality [1278]. Moral [1499, 1240, 1285, 1082, 1525, 1033, 1779, 1609, 1460, 1314, 1766, 107, 1268, 842, 973, 1678, 1409, 1262, 974, 1381, 1035, 1167, 1600, 1353, 1420, 1374, 969, 16, 331, 448, 913, 865, 886, 333, 790, 896, 793, 866, 1224, 1, 1558, 1240, 1355]. Moralbegründung [107]. morale [75, 331, 438, 333]. moralisk [16]. moraliska [1]. moralistic [759]. Morality [1026, 1472, 877, 438, 749, 75]. Morally [1659, 1635]. Morin [176]. Moritz [926]. Morning [376]. morrow [574]. most [1724]. motif [151]. Motivating [1655]. Motivation [1267, 1037, 1426, 1035, 1230]. Motivational [1603]. Motive [236, 113, 231].

Motive-Research [113]. motives [236]. motivforskning [113]. move [766]. Movement [515, 422]. Moving [1721]. Mozarts [6, 6]. Mr. [483]. Multidimensional [680]. Multimodal [1449]. Multiple [993, 1277, 979]. Multivocal [232]. Munthe [1382]. music [809]. Must [1044]. Mutual [1024]. My [1005, 1455, 324, 951]. myterna [12]. Myth [1720, 1405, 1846, 1786]. myths [12].

N [646, 898]. några [26]. nach [175, 293]. Naess [422, 473]. Nagel [1267, 1266, 1265]. Name [283, 438]. Names [1502, 516, 745, 545, 363]. Naming [1746, 1741, 1815, 1747, 1749, 1742, 1748]. Nantes [860]. national [13]. nationalekonomien [13]. Nationality [1207]. Natur [252]. Natural [68, 1070, 210, 840, 729, 993, 1726]. Naturalism [1763, 1118]. Naturalistic [1394]. Naturalizing [1254]. Nature [1630, 1256, 1698, 1241, 1192, 186, 1472, 1197, 252]. naturelle [68]. naturvetenskapliga [26]. Nec [1628]. Necessary [1725, 509, 1450, 1429]. Necessitation [1790]. Necessities [1284]. Necessity [1433, 1746, 1741, 1742, 1747, 1554, 1530, 1749, 752, 1748]. Negation [968, 1154, 299, 383, 993, 472]. Negative [792, 1494, 299]. Neglected [1498]. Negotiation [1822]. neighborhood [709, 724]. neither [1205]. Nelson [326, 326]. Neo [1833]. Neo-Fregean [1833]. Nesting [1649]. networks [1008]. Neural [1701]. Neurath [422]. neurophysiological [685, 660]. neuroscience [885]. Neutral [846, 1759]. Never [1793]. Newman [1438, 1248]. Newtonians [197]. Nicholas [1318]. Nicolas [1415]. Nietzsche [1110]. Nils [1241, 1344]. Nilson [661]. Nis [1532]. No [1468, 1236, 1708, 1453, 1398, 413, 854, 958, 975, 970, 1488]. No-categoricity [413]. nom [438]. nominalism [782]. Nominalist [1837]. Non [1250, 1204, 1658, 1195, 1109, 1515, 684, 1457, 1572, 1222, 1283, 1493, 998]. Non-academic [1572]. Non-aesthetic [695]. Non-Archimedean [1195]. Non-causal [1658]. Non-classical [684]. Non-Cognitivism [1204]. Non-Contradiction [1109]. Non-Epistemic [1515]. Non-Exegetical [1250]. Non-existence [1457]. Non-Humean [1222]. Non-Monotonic [1493]. Non-prioritized [998]. Non-Reductive [1283]. Nonmonotonic [1191, 1757, 1756]. Nonsense [469, 950, 544]. nonsense-logics [544]. nor [1205]. Nordenfelt [1059]. norm [272, 278]. Normal [1089, 709, 724]. Normalcy [1562]. Normative [1209, 1231, 1416, 1581, 955, 1203, 1823, 1255]. Normativism [1373]. Normativity [1480, 1273, 1827, 1808, 1800, 1441]. Normes [860]. Norms [1357, 603, 409, 473]. Not-Doing [854]. Note [474, 323, 324, 279, 288, 337, 573, 1293, 1009, 747, 391, 1140, 347, 719, 707, 598, 786, 822, 479, 644, 1791, 474]. Notes [118, 124, 130, 135, 140, 154, 145, 150, 170, 160, 165, 177, 188, 184, 193, 194, 202, 207, 215, 216, 224, 234, 235, 270, 275, 276, 282, 309, 315, 320, 327, 338, 339, 358, 356, 939, 1003, 399, 402, 938, 1243, 258, 353, 356]. Notice [80, 86, 95, 100, 1136, 973, 974]. Notices [44, 54]. Notifications [5, 9, 14, 20, 27, 30]. Notion [240, 218, 924, 350, 492]. Notions [241, 427, 1273, 642, 1639, 51]. Notis [67]. nouveau [372]. Novel [1677]. November [966, 965]. Nozick [798]. Number [1474, 1362, 1835, 260]. Numbers [1221, 784]. Numerical [908]. Nussbaum [1207]. Nutida [25]. nyere [33]. Nygren [113]. Nyman [263].

obigen [175]. Obituary [1303, 1456, 1295]. Object [127, 870, 813]. objectif [75]. Objection [1438, 686, 1847]. Objections [1142]. Objective [1623, 204, 1197, 75, 901]. Objectivism [901]. objectivitè [83]. objectivity [83, 17, 62]. Objects [446, 1361, 482, 1835, 1836, 1017, 1642, 1832, 963, 1830]. objektivism [17]. Objektivität [62]. Obligation [1433, 1804, 322]. Obligations [1242, 956, 649, 750, 1051]. Obscurantism [1719, 1445]. Obscurity [1694]. Obtained [919]. Ockham [782]. Off [1402]. Offsetting [1630]. Offspring [1798]. Offspring-Determination [1798]. Olav [1127]. Old [1236, 46]. olikhet [26]. Olivecrona [1250]. Olof [899]. Olsson [1201, 1415, 1339, 1676]. Olsson-Yaouzis [1415]. Om [18, 15, 15]. Omitting [760]. Omnitemporal [713]. område [2, 7]. One [1252, 508, 248, 395, 316]. one-many [395]. one-one [395]. O’Neill [1200]. Oneself [1124, 550]. Only [1252]. Ontic [1317]. Ontologese [1542]. Ontological [1789, 585, 452, 932, 305]. Ontologically [730]. ontologischen [305]. Ontology [729, 1429, 1699, 820, 1173]. opacity [737]. Open [1431]. Operational [298]. operations [1757, 1756, 955]. operator [682]. Operators [630, 888, 907]. opportunity [1795]. Opposing [1595]. opposition [321]. Oppression [1415]. Options [1761, 1853]. Order [565, 1307, 609, 1044, 619, 587, 413, 429, 1762]. orders [437]. Organisms [1141]. Organization [1346]. origin [696, 838, 697]. original [1140]. originality [1810]. Orman [965]. Ostenfeld [1532]. Other [316, 1388, 1580, 1158, 1269, 1234, 1134, 550, 245, 734, 970]. otherwise [851]. Otto [422]. Ought [462, 348]. Oughts [1297, 1203]. Our [527, 1394, 1260, 1327, 1385, 1718]. Outline [436]. Outlines [394]. Outrageous [1387]. Ove [1296, 1193]. Overcoming [1130]. Overview [1654, 1830]. Own [1394, 1327]. Owne [1237]. Oxenstierna [88, 15, 293]. Oxenstiernas [15]. Oxford [1533, 1590, 1605, 1175, 1176, 1650, 1166, 1181].

P [1174, 1489, 1234]. P. [1174]. pages [1168, 1172, 1166]. Pain [1774]. Palm [1193]. Pap [586]. Paper [415, 307, 661]. paperback [1176]. papers [1020]. Papineau [1166]. Paraconsistent [1074]. Paradox [1433, 1178, 1060, 1628, 1332, 1140, 166, 1491, 1149, 376, 1487, 586, 1237, 1117, 546, 1238, 1306, 1439, 729, 803, 492, 869, 1799, 1215, 1247, 1309]. Paradoxes [89, 1180, 1010, 1449, 547, 828, 802, 1262]. Paradoxicality [1278]. Paradoxie [37]. Paraphrasability [1777]. Paraphrase [1064]. Parekh [1507]. Parenthesis [1540]. Parenting [1798]. Parfit [1474, 1475, 1473, 1470, 1472, 1619, 1541, 1470]. Parity [1298, 1475, 1452, 1611]. Parliamentarism [559]. Parsing [1411]. Part [1301, 1315, 655]. Partial [1402, 1101, 347, 1853]. Participation [1624].

Particular [1654]. Particularism [1033, 1148, 1228, 1268, 1031]. particularity [1826]. Particulars [920, 1745]. Partitions [1040]. Partnership [1585]. Parts [1141, 1040]. Past [410, 1721, 1790, 891]. Patient [1797]. Patrice [860]. Patricia [1489, 1326]. Patterns [1595, 304]. Paul [1576]. pbk [1590]. Peano [1582]. peculiarity [1816]. Peer [1091, 1384, 1829]. People [1088, 835, 245]. Pepper [416]. Perceiving [1128, 1304]. Perception [846, 1701, 1189, 1379, 1438, 1596, 1524, 1667, 258, 783, 1759, 290, 181]. Perceptual [1504]. Percipi [412, 415, 377]. percipi-Principle [377]. Perdurantism [1597]. perfect [648]. Perfection [1690, 1667]. perfectly [772]. Performance [1480]. Performative [521, 529]. Performatives [495, 800]. Performatory [496]. Perish [1652]. permission [825]. permutations [420]. Perry [1605]. Persistence [1415]. Persistency [1611]. Person [520, 941, 873]. Personal [1259, 1709, 1796, 1473, 1577, 1099, 645, 1780, 1370]. Persons [1141, 1766, 1344]. Perspectival [1528]. Perspective [1469, 1053, 1490]. perspectives [941]. Persson [846, 893, 974]. Petar [1237]. Peter [1214, 1520, 1193, 897]. Petersen [1255, 1453]. Petits [504]. Petrazycki [455]. Petzäll [380]. peut [248]. peut-on [248]. Pflicht [121, 296]. Phaedo [721]. Phalén [41]. Phaléns [41]. Phänomenalismus [131]. Phase [1701]. Phases [1047]. Phenomena [119, 247]. Phenomenalism [131]. phénomène [75]. Phenomenology [1557, 1657]. phenomenon [75]. Philip [1192, 1172]. Philosoph [136]. Philosopher [1684, 136]. Philosophers [1151, 1242, 1609, 1192, 1084, 1233]. Philosophic [871, 34]. Philosophical [758, 1021, 1059, 1226, 780, 1369, 1533, 1563, 941, 953, 1157, 1162, 1235, 1294, 1346, 1501, 1527, 1553, 1592, 1694, 1782, 1810, 36, 1532, 34, 1730, 1490, 1118, 833, 646, 1072, 909, 13, 824, 137, 798]. philosophically [1759]. Philosophie [74, 286, 89, 50, 213, 1325, 61, 218, 329, 371, 96, 185, 295, 208]. philosophiegeschichtlichen [36]. philosophische [137]. philosophize [1219]. Philosophy [1250, 1356, 217, 1028, 1137, 1200, 1225, 1199, 1217, 1216, 113, 1732, 1394, 218, 323, 74, 77, 219, 329, 1022, 1038, 1061, 1068, 1077, 1106, 1113, 1169, 1177, 1185, 1249, 1246, 1264, 1279, 1272, 1302, 1288, 1308, 1319, 1342, 1328, 1351, 1358, 1365, 1371, 1364, 1400, 1383, 1399, 1407, 1423, 1427, 1443, 1448, 1437, 1454, 1463, 1469, 1476, 1481, 1511, 1510, 1509, 1540, 1539, 1534, 1560, 1564, 1616, 1606, 1637, 1682, 1728, 1829, 1174, 50, 1045, 1096, 1389, 1310, 1158, 921, 371, 1627, 1192, 1470, 213, 185, 313, 1172, 1405, 854, 295, 841, 1078]. philosophy [1091, 1754, 940, 61, 964, 286, 89, 952, 990, 1097, 1661, 1758, 33, 677, 829, 904, 959, 96, 382, 579, 823, 208, 1266, 853, 1184, 1590, 1532, 962, 1090, 1223, 1742, 1435, 1506]. Philosophyland [1377]. Phosphorus [819]. Photographic [1602, 1826]. Physical [446, 1554, 174, 796]. Physicalism [1773, 1783, 131]. physicalistic [249, 264]. Physicalists [1378]. physicist [40, 41]. Physikalismus [131]. Physikalistischer [264]. Physiker [40, 41]. physique [241]. Pictorial [1465, 1602, 705]. Picture [1751, 831]. Pihlström [1422]. Place [1143, 329]. Plagiarism [1563, 1235, 1553, 1626, 1821]. Plagiarize [1652]. Planetary [250]. Plantinga [845]. Plato [870, 721, 666, 288, 111, 534, 954]. Platon [11]. Platonic [82]. Platonists [1418, 1234]. Platons [111]. Plausible [1544, 1557]. pleasing [835]. Pleasure [1667]. Pluralism [1302, 1597, 1422, 1817]. Poetics [1628]. point [17]. Points [482]. Polarity [510]. polemics [101]. Policy [1369, 1358, 1715]. Politeia [111, 111]. Political [1688, 1296, 1110, 1723, 1424, 12, 1078, 1090]. politically [1843]. politiska [12]. polyglottism [941]. popper [946]. Population [1474]. position [695, 765]. positive [792, 620]. positivism [264]. Positivismus [264]. Possessive [1411]. Possibility [1366, 763, 1597, 854]. Possible [1252, 607, 230, 732, 914, 471]. Post [1567]. Post-Truth [1567]. Posteriori [1284]. Posteriority [1451]. Posts [21, 28, 31, 15]. Postulate [1764, 498, 347]. potentiality [848]. Power [1667, 848]. Powers [1650]. pp [1533, 1624, 1184, 1590, 1532, 1183, 1173, 1551, 1605, 1175, 1192, 1176, 1650, 1167, 1181, 1191, 1558]. pp. [1193]. Practical [1634, 1285, 1299, 1210, 567, 1472, 1715, 334]. practice [766, 1796, 793]. practices [1848, 1850]. Pragmatic [1505, 1601, 1422]. Pragmatism [921, 1287]. Praiseworthy [1478]. Praising [626]. praktische [334]. Prawitz [1021, 1015, 1017, 1791]. Pre [277]. Pre-sentence [277]. Precaution [1382]. Précis [1831]. preciseness [253]. Precision [1371]. précriminelles [277]. Predatory [1431]. Predicament [1331]. Predicate [565, 240, 413]. Predicate-in-Notion [240]. Predicates [1645, 620]. Predication [1109, 611]. Prediction [553]. predictions [189]. Preemption [1403]. Preface [981, 992, 1002, 937, 1123, 1063, 918, 1029, 1046, 869]. Preference [1452, 987, 986]. Preferences [1025, 988, 891]. prescriptive [463, 464]. present [410, 900]. Presentism [1450, 1239]. Presentists [1418]. Press [1533, 1590, 1183, 1173, 1605, 1175, 1192, 1168, 1176, 1172, 1650, 1166, 1167, 1182, 1558]. presumptions [470]. Presuppositions [470, 1023]. Prevention [318, 318]. Previews [1412]. Price [1382, 1483, 13]. Prima [649]. Primacy [846, 1300]. Primary [227]. prime [1847]. prime-mammal [1847]. primitive [262]. Principle [1544, 240, 1278, 412, 1147, 377, 1827, 1808, 1086, 815, 415]. Principles [1767, 1775, 1099, 1003, 1031, 807, 1791, 1175, 386]. Prior [645, 674, 646, 808]. Priori [509, 1750, 1041, 1695]. Prioritization [927]. prioritized [998]. Priority [1451, 1597, 1417]. Prisbildning [13]. prison [291]. prisoner [883]. Privacy [1193]. private [2, 7]. private-rights [2, 7]. privaträttens [2, 7]. Prize [1266, 1267, 1265, 1391, 1535, 962]. Proaction [1632]. Probabilistic [1581, 1197, 747, 762]. probabilities [901]. Probability [1201, 312, 489, 56, 346, 460, 467, 583, 924, 408, 488, 923]. Problem [1562, 1716, 1556, 176, 36, 50, 1079, 1337, 462, 1036, 1049, 1384, 1676, 1458, 1706, 173, 1338, 410, 973, 1649, 1466, 313, 257, 428, 974, 1840, 421, 157, 1422, 25, 48, 245, 24, 8, 660, 90, 830, 13, 700]. Problem-Critique [50]. Problematik [36]. problème [218, 48, 265]. Problèmes [96]. Problemkritik [50]. Problems [1213, 218, 494, 1580, 267, 229, 1761, 25, 426, 1778, 1140, 550, 303, 265, 96, 1181]. procedure [432]. procedures [960, 498]. proceed [959]. Process [1067, 1337, 1375, 919]. Processes [537, 247]. processing [993]. Procrustean [1798]. Product [1067]. Professeur [263]. Professor [263, 717, 1573, 279, 714]. Profile [1643]. Programmatic [1202]. Progress [1365, 1469, 1476]. Project [1689]. Projection [1465, 942]. Prolegomena [50, 504]. prolégomènes [504]. prolegomenon [1807]. Promises [480]. Proof [1516, 1487, 1111, 1726, 452, 305, 432, 738]. Proof-Theoretic [1516, 1487, 1726]. Proofs [1017, 736, 874, 900]. Propaganda [32]. Proper [1634, 1141, 946]. Properties [1494, 1035, 729, 950, 942, 674]. Property [1408]. Proposal [1719, 1761]. proposed [272, 278, 382]. Proposition [1352, 1117, 1359]. Propositional [607, 657, 1360, 1153, 595, 1347, 347]. Propositionalism [1503]. Propositions [271, 1039, 1503, 330, 1518, 604, 186, 1647, 172]. Protect [1770]. Protection [1797]. Provability [983, 776, 752, 1825]. Provide [1420]. Providing [1731]. Prudence [891, 1322, 890]. pseudo [545]. pseudo-names [545]. Pseudophilosophy [1620]. Pseudoscience [1495, 1331, 1620]. Psychoanalysis [1331]. Psychological [1259, 246, 190, 277, 249, 1812, 264]. Psychologie [290, 211, 51, 156, 255, 161]. psychologiques [277]. Psychologism [1485]. psychologistischer [264]. Psychology [161, 110, 1479, 421, 51, 290, 156, 255, 211, 1285]. psychometrics [267]. Psychopaths^{≫} [198]. Psychopathy^{≫} [198]. Psychophysical [584]. Psychophysics [161]. Psychophysik [161]. Public [1369, 1358]. Publications [1021, 1191]. publicity [862]. Publish [1522]. Published [1672]. Publishing [1328]. Punir [226, 228]. Punish [226, 228, 225]. Punishment [227, 203, 381]. Purely [796]. Putnam [1199, 1394, 1395, 1392, 1396, 1395, 1391, 771]. Puzzle [1545, 1833, 1738, 887]. Puzzlement [1452]. Puzzling [1668].

Qasim [1731]. qualia [906, 900]. Qualities [541, 881, 620, 693]. quality [692, 461]. Quantenlogik [519]. Quantification [1432, 454, 1079, 643]. Quantifiers [565, 903, 908, 657, 643, 479]. Quantitative [285]. Quantities [1554]. Quantum [1712, 519, 382]. quantum-mechanics [382]. Quelques [51]. Question [1228, 1357, 1444, 233, 102, 954]. Questions [903, 1744, 1017, 1255]. qui [438]. Quine [967, 965, 1188, 1388, 964, 494, 1023, 1361, 1621, 1012, 712, 1182, 1430]. Qunawi [1736]. quu [868].

R [841, 1541]. R. [1058]. Rabinowicz [891]. Radical [804, 1087, 1699, 1238]. Raimo [855]. Raining [601]. Ramsey [1650, 1500, 922, 924, 1505, 1170, 933, 921, 920, 923]. Randomness [442, 489]. rapports [331]. Rasmussen [1241, 1370]. Rational [1731, 1697, 988, 817]. Rationalist [1656, 834]. Rationality [487, 1415, 1194, 1150, 404, 1181]. Rationalization [1681]. Rawls [1115, 816, 815, 773]. Razi [1698]. Razón [1685]. Razón/Corazón [1685]. Re [841, 314]. Re-enactment [841]. Reader [1044, 1174]. Reading [874]. Real [1643, 1326, 689]. Realism [1082, 727, 243, 1361, 1419, 1621, 1353, 1420, 1593, 931, 1147, 33, 876, 866]. Realisme [33]. realisms [771]. Realistic [1069]. Realität [302]. Realities [1637]. Reality [603, 175, 142, 147, 302]. Realizations [1277, 1706]. Really [1534, 887]. Reason [1285, 227, 1735, 1685, 137]. Reason/Heart [1685]. Reasonable [1822]. Reasoning [1464, 1681, 1493, 1581, 1472, 1757, 836, 1756]. Reasons [1379, 1357, 1081, 1503, 392, 1717, 1152, 1034, 546]. Rebut [1619]. Received [105, 171, 178, 195, 206, 310, 328, 340, 359, 365, 375, 385, 424, 441, 451, 458, 466, 475, 485, 493, 514, 526, 532, 557, 570, 591, 600, 625, 670, 681, 698, 688]. Recherche [68]. recipes [503, 984]. Recipient [1266]. Recognition [1499, 1465]. Reconsidered [1659]. Reconsidering [1603, 1631, 1099]. Reconstructing [1240]. Reductio [1619]. Reduction [710, 725, 685, 391, 660, 1791]. Reductionism [1374]. Reductive [1283]. Redundancy [1142]. Referee [1599]. Referees [1390]. Reference [1430, 1372, 1154, 1243, 1837, 1605, 1643, 1744, 1440, 640, 1750, 1196, 1789, 743, 847, 249, 517, 1447]. Referential [1548]. Referentialism [1291]. Referring [1124]. Refined [1589, 707]. reflection [1003]. Reflections [469, 985, 46, 1748, 266, 318, 381, 386, 744, 1102]. Reflexionen [386]. Réflexions [318]. reflexive [1006]. Reflexivity [1610]. Refugees [1507]. Refutation [1368, 989]. Regimenting [1152]. Regress [1467]. Regularity [342]. Rehearsals [1412]. Reichenbach [386]. Reichenbachs [386]. Reid [1225, 1183]. Reidel [855]. Reimagined [1640]. Rejected [1672]. rejection [877]. Relata [1274]. Related [595, 233, 444]. Relation [870, 1556, 462, 1541, 1340, 57, 395, 955]. Relational [1438, 1772, 656]. Relations [1240, 1429, 564, 622, 1229, 911, 1819, 331, 716, 654, 863, 979]. Relationship [246]. Relative [1009, 987, 797]. relativism [1188, 795]. Relativitätstheorie [189, 41]. Relativity [1736, 41, 189, 1423, 292, 505, 492, 497]. Relevance [850, 826]. Relevant [879, 1671]. Reliabilism [1556, 1337, 1338]. Reliability [1524, 1339, 1111]. reliable [919, 915]. Reliance [1634]. Religion [220, 3]. religious [1205]. Remark [280]. Remarks [1395, 1747, 41, 245, 415, 661, 307, 959, 699, 410, 491, 83, 279, 268, 57]. Remarques [83]. Repair [1240]. Repentant [1580]. Replaceability [1718]. Replies [1841, 1142]. Reply [687, 717, 1468, 464, 1215, 1488, 1614, 1453, 951, 1336, 897, 895, 891, 893, 1460, 1339, 1419, 1135, 1131, 1133, 1129, 1125, 1127, 1408, 1338, 1401, 1054, 974, 1381, 483]. Report [1599]. Reported [1698]. Reports [853, 854]. Represent [1536]. representability [913]. Representation [1412, 705]. Representational [1170]. Representationalist [1783]. Representations [1538]. Reproductive [1798]. Republic [288]. Republicanism [1526]. Repugnance [1171]. Repugnant [1475, 1471, 1171, 1088]. requirements [137]. Rescuing [1373]. Research [1028, 113, 1346, 1491, 1584, 259, 301]. resemblance [597, 579]. resentment [884]. Resorted [1136]. Respect [1766]. Response [1438, 1291, 1800]. Response-dependence [1800]. Responses [1409]. Responsibility [1702, 229, 1492, 842]. Restriction [1518]. results [684]. Résumé [246]. Resümee [246]. retardation [492]. Retracted [1607]. revealing [277]. Revelation [1735]. révélatrices [277]. revenge [851]. Revievs [99]. Review [71, 123, 821, 841, 846, 860, 864, 875, 878, 882, 889, 910, 929, 936, 1112, 1200, 1201, 1198, 1199, 66, 818, 853, 1415, 855, 1533, 917, 711, 726, 1624, 307, 1184, 765, 512, 871, 416, 1406, 1363, 1590, 1532, 1183, 1447, 787, 1382, 1520, 1551, 1605, 798, 885, 1192, 898, 697, 1650, 422, 769, 1181, 1193, 854, 1191, 1182, 1558, 1091, 175, 1829]. Reviewed [1672]. Reviewers [1672]. Reviews [1462, 1482, 1356, 43, 53, 60, 79, 85, 94, 104, 109, 117, 129, 134, 139, 144, 149, 153, 159, 164, 169, 183, 187, 192, 214, 274, 281, 308, 319, 357, 345, 364, 368, 669, 676, 703, 715, 753, 779, 806, 837, 905, 961, 971, 980, 1013, 1059, 1058, 1071, 1076, 1083, 1095, 1155, 1161, 1212, 1214, 1223, 1224, 1207, 1225, 1213, 1226, 113, 1090, 1388, 1430, 1369, 1240, 1255, 1285, 1060, 1428, 1508, 1483, 1256, 1318, 1043, 223, 1370, 1441, 1241, 176, 1233, 957, 1343, 1442, 1144, 1174, 1248, 1286, 1389, 1119]. Reviews [1239, 1310, 1344, 1435, 1429, 1489, 1311, 1173, 1357, 1042, 1120, 1145, 1461, 1507, 1287, 1121, 1175, 1326, 1206, 1168, 1262, 1490, 1261, 1232, 1176, 1234, 1172, 1254, 1057, 1247, 1166, 1167, 1436, 1044, 1506, 1421, 1309, 1317, 1325, 1355, 1422, 201]. Revised [663]. Revision [1464, 1073, 1493, 1074, 801, 997, 996, 1001, 998, 984]. Revisionary [1139]. Revisited [1245, 1632, 1723, 1621, 933, 615]. Revisiting [1115]. Revolution [1415, 632]. Rhetoric [1249, 1183, 543]. Richard [1508, 1213]. Ricoeur [1576]. Right [226, 1451, 1779, 1312, 228, 225, 1801]. Rightlessness [1461]. Rights [1107, 559, 1108, 1461, 1669, 2, 7]. rigid [804]. Rigidity [1649]. risk [990, 1311, 1382]. Rites [1411]. Ritter [357]. Rivals [1477]. RLE [1193]. RLE.- [1193]. Road [1604]. Robert [1310, 1287, 1225, 798]. Rodin [1213]. Roeser [1311]. Roger [1182]. Role [1747, 1485, 1530, 1662, 942, 935, 1765, 790]. Rolf [1474, 1394, 1267, 1395, 1475, 1535, 1392, 1471, 1473, 1393, 1470, 1472, 1396, 1266, 1265, 1391, 425, 962]. Rønnow [1370]. Rønnow-Rasmussen [1370]. Roots [1105, 743]. Rosmini [361]. Ross [1140, 1215, 803]. Round [1328]. route [1005]. routine [947]. Routledge [1174, 1551]. Roux [1174]. Rudolf [167]. Rudolph [422]. Rule [1268, 1010, 399, 402, 357]. Rule-Following [1268]. Rules [1499, 1032, 1848, 566, 411, 641, 1792]. Runde [1184]. Russell [326, 324, 1325, 326, 778, 805, 729, 1425, 1438, 1039, 306, 777, 1117, 979]. Russellsche [306]. Ruth [1535]. Rutherford [1223]. Ryberg [1255].

S [1255, 1533, 841]. S4 [718, 706, 701, 1725]. S4.1 [593]. S4.3 [718]. S5 [1725]. Sa [47]. Sabine [1311]. Sacrifices [1044]. Safety [1410, 1488]. Saharon [1654, 1692]. said [651]. Sailing [1684]. Saint [333]. Sajama [853]. Salerno [1309]. Sally [1226]. Salt [1543]. Same [1236, 1446]. Sami [1422]. Sampling [312]. Samuel [219]. Sandøe [897]. Sankt [1532]. Satisfaction [1244]. Satz [322, 296]. Sauer [1558]. Saul [1447, 1262, 1742]. Save [1044]. Saying [1287]. scale [554]. scale-types [554]. scaling [267]. Scandinavia [1627, 361]. Scandinavian [369, 379, 396, 406, 423, 440, 465, 500, 513, 525]. Scandinavie [361]. scenarios [1816]. Sceptical [1736, 1735, 1738, 1397, 1737, 1816]. Scepticism [1459, 1732, 1350, 1341, 1736, 1733, 1501, 1729, 1730, 1087, 1401, 1160, 728]. Scepticisms [1731]. Schaffer [1283]. Schiller [1785]. Schmidtz [1499]. Schock [1474, 1266, 1394, 1267, 1395, 1475, 1265, 1391, 1535, 1392, 1471, 1473, 962, 1393, 1470, 1472, 1396]. Scholastic [323]. schooling [289]. Schools [1162]. Schopenhauer [823, 295]. Schopenhauers [295]. Schurz [1435]. schwedischen [185]. Science [1495, 1048, 1106, 1319, 50, 306, 1435, 1042, 1506, 69, 1770, 677, 249, 26, 303, 137, 106, 1225, 1363]. Sciences [62, 209, 83]. Scientific [922, 1723, 1621, 1787, 1205]. Scientifically [1118]. scientist [977]. Scientists [1506]. Scope [1101, 1018]. Scotus [1707]. Screened [996, 997]. Screening [1402]. Sculptural [1602]. Se [1607, 1648]. Seanor [898]. Search [1835, 68]. Searle [873]. Second [1684, 1211, 619, 587]. Secondary [881]. security [934]. Sed [1628]. Seeing [880, 1515, 1360, 772, 1602]. seelische [246]. Segerstedt [15]. Segerstedts [15]. Selection [617, 740]. Selective [655]. Self [1664, 1684, 1548, 1798, 1472, 1814, 943, 1198, 1060]. Self-Consciousness [1060]. Self-Control [1664]. Self-Determination [1798]. Self-Interest [1472]. Self-Referential [1548]. Sellars [1130, 1405]. Semantic [1394, 330, 641, 1570, 1784, 1757, 1756, 633, 647, 655]. Semantical [1250, 997]. Semantics [1732, 1538, 1395, 1067, 1251, 1516, 1070, 1393, 1704, 463, 464, 1759, 908, 712, 1726]. semi [1001]. semi-revision [1001]. Semicompatibilism [1776]. Semiotic [1643]. Sensationalism [377]. Sensations [541, 537, 490]. Sense [1341, 446, 1243, 1498, 1440, 268, 58, 259, 444]. Senses [482, 1792]. sensible [242, 1745]. Sensitive [1276]. Sensitivity [1410]. sentence [322, 296, 277]. Sentences [581, 496, 1170, 1548, 330, 712, 391, 1759]. sentiment [90]. separated [911]. Seppo [853]. September [193]. Serena [1507]. Series [841, 1275]. Serious [1418, 1048]. Seriously [1140, 1291]. ses [331]. Set [1655, 1437, 629, 684]. Setpoints [383]. sets [347, 498]. Seventy [1257]. Seventy-Five [1257]. Several [630]. Shall [1043]. Shared [1213]. Sharpening [824]. Sharpness [1414]. Sheer [1650]. Shelah [1654, 1692]. Shifting [1514]. Shifts [1351]. shifty [1203]. Shoemaker [1271, 582]. shoes [986]. short [983]. Should [1260, 1055, 1332, 1052, 1496, 1843, 1672, 1696]. Shouldn’t [1555]. Show [1686]. Shown [1036]. Sibley [695]. Sick [1580]. Side [1445]. Sideboard [1671]. Sider [1542]. signalling [830]. significance [969]. Significant [1030]. Signification [370]. Signs [1688]. Similarity [1640, 1588, 26]. Simple [1721, 629, 751]. Simulation [1624, 1363]. simultaneity [497]. Singular [1372, 1695, 1274]. Sinner [1580]. Sinnes [444]. Sinnesdatum [279, 268]. sinnesjämvikt [18]. Situation [141]. situations [277]. Size [1102, 1103, 40, 41, 969]. själslivets [24]. Skepticism [1413, 1245, 1734, 1118, 1406]. Skolem [1178]. Slavery [1617]. Sleep [596]. Slogan [1099]. slurs [1792]. slutinlägg [15]. small [999, 504]. Smenkhkare [1744]. Smilansky [1262]. Smith [1318, 973]. So-called [318, 220]. Soames [1368]. Sobel [1303]. Social [62, 283, 1331, 855, 1763, 209, 1489, 203, 1803, 83, 75, 1852, 1850, 106, 12]. sociales [83]. socials [12]. Sociétés [213]. Societies [213]. Society [297]. Socio [1301, 1315]. Socio-technical [1301, 1315]. Sociological [230]. Sociology [285, 1560, 127, 97]. Sociometry [266]. Socrates [251, 733]. Socratic [546]. soi [318]. soi-disant [318]. Sokrateslarjungen [11]. Solid [1000]. Solidarity [1586]. sollen [322]. Sollsätzen [383]. Solubility [1543]. Solution [1433, 1332, 1676, 1349, 830]. Solutions [1384, 96]. Solving [1449]. Some [1250, 426, 1395, 1203, 285, 219, 744, 1747, 266, 494, 307, 1243, 1498, 415, 410, 1102, 1737, 258, 575, 301, 491, 381, 1017, 303, 240, 386, 286, 911, 478, 498, 51, 252]. Sommers [531]. son [47]. Sonderholm [1460]. Soper [1172]. Sorabji [1213]. Søren [6, 239]. Sorites [1290, 1289, 1290]. Sorter [1136]. Sorting [1058]. Soul [1698, 1192, 24]. Soul-Switching [1698]. Source [1379]. Sources [416]. Sovereignty [1585]. Sozialwissenschaften [62]. Space [342]. spacetime [876]. Sparks [1508]. Speaking [468, 730]. Special [1775, 189, 1154, 233, 249, 505, 492, 497]. Species [679, 337]. Spectral [1602]. Spectrum [1611]. Speech [1463, 1572]. speziellen [189]. Spinoza [611, 1085]. spirit [252]. Spiritual [220]. Spotlight [1553]. Sprache [50]. Sprachphilosophie [50]. Sprachwissenschaft [306]. St [333]. St. [325]. Staat [111]. Stance [1416]. Standing [1634]. Standpoint [1014, 1232]. Standpunkt [306]. Star [376]. stare [867]. Starving [1044, 791]. State [1724, 277, 331, 745, 111]. Statements [560, 408]. States [1519, 249]. Statistical [1562, 460, 467, 1613]. Status [1777, 481, 1177, 1725, 932, 647]. Steglich [1453]. Steglich-Petersen [1453]. Stegmüller [345]. Sten [1755, 1756]. Stenius [833, 829, 828, 827, 825, 1293]. Stephan [1176]. Stephen [1489, 416]. stereotypes [1849]. Still [1453]. stimuli [712]. Stimulus [494]. Stipulation [1695]. Stockholm [113]. Stone [1163]. Stop [1560]. strain [759]. Strange [1269]. Strangers [1286]. Strategic [1538]. Stratified [1073, 572]. Strawson [728, 805, 1341, 520, 582, 1814]. Strict [816, 233]. Strong [1529, 472]. Structural [1702, 255]. structurale [255]. structuralist [1772]. Structure [1689, 174, 1376, 1472, 436, 705, 876, 336]. Structures [564, 1417, 679, 707, 678]. Struggles [1461]. stubborn [737]. student [861]. Studentenvereinigung [120]. Students [1579]. Studies [1532, 697, 1172, 854]. Study [1797, 1286, 389, 430, 456, 841, 818, 233, 765, 911, 331, 418, 504, 1807, 199, 1057, 1436]. Suárez [1738]. Sub [1610]. Sub-utterances [1610]. Subcontrariety [1154]. Subdisciplines [1740]. Subdivision [627]. Subjective [881, 1523, 204, 924]. subjectivism [81]. Subjektivismus [81]. Subordinate [1530]. Substance [1198, 611, 623, 527]. Substances [621, 1738, 623]. Substantial [284, 1109]. Substitutional [1079, 643]. Success [1538, 919]. Successful [1646]. suédoises [418]. suffering [120]. sum [540, 335]. summary [199]. summing [893, 892].

Sums [1838]. Superduper [1378]. supereminence [931]. Supererogation [1478, 839]. Superfluity [1160]. Supervaluationism [1813]. Supervenience [1378, 1270]. Supplement [1767]. Surveillance [1551]. survey [1006, 388]. Survival [1577]. Survive [1116]. Suspension [1697]. Sven [1489, 1296, 1193]. Svenska [113]. Swamping [1721]. Sweden [780]. Swedenborg [252]. Swedenborgs [252]. Swedish [1137, 1096, 12, 5, 9, 14, 15, 20, 27, 21, 28, 31, 22, 29, 1217, 1216, 25, 2, 7, 1061, 1077, 1113, 1185, 1246, 1272, 1308, 1342, 1364, 1399, 1427, 1454, 1481, 1510, 1509, 1539, 1564, 1606, 1682, 1758, 30, 24, 16, 418, 18, 8, 19, 185, 23, 199, 13, 17, 26, 1]. Switching [1698]. Sydenham [238]. Syllogism [334]. Syllogismus [334]. Symbol [246]. Symbolbegriff [34]. Symbolbegriffs [63]. Symbolic [313, 63]. symbols [34]. Symmetry [1669]. sympathy [262]. syndérèse [333]. syndication [333]. Syntactic [767]. Synthesis [243, 921]. Synthetic [509, 824, 353]. System [233, 1301, 1315, 840, 156, 438]. système [156, 438]. Systems [250, 776, 739, 429, 382, 498].

T [15, 438]. Table [1328]. tableaux [655]. Taking [704, 1140, 1291]. Talja [854]. Talking [1059]. Tamar [1533]. Tännsjö [1167, 1357]. Tarskian [1791]. Taste [696, 1709, 1678, 695]. Tatievskaya [1325]. Teach [1052]. teacher [861]. Technical [1312, 1329, 1301, 1315]. Technicity [1301, 1315]. Technological [1311]. Technology [1346]. Teleologie [173]. teleology [173]. Tell [1686]. Temkin [1102]. Temporality [1631]. Temps [860, 504]. Temptations [1445]. Ten [1390]. Tennant [1778]. Tense [1433, 572, 653, 707, 738, 710, 725]. tense-logic [738]. tensor [394]. teoretiska [13]. term [237, 268]. Terminology [1157]. Terms [1372, 1314, 1396, 620, 682, 444]. ternary [682]. Terry [1286]. Tersman [1224]. Test [1500, 1679, 933]. Tested [1118]. Testimony [1614, 1587]. text [372]. texte [372]. Texts [360, 1191]. Thaddeus [1436]. Their [1404, 1670, 482, 1643, 1340, 801, 950, 970]. Them [1429, 1600]. Themes [1250]. theodicy [793]. Theology [1731]. theorem [619, 420, 702, 969]. theorems [767]. Theoretic [1516, 1487, 629, 1726, 1791]. Theoretical [853, 161, 1665, 24, 908, 13, 189]. theoretischen [161]. Theoria [846, 175, 1053, 279, 35, 1137, 1217, 1216, 1334, 175, 1061, 1077, 1113, 1156, 1185, 1227, 1246, 1257, 1258, 1272, 1308, 1342, 1328, 1364, 1399, 1427, 1454, 1481, 1510, 1509, 1573, 1096, 1158, 1335, 1228, 1296]. Theorie [306, 146, 311]. Theories [481, 1549, 922, 447, 592, 921, 561, 847, 584, 1006, 913, 324]. Theorist [1696]. Theory [1115, 1655, 1640, 1187, 1508, 855, 1438, 41, 1011, 189, 306, 1243, 551, 921, 146, 1723, 1150, 455, 1323, 1380, 629, 1280, 1466, 516, 1117, 1056, 157, 430, 1424, 1654, 1317, 1066, 1817, 801, 311, 1793, 436, 729, 583, 858, 1802, 766, 684, 682, 1008, 888, 503, 417, 740, 1796, 638, 253, 1016, 410, 885, 1785, 690, 815, 545, 641, 773, 141, 505, 492, 497, 979, 262, 1167]. There [1567, 1079, 909, 958, 438, 800]. therefore [335]. Theses [1137, 1217, 1216, 1061, 1077, 1113, 1185, 1246, 1272, 1308, 1342, 1364, 1399, 1427, 1454, 1481, 1510, 1509, 1539, 1564, 1606, 1682, 1096, 1758]. Thesis [1529, 1804]. Thin [1830, 1372, 1839, 1838, 1835, 1836, 1832]. things [730, 689, 772, 628]. think [335]. Thinking [1547, 1485, 334, 1722, 896, 141, 211, 1166]. Third [213]. Thomas [1225, 1255, 1266, 1267, 1265, 895, 1183]. Thomason [707]. Thorild [114]. Those [1672, 1671]. Thought [1695, 181, 247, 1745, 1489]. Thoughts [1557, 1297, 1203, 286]. Three [1549, 575, 984, 863, 1639, 453, 664, 373]. Three-valued [575, 453, 373]. tidsproblemet [176]. til [4]. till [26]. Time [1025, 789, 1631, 1597, 295, 835, 982, 713, 719, 504, 1100, 652, 673, 256]. Timothy [1320]. Titel [175]. title [175]. To-morrow [574]. today [422]. Token [1610]. Token-Reflexivity [1610]. tolerance [1220]. Tomatoes [484]. Toni [1370]. Too [1378]. tool [824]. Topics [1333]. Topot [787]. Torbjörn [1357, 1167]. Torgny [15]. Torture [1355]. Totality [1519]. Toulmin [488, 392]. Traces [1683, 1685]. Tracing [1691]. Tractatus [1153, 746, 1280]. Tradition [1730]. Traditions [1213]. Transcendental [1160]. transform [422]. transformation [836]. Transformative [1733]. transiency [785]. Transitions [1685]. Transitivity [1402, 756]. Translation [1563, 1149]. Transmission [1731]. Transmit [1587]. Transparent [1836]. Treatise [1256]. Treatment [198]. trend [249]. Triad [1597]. Tribunal [1628]. trick [977]. Tricks [1236]. Trilemma [1270]. Trivial [1577]. Trivialism [1834]. Tropes [1657]. Tropman [1419]. Trouble [1286]. Troubles [845, 311]. True [1339, 1714, 1321, 1338, 560, 1359, 1846, 212, 409, 473, 1042]. Trust [1555, 1793]. Truth [1201, 1430, 1502, 68, 1067, 1014, 341, 1505, 1567, 1725, 1039, 1819, 1714, 1548, 921, 1486, 531, 58, 330, 1744, 1056, 430, 1622, 46, 1019, 665, 743, 48, 372, 1807, 907, 1811, 517, 652, 785, 373, 261, 923, 1605, 1318]. Truth-conditional [1067]. Truth-Makers [1744]. Truth-value [1819, 652]. truth-values [517]. Truthmaker [1519, 1817, 1811]. Truthmaking [1450]. Try [1425]. Trying [1529, 511]. Tuomela [855]. Turku [853, 854]. Turn [1251]. Turner [1489]. Twelfth [1732]. Twelfth-Century [1732]. Twenty [1742]. Twenty-First [1742]. Twin [1394, 1460, 1409, 1381]. Twin-Earth [1394, 1381]. Two [1252, 1313, 1595, 447, 1279, 1329, 1273, 553, 699, 97, 1409, 815, 497, 751, 561, 900, 382, 290]. Type [1332, 434]. Typed [1332]. Types [1777, 97, 629, 760, 554]. Typologies [1730].

Übersicht [388]. Überzeugung [162]. Ubuntu} [1699]. Ultimate [1477, 1023]. Unavailability [1618]. Unbelievable [1297]. Unconscious [542]. uncountable [760]. undecidable [748]. Undercutting [1459]. Underdetermination [1459, 1816]. Underdetermination-Based [1459]. understand [40, 41]. Understanding [1562, 1614, 1248, 1148, 1587, 1761, 1531, 1179, 1164, 1787, 874, 1167]. Undoing [1237]. Unexplainable [1658]. Unfilled [1352]. Unger [1245]. Unified [1435, 1602]. Uniform [1384]. Unifying [1068, 1639]. Uniqueness [1725, 1446, 844, 746]. United [249]. Unity [1117, 943]. Universal [659, 843, 680]. Universalism [1594]. Universality [1653].

Universalizability [893, 892]. Universals [920, 1657]. Universität [120]. Université [860]. University [853, 1533, 1590, 1183, 1173, 1605, 1175, 1192, 1168, 1176, 1172, 1650, 1167, 854, 1182, 1795, 1090, 120]. Univocal [232]. Unjust [1104]. Unmasked [1282]. Unpublished [1293]. unreal [761]. Unreality [622]. Unreliable [1381]. Unsharp [1414]. unsound [1793]. Untersuchung [121]. Untersuchungen [326]. unveiled [718]. upon [289]. upprätthållande [15]. Urban [1240]. Urteil [162, 294]. Usage [370]. Use [1644, 496, 448]. Useful [1491]. Uses [1501, 428, 946]. Utilitarian [987, 807, 686, 708, 723, 883]. Utilitarianism [741, 750, 987, 816, 744, 598, 986]. utilities [817]. utility [1086]. utmost [26]. Utterances [1610, 1292]. Uwe [1317].

V [967, 1388]. Väänänen [1261]. Vacuous [341]. Vague [1645]. Vagueness [1289, 1290, 1514, 1674, 976, 597, 1318]. Valeur [311]. Validity [1204]. Valpola [326]. Value [1700, 1241, 1337, 1676, 1674, 1513, 1338, 1229, 186, 172, 351, 157, 311, 987, 693, 1819, 638, 913, 13, 652, 1370, 416]. Value-Judgement [351]. valued [453, 575, 373]. Values [1195, 244, 510, 1801, 839, 517]. Valuing [1281]. Vannérus [196]. värde [13]. Variability [1649]. Variable [508]. Variants [341]. Variation [740]. Varieties [325]. vegetables [484]. Veli [326]. Veränderung [146]. verbal [1818]. Verbs [588]. Verhaltens [211]. Verhältnis [57]. Verification [285, 362]. verificationism [1015]. Vérité [68, 48, 372]. Verklighetsbegreppet [175]. Verlag [1532]. Verpflitchtung [322]. Version [1030, 1323, 1015]. versteht [40, 41]. versus [1430, 1099]. verwandte [444]. VI [145]. via [1797, 555]. Viennese [19]. View [1659, 1554, 1486, 302, 17]. viewed [190]. Viewpoint [306]. Views [447, 1296]. VII [150, 1173, 1167]. viljeförklaring [2, 7]. Vincent [47]. Vindication [506]. Violations [1203]. Virtue [1340, 1785]. Vision [1011]. Visual [1701, 1657, 694]. vividness [386]. vocabulary [697]. Voices [1627]. Vol [1532, 1191]. Vollmann [1624]. Volume [1741]. vom [306, 294]. Vorausseteungen [189]. Vortrag [120]. Voting [1669]. vs [1016].

W [1388, 1428, 1508, 1233, 1343, 1310]. W. [967]. Wahrheitsbegriff [46, 261]. Wahrnehmen [181]. Wahrnehmung [290]. Wahrscheinlichkeitsbegriff [56]. Walker [1240]. Want [1696]. War [1213, 1355]. Warming [1421]. Warrant [1530]. Was [40, 81, 41]. Water [1396]. Way [621, 1486]. ways [1164, 863]. Weak [1763, 472, 675]. Wedberg [1294, 757, 758, 175]. Wedbergs [175]. weight [969]. Weighting [742, 764, 814]. Welcome [1377]. Welfare [1088]. Well [1340, 641, 803]. Well-being [1340]. well-formed [641, 803]. Welt [167]. Wenn [180]. were [986]. Westermarck [292]. Wetterström [895]. WH [1744]. WH-Questions [1744]. Wheatley [483]. Whether [1359]. Which [1700, 233]. Whistle [1566]. Whistle-Blowing [1566]. White [1448]. Who [791, 1052, 1496, 1672, 1455, 1064]. Whole [1244, 393]. whom [791, 1098].

Wie [422]. wienerkretsen [19]. Wikforss [1135]. Wikner [176]. Will [231, 700, 853, 295]. Willard [965]. Wille [295]. William [357, 64, 1316, 262]. Williamson [1320, 1724]. Window [1731]. Winsberg [1363]. Wirklichkeitsbegriff [175]. Wirklichkeitsbegriffs [142, 147]. Wisdom [1532]. Wissenschaft [50, 137]. Wissenschaftsgeschichte [69]. Wissenschaftskrisis [69]. Wissenschaftstheorie [871]. Without [1746, 1289, 1290, 1175, 754, 1829, 1768, 1844, 1785, 533, 901]. Witt [1226]. Wittgenstein [1425, 922, 982, 1039, 1268, 746, 1117, 972, 579, 823, 1325]. Wittgenstein-Russell [1039]. Wittgensteinian [781]. Wlodzimierz [891]. Wogau [930, 279]. Wolff [1369]. Wolfgang [345]. Women [1085, 1288, 1627]. Word [518, 307, 287]. Words [1550, 1536, 428]. Work [1425, 1662, 1326, 15, 924, 1057]. working [199]. Workplace [1193]. works [758, 953, 928, 820, 612]. World [1299, 1421, 1164, 167]. Worlds [1252, 607, 732, 914, 1780, 648]. Worried [1136]. Worthy [948]. Wouldn’t [1425]. Write [1455]. Writing [1327]. Writings [775, 833, 646]. Written [360]. Wrong [1630, 1659]. Wrongdoing [1240]. Wyatt [1709].

X [175, 288, 1173, 471, 175, 1184]. X.5 [1191]. Xanthippe [251]. Xavier [39]. Xeme [208]. XI [170]. XIII [184]. xiv [1192]. XIX [215]. XL [440, 465]. Xli [500]. XLII [513]. XLIII [525]. Xth [208, 217]. XV [193]. XVI [194, 1533, 1191, 279]. XVII [202]. XVIII [207]. XX [216, 1590, 1181, 1182]. XXI [224]. XXII [234, 235, 1184]. XXIV [270]. XXIX [315]. XXV [275]. XXVI [276]. XXVII [282]. XXVIII [309]. XXX [320]. XXXI [327]. XXXII [338]. XXXIII [339]. XXXIX [423]. Xxxv [369]. XXXVII [396]. XXXVIII [406].

Y. [1154]. Yaouzis [1415]. Year [1741, 1077, 46]. Years [1746, 1257, 268]. Yemima [1199]. yield [471]. York [1590, 1551]. yttersta [26].

Zahlendiagramme [260]. Zap [1116]. Zeit [256, 295, 176]. Zolt [1764]. Zombie [1476]. zum [279, 280, 268]. Zur [37, 46, 161, 290, 121, 57, 69, 63, 286, 102, 211]. Zwei [290]. Zweifel [335]. zweiwertigen [261].

References

Westermarck:1935:MI

[1] Edvard Westermarck. Den moraliska intuitionen. (Swedish) [Moral intuition]. Theoria, 1(1–2):1–31, April 1935. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hagerstrom:1935:BVPa

[2] Axel Hägerström. Begreppet viljeförklaring på privaträttens område. (Swedish) [The concept of a declaration of intent on the private-rights area]. Theoria, 1(1–2):32–57, April 1935. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kalla:1935:AER

[3] Elno Kalla. Albert Einsteins religion. Theoria, 1(1–2):58–67, April 1935. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kuhr:1935:BTB

[4] Victor Kuhr. Bemærkninger til Belysning af et karakterologisk Grundbegreb. (Danish) [Comments on the illumination of a characterological basic concept]. Theoria, 1(1–2):68–71, April 1935. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1935:ASNa

[5] Anonymous. Anmälningar. (Swedish) [Notifications]. Theoria, 1(1–2): 72–82, April 1935. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Brandt:1935:SKO

[6] Frithiof Brandt. Søren Kierkegaard og Mozarts Don Juan. (Danish) [Søren Kierkegaard and Mozarts Don Juan]. Theoria, 1(1–2):83–120, April 1935. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hagerstrom:1935:BVPb

[7] Axel Hägerström. Begreppet viljeförklaring på privaträttens område. (Swedish) [The concept of a declaration of intent on the private-rights area]. Theoria, 1(1–2):121–138, April 1935. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Marc-Wogau:1935:MPS

[8] Konrad Marc-Wogau. Månillusionens problem. (Swedish) [The moonillusion problem]. Theoria, 1(1–2):139–182, April 1935. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1935:ASNb

[9] Anonymous. Anmälningar. (Swedish) [Notifications]. Theoria, 1(1–2): 183–188, April 1935. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1935:D

[10] Anonymous. Diskussionsinlagg. Theoria, 1(1–2):189–191, April 1935. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Rudberg:1935:SP

[11] Gunnar Rudberg. Sokrateslarjungen Platon. Theoria, 1(3):193–202, December 1935. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ahlberg:1935:SOP

[12] Alf Ahlberg. De socials och politiska myterna. (Swedish) [Social and political myths]. Theoria, 1(3):203–230, December 1935. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sundbom:1935:POV

[13] Iver Sundbom. Prisbildning och värde : Ett filosofiskt problem i den teoretiska nationalekonomien. (Swedish) [Price formation and value: a philosophical problem in the theoretical national economy]. Theoria, 1(3): 231–241, December 1935. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1935:ASNc

[14] Anonymous. Anmälningar. (Swedish) [Notifications]. Theoria, 1(3):242– 259, December 1935. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1935:DDT

[15] Anonymous. Diskussionsinlägg: Docenterna Torgny T. Segerstedts och Gunnar Oxenstiernas slutinlägg i diskussionen om Docent Oxenstiernas arbete: Om betingelserna för det analytiska konsekvensbegreppets upprätthållande. (Swedish) [Discussion posts: Docent Torgny T. Segerstedt’s and Gunnar Oxenstierna’s endpoints in the discussion of Docent Oxenstierna’s work: On the conditions for the maintenance of analytical consequences]. Theoria, 1(3):260–263, December 1935. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kinberg:1936:HOM

[16] Olof Kinberg. Hjärnfunktion och moralisk funktion. (Swedish) [Brain function and moral function]. Theoria, 2(1–2):3–26, April 1936. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tegen:1936:KOG

[17] Einar Tegen. Kritisk objektivism : En grundståndpunkt och en kritik. (Swedish) [Critical objectivity: A fundamental point of view and a criticism]. Theoria, 2(1–2):27–57, April 1936. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Lagerborg:1936:OSS

[18] Rolf Lagerborg. Om sinnesjämvikt. (Swedish) [On equanimity]. Theoria, 2(1–2):58–71, April 1936. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Neurath:1936:LEO

[19] Otto Neurath. Den logiska empirismen och wienerkretsen. (Swedish) [Logical empiricism and the Viennese Circle]. Theoria, 2(1–2):72–82, April 1936. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1936:ASNa

[20] Anonymous. Anmälningar. (Swedish) [Notifications]. Theoria, 2(1–2): 83–96, April 1936. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1936:DSDa

[21] Anonymous. Diskussionsinlägg. (Swedish) [Discussion posts]. Theoria, 2 (1–2):97–105, April 1936. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1936:MSMa

[22] Anonymous. Meddelanden. (Swedish) [Messaging]. Theoria, 2(1–2):106, April 1936. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Russell:1936:EGS

[23] Bertrand Russell. Empirismens gränser. (Swedish) [The limits of empiricism]. Theoria, 2(1–2):107–127, April 1936. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kaila:1936:FSK

[24] Eino Kaila. Det främmande själslivets kunskapsteoretiska problem. (Swedish) [The theoretical problem of the foreign soul]. Theoria, 2(1– 2):128–155, April 1936. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Aspelin:1936:NPI

[25] Qunnar Aspelin. Nutida problem inotn kunskapsläran. (Swedish) [Current problems in the field of learning]. Theoria, 2(1–2):156–172, April 1936. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Vannerus:1936:LOO

[26] Allen Vannérus. Likhet och olikhet hos det yttersta : i anknytning till några aktuella naturvetenskapliga grundbegrepp. (Swedish) [Similarity and inequality in the utmost: in connection with the current science of basic concepts]. Theoria, 2(1–2):173–200, April 1936. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1936:ASNb

[27] Anonymous. Anmälningar. (Swedish) [Notifications]. Theoria, 2(1–2): 201–206, April 1936. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1936:DSDb

[28] Anonymous. Diskussionsinlägg. (Swedish) [Discussion posts]. Theoria, 2 (1–2):207–237, April 1936. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1936:MSMb

[29] Anonymous. Meddelanden. (Swedish) [Messaging]. Theoria, 2(1–2):238, April 1936. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kaila:1936:ASN

[30] Eino Kaila. Anmälningar. (Swedish) [Notifications]. Theoria, 2(3):233– 234, December 1936. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1936:DSDc

[31] Anonymous. Diskussionsinlägg. (Swedish) [Discussion posts]. Theoria, 2(3):235–364, December 1936. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ranulf:1936:P

[32] Svend Ranulf. Propaganda. Theoria, 2(3):239–256, December 1936. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Holm:1936:IOR

[33] Sverre Holm. Idealisme og Realisme i nyere fransk Filosofi. (Danish) [Idealism and realism in new French philosophy]. Theoria, 2(3):257–278, December 1936. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Marc-Wogau:1936:SPE

[34] Konrad Marc-Wogau. Der Symbolbegriff in der Philosophic Ernst Cassirers. (German) [The concept of symbols in the philosophical Ernst Cassirers]. Theoria, 2(3):279–332, December 1936. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1937:T

[35] Anonymous. Theoria 1937. Theoria, 3(1):1–2, April 1937. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hoffmann:1937:PPM

[36] Ernst Hoffmann. Über die Problematik der philosophiegeschichtlichen Methode. (German) [On the problem of the philosophical–historical method]. Theoria, 3(1):3–37, April 1937. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Saarnio:1937:HP

[37] Uuno Saarnio. Zur heterologischen paradoxie. Theoria, 3(1):38–56, April 1937. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Janet:1937:LI

[38] Pierre Janet. Le langage inconsistant. Theoria, 3(1):57–71, April 1937. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansen:1937:XL

[39] Valdemar Hansen. Xavier Léon. Theoria, 3(1):72–75, April 1937. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1937:WVP

[40] Anonymous. Was versteht der Physiker unter der “Grösse” eines Körpers?. (German) [What does the physicist understand by the “size” of a body?]. Theoria, 3(1):76–89, April 1937. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Frank:1937:WVP

[41] Philipp Frank. Was versteht der Physiker unter der “Grösse” eines Körpers?: Bemerkungen zu A. Phaléns Kritik der Einsteinschen Relativitätstheorie. (German) [What does the physicist understand by the “size” of a body?: Remarks on A. Phalén’s critique of Einstein’s relativity theory]. Theoria, 3(1):90–114, April 1937. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1937:Da

[42] Anonymous. Discussions. Theoria, 3(1):115–127, April 1937. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1937:Ra

[43] Anonymous. Reviews. Theoria, 3(1):128–156, April 1937. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1937:Na

[44] Anonymous. Notices. Theoria, 3(1):157, April 1937. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1937:GJ

[45] Anonymous. G. A. Jaederholm. Theoria, 3(1):158–159, April 1937. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Cassirer:1937:DWB

[46] Ernst Cassirer. Descartes’ Wahrheitsbegriff: Betrachtungen zur 300-Jahresfeler des “Discours de la Méthode”. (German) [Descartes’ concept of truth: Reflections on the 300-year-old field of the “Discourse on the Method”]. Theoria, 3(2–3):161–187, August 1937. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hedenberg:1937:VVG

[47] Sven Hedenberg. Vincent van gogh. sa maladie et son art. Theoria, 3(2– 3):188–205, August 1937. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hempel:1937:PVF

[48] C. G. Hempel. Le problème de la vérité. (French) [The problem of truth]. Theoria, 3(2–3):206–244, August 1937. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1937:BGB

[49] Anonymous. Bibliographie. (German) [Bibliography]. Theoria, 3(2–3): 245–246, August 1937. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Honigswald:1937:PWS

[50] Richard Honigswald. Über Philosophie als Wissenschaft von der Sprache: Prolegomena zu einer Problemkritik der Sprachphilosophie. (German)

[On philosophy as science from language: Prolegomena to a problemcritique of language philosophy]. Theoria, 3(2–3):247–266, August 1937. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Elmgren:1937:QNP

[51] John Elmgren. Quelques notions sur la psychologie de la haine. (French) [Some notions about the psychology of hatred]. Theoria, 3(2–3):267–282, August 1937. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1937:Db

[52] Anonymous. Discussions. Theoria, 3(2–3):283–342, August 1937. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1937:Rb

[53] Anonymous. Reviews. Theoria, 3(2–3):343–378, August 1937. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1937:Nb

[54] Anonymous. Notices. Theoria, 3(2–3):379, August 1937. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1938:AH

[55] Anonymous. Axel Herrlin. Theoria, 4(1):1–2, April 1938. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

vonWright:1938:WME

[56] G. H. von Wright. Der Wahrscheinlichkeitsbegriff in der modernen Erkenntnisphilosophie. (German) [The probability concept in modern epistemology]. Theoria, 4(1):3–20, April 1938. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Oxenstierna:1938:BVG

[57] Gunnar Oxenstierna. Bemerkungen über das Verhältnis der Geometrie zur Anschauung. (German) [Remarks on the relation of geometry to intuition]. Theoria, 4(1):21–38, April 1938. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ness:1938:CST

[58] Arne Ness. Common-sense and truth. Theoria, 4(1):39–58, April 1938. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1938:Da

[59] Anonymous. Discussions. Theoria, 4(1):59–84, April 1938. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1938:Ra

[60] Anonymous. Reviews. Theoria, 4(1):85–96, April 1938. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Brehier:1938:CHP

[61] Émile Bréhier. La causalité en histoire de la philosophie. (French) [Causality in the history of philosophy]. Theoria, 4(2):97–116, August 1938. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1938:OBS

[62] Anonymous. Objektivität und Begriffsbildung in den Sozialwissenschaften. (German) [Objectivity and conceptualization in social sciences]. Theoria, 4(2):117–144, August 1938. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Cassrier:1938:LSG

[63] Ernst Cassrier. Zur Logik des Symbolbegriffs. (German) [The logic of the symbolic concept]. Theoria, 4(2):145–175, August 1938. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansen:1938:GLW

[64] Valdemar Hansen. Un grand livre sur William James. (French) [A great book on William James]. Theoria, 4(2):176–180, August 1938. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1938:E

[65] Anonymous. Ethics. Theoria, 4(2):181–190, August 1938. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Aspelin:1938:R

[66] Gunnar Aspelin. Review. Theoria, 4(2):191–192, August 1938. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1938:N

[67] Anonymous. Notis. Theoria, 4(2):192, August 1938. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Cassirer:1938:BAD

[68] Ernst Cassirer. Über Bedeutung und Abfassungszeit von Descartes’ “Recherche de la Vérité par la lumière naturelle” : Eine kritische Betrachtung. (German) [On the meaning and era of Descartes’ “Search for Truth through Natural Light”: Critical consideration]. Theoria, 4(3): 193–234, December 1938. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Wittenberg:1938:WDI

[69] Erich Wittenberg. Die Wissenschaftskrisis in Deutschland im Jahre 1919: Ein Beitrag zur Wissenschaftsgeschichte. (German) [The science crisis in Germany in 1919: A contribution to the history of science]. Theoria, 4(3): 235–264, December 1938. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1938:Db

[70] Anonymous. Discussions. Theoria, 4(3):265–296, December 1938. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1938:Rb

[71] Anonymous. Review. Theoria, 4(3):297–318, December 1938. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tegen:1939:HM

[72] Einar Tegen. Harry meurling. Theoria, 5(1):1, January 1939. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Rodhe:1939:CC

[73] Sven Edvard Rodhe. Correspondence and coherence. Theoria, 5(1):3–40, January 1939. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

DeVieeschauwer:1939:LHP

[74] Herman J. De Vieeschauwer. L’apriorisme dans l’histoire de la philosophie. (French) [Apriorism in the history of philosophy]. Theoria, 5(1): 41–60, January 1939. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kinberg:1939:MCP

[75] Olof Kinberg. De la morale comme phénomène social objectif. (French) [Morality as an objective social phenomenon]. Theoria, 5(1):61–80, January 1939. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ranulf:1939:E

[76] Svend Ranulf. Epistemolgy. Theoria, 5(1):81–86, January 1939. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

DeVleeschauwer:1939:HP

[77] H. J. De Vleeschauwer. History of philosophy. Theoria, 5(1):86–96, January 1939. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ledent:1939:E

[78] Adrien Ledent. Ethics. Theoria, 5(1):97–100, January 1939. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1939:Ra

[79] Anonymous. Reviews. Theoria, 5(1):101–107, January 1939. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1939:Na

[80] Anonymous. Notice. Theoria, 5(1):108–110, January 1939. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Casslrer:1939:WSG

[81] Ernst Casslrer. Was ist “Subjektivismus”?. (German) [What is subjectivism?]. Theoria, 5(2):111–140, January 1939. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Lonborg:1939:CPD

[82] Sven Lönborg. The chronology of the Platonic dialogues. Theoria, 5 (2):141–160, January 1939. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Aron:1939:ROS

[83] Raymond Aron. Remarques sur l’objectivitè des sciences sociales. (French) [Remarks on the objectivity of the social sciences]. Theoria, 5 (2):161–194, January 1939. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1939:E

[84] Anonymous. Epistemology. Theoria, 5(2):195–210, January 1939. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1939:Rb

[85] Anonymous. Reviews. Theoria, 5(2):211–226, January 1939. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1939:Nb

[86] Anonymous. Notice. Theoria, 5(2):227, January 1939. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tegen:1939:AH

[87] Einar Tegen. Axel Hägerström 6/9 1868–7/7 1939. Theoria, 5(3):229– 232, January 1939. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hedenius:1939:GO

[88] Ingemar Hedenius. Gunnar oxenstierna. Theoria, 5(3):233–234, January 1939. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Groethuysen:1939:PHP

[89] Bernard Groethuysen. Les paradoxes de l’histoire de la philosophie. (French) [The paradoxes of the history of philosophy]. Theoria, 5(3): 235–264, January 1939. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Petzall:1939:CPS

[90] Åke Petzäll. Contribution to the problem of the sentiment of guilt. Theoria, 5(3):265–280, January 1939. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hedenius:1939:BKI

[91] Ingemar Hedenius. Begriffsanalyse und kritischer Idealismus (I). (German) [Conceptual analysis and critical idealism (I)]. Theoria, 5(3):281– 313, January 1939. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hedenius:1939:E

[92] Ingemar Hedenius, Adrien Ledent, Svend Ranulf, and Erich Wittenberg. Epistemology. Theoria, 5(3):314–338, January 1939. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Dupreel:1939:E

[93] E. Dupréel. Ethics. Theoria, 5(3):338–341, January 1939. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1939:Rc

[94] Anonymous. Reviews. Theoria, 5(3):342–354, January 1939. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1939:Nc

[95] Anonymous. Notice. Theoria, 5(3):355, January 1939. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Nyman:1940:PSP

[96] Alf Nyman. Problèmes et solutions en philosophie. (French) [Problems and solutions in philosophy]. Theoria, 6(1):1–42, April 1940. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ranulf:1940:TTS

[97] Svend Ranulf. Two types of sociology. Theoria, 6(1):43–57, April 1940. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kraft:1940:D

[98] Victor Kraft and Konrad Marc-Wogau. Discussions. Theoria, 6(1):58– 74, April 1940. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1940:Ra

[99] Anonymous. Revievs. Theoria, 6(1):75–107, April 1940. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1940:N

[100] Anonymous. Notice. Theoria, 6(1):108, April 1940. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Aspelin:1940:PFB

[101] Gunnar Aspelin. The polemics in the first book of Locke’s essay. Theoria, 6(2):109–122, August 1940. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Moritz:1940:LFG

[102] Manfred Moritz. Zur Logik der Frage. (German) [On the logic of the question]. Theoria, 6(2):123–149, August 1940. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

DeCoster:1940:D

[103] Sylvain De Coster. Discussions. Theoria, 6(2):150–163, August 1940. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1940:Rb

[104] Anonymous. Reviews. Theoria, 6(2):164–169, August 1940. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1940:BR

[105] Anonymous. Book received. Theoria, 6(2):170, August 1940. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

aakerman:1940:MIS

[106] Johan åkerman. The meaning of induction in social science. Theoria, 6(3):171–190, December 1940. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kraft:1940:MGM

[107] Victor Kraft. Über Moralbegründung. (German) [On moral establishment]. Theoria, 6(3):191–226, December 1940. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1940:D

[108] Anonymous. Discussions. Theoria, 6(3):227–237, December 1940. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1940:Rc

[109] Anonymous. Reviews. Theoria, 6(3):238–243, December 1940. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kinberg:1941:CEG

[110] O. Kinberg. The conception of environment in genetic bio-psychology. Theoria, 7(1):1–19, April 1941. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Marc-Wogau:1941:SBG

[111] Konrad. Marc-Wogau. Der Staat und der Begriff des Guten in Platons Politeia. (German) [The state and the concept of good in Plato’s Politeia]. Theoria, 7(1):20–45, April 1941. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1941:Da

[112] Anonymous. Discussions. Theoria, 7(1):46–71, April 1941. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Aspelin:1941:RAN

[113] Gunniar Aspelin. Reviews: Anders Nygren: Filosofi och motivforskning (Philosophy and Motive-Research). Svenska Kyrkans Diakonistyrelses Förlag. Stockholm 1940. Theoria, 7(1):72–74, April 1941. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Cassirer:1941:THG

[114] Ernst Cassirer. Thorild und Herder. (German) [Thorild and Herder]. Theoria, 7(2):75–92, August 1941. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Rodhe:1941:KIG

[115] Sven Edvard Rodhe. Über Kants Ideenlehre. (German) [On Kant’s doctrine of ideas]. Theoria, 7(2):93–124, August 1941. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1941:Db

[116] Anonymous. Discussions. Theoria, 7(2):125–153, August 1941. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1941:Ra

[117] Anonymous. Reviews. Theoria, 7(2):154–165, August 1941. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1941:BNa

[118] Anonymous. Bibliographical notes. Theoria, 7(2):166–170, August 1941. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ross:1941:ICD

[119] Alf Ross. On the illusion of consciousness: Do mental data (‘phenomena of consciousness’) exist? Theoria, 7(3):171–202, December 1941. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Widmark:1941:GMV

[120] Erik M. P. Widmark. Gehirnentwicklung und Menschenleiden : Vortrag in der Studentenvereinigung der Universität in Kopenhagen am 23. April 1938. (German) [Brain development and human suffering: lecture at the University of Copenhagen on 23rd April, 1938]. Theoria, 7(3):203–218, December 1941. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Moritz:1941:GPU

[121] Manfred Moritz. Gebot und Pflicht : Eine Untersuchung; zur imperativen Ethik. (German) [Commandment and duty: An inquiry; on imperative ethics]. Theoria, 7(3):219–257, December 1941. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1941:Dc

[122] Anonymous. Discussions. Theoria, 7(3):258–267, December 1941. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1941:Rb

[123] Anonymous. Review. Theoria, 7(3):268–277, December 1941. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1941:BNb

[124] Anonymous. Bibliographical notes. Theoria, 7(3):278–279, December 1941. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Aspelin:1942:EL

[125] Gunnar Aspelin. Efraim liljeqvist. Theoria, 8(1):1–2, April 1942. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Segerstedt:1942:CC

[126] Torgny T. Segerstedt. Customs and codes. (I). Theoria, 8(1):3–22, April 1942. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Pfannenstill:1942:MOS

[127] Bertil Pfannenstill. Method and object in sociology. Theoria, 8(1):23– 57, April 1942. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Marc-Wogau:1942:D

[128] Konrad Marc-Wogau. Discussions. Theoria, 8(1):58–71, April 1942. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1942:Ra

[129] Anonymous. Reviews. Theoria, 8(1):72–75, April 1942. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1942:BNB

[130] Anonymous. Bibliographical notes: Bibliographical notes. Theoria, 8(1): 76–83, April 1942. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kaila:1942:PPG

[131] Eino Kaila. Physikalismus und Phänomenalismus. (German) [Physicalism and phenomenalism]. Theoria, 8(2):85–125, August 1942. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Segerstedt:1942:CCI

[132] Torgny T. Segerstedt. Customs and codes. (II). Theoria, 8(2):126–153, August 1942. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kraft:1942:D

[133] Victor Kraft. Discussions. Theoria, 8(2):154–171, August 1942. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1942:Rb

[134] Anonymous. Reviews. Theoria, 8(2):172–196, August 1942. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1942:BNa

[135] Anonymous. Bibliographical notes. Theoria, 8(2):197–200, August 1942. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

vonWright:1942:GCL

[136] G. H. von Wright. Georg Christoph Lichtenberg als Philosoph. (German) [Georg Christoph Lichtenberg as philosopher]. Theoria, 8(3):201–217, December 1942. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Wittenberg:1942:GPB

[137] Erich Wittenberg. Grundvoraussetzungen für eine philosophische Begründung der Ethik als Wissenschaft. (German) [Basic requirements for a philosophical reason of ethics as science]. Theoria, 8(3):218–261, December 1942. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1942:D

[138] Anonymous. Discussions. Theoria, 8(3):262–285, December 1942. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1942:Rc

[139] Anonymous. Reviews. Theoria, 8(3):286–309, December 1942. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1942:BNc

[140] Anonymous. Bibliographical notes. Theoria, 8(3):310–313, December 1942. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Szekely:1943:BSD

[141] L. Szekely. Die Bedeutung der Situation fur das Denken : Ansätze zu einer Feldtheorie des Denkens. (German) [The meaning of the situation for thinking: Approaches to a field theory of thinking]. Theoria, 9(1): 1–21, April 1943. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Leander:1943:AWG

[142] Folke Leander. Analyse des Wirklichkeitsbegriffs (I). (German) [Analysis of the concept of reality (I)]. Theoria, 9(1):22–38, April 1943. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1943:Da

[143] Anonymous. Discussions. Theoria, 9(1):39–67, April 1943. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1943:Ra

[144] Anonymous. Reviews. Theoria, 9(1):68–83, April 1943. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1943:BNVa

[145] Anonymous. Bibliographical notes (VI). Theoria, 9(1):84–89, April 1943. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Marc-Wogau:1943:TVC

[146] Konrad Marc-Wogau. Die Theorie der Veränderung bei C. D. Broad. (German) [The theory of change in C. D. Broad]. Theoria, 9(2):91–123, August 1943. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Leander:1943:AWI

[147] Folke Leander. Analyse des Wirklichkeitsbegriffs (II). (German) [Analysis of the concept of reality (II)]. Theoria, 9(2):124–144, August 1943. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1943:Db

[148] Anonymous. Discussions. Theoria, 9(2):145–173, August 1943. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1943:Rb

[149] Anonymous. Reviews. Theoria, 9(2):174–183, August 1943. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1943:BNVb

[150] Anonymous. Bibliographical notes (VII). Theoria, 9(2):184–187, August 1943. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Moritz:1943:GIE

[151] Manfred Moritz. Ein Grundmotiv der idealistischen Ethik. (German) [A basic motif of idealistic ethics]. Theoria, 9(3):189–209, December 1943. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1943:Dc

[152] Anonymous. Discussions. Theoria, 9(3):213–219, December 1943. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1943:Rc

[153] Anonymous. Reviews. Theoria, 9(3):220–233, December 1943. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1943:BN

[154] Anonymous. Bibliographical notes. Theoria, 9(3):234–238, December 1943. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Nyman:1944:HL

[155] Alf Nyman. Hans Larsson. Theoria, 10(1):1–6, April 1944. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Nyman:1944:SAP

[156] Alf Nyman. Sur le système d’axiomes de la psychologie. (French) [On the system of axioms in psychology]. Theoria, 10(1):7–27, April 1944. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tegen:1944:BPT

[157] Efnar Tegen. The basic problem in the theory of value. Theoria, 10(1): 28–52, April 1944. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1944:D

[158] Anonymous. Discussions. Theoria, 10(1):53–57, April 1944. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1944:Ra

[159] Anonymous. Reviews. Theoria, 10(1):58–82, April 1944. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1944:BNa

[160] Anonymous. Bibliographical notes. Theoria, 10(1):83–90, April 1944. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kaila:1944:LPB

[161] Eino Kaila. Logik und Psychophysik: Ein Beitrag zur theoretischen Psychologie. (German) [Logic and psychophysics: A contribution to theoretical psychology]. Theoria, 10(2):91–119, August 1944. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hedenius:1944:UUG

[162] Ingemar Hedenius. Überzeugung und Urteil. (German) [Conviction and judgment]. Theoria, 10(2):120–170, August 1944. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Moritz:1944:D

[163] Manfred Moritz. Discussions. Theoria, 10(2):171–175, August 1944. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1944:Rb

[164] Anonymous. Reviews. Theoria, 10(2):176–189, August 1944. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1944:BNb

[165] Anonymous. Bibliographical notes. Theoria, 10(2):190–196, August 1944. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

From:1944:MP

[166] Franz From. Mandeville’s paradox. Theoria, 10(3):197–215, December 1944. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Wedberg:1944:LCW

[167] Anders Wedberg. The logical construction of the world: A critical analysis of Rudolf Carnap’s Der logische Aufbau der Welt. Theoria, 10(3): 216–246, December 1944. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sandstrom:1944:D

[168] Tora Sandström. Discussions. Theoria, 10(3):247–249, December 1944. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1944:Rc

[169] Anonymous. Reviews. Theoria, 10(3):250–257, December 1944. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1944:BNX

[170] Anonymous. Bibliographical notes: (XI).: (may 1st to August 31st, 1944). Theoria, 10(3):258–263, December 1944. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1944:BR

[171] Anonymous. Books received. Theoria, 10(3):264, December 1944. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Segerstedt:1945:IPJ

[172] Torgny T. Segerstedt. Imperative propositions and judgments of value. Theoria, 11(1):1–19, April 1945. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Morin:1945:PTG

[173] Harald Morin. Ein Problem der Teleologie. (German) [A problem of teleology]. Theoria, 11(1):20–39, April 1945. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Lovestad:1945:SPL

[174] Ludvig Lövestad. The structure of physical laws. Theoria, 11(1):40– 70, April 1945. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Fries:1945:WNH

[175] Martin Fries. Der Wirklichkeitsbegriff nach Hägerström (Verklighetsbegreppet enligt Hagerström). Antwort auf Anders Wedbergs Besprechung des Buches mit dem obigen Titel (Theoria X, 250–52). (German) [The concept of reality according to Hägerstrøm. Answer to Anders Wedberg’s review of the book with the above title (Theoria X, 250–252)]. Theoria, 11(1):71–74, April 1945. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Fries:1945:RHM

[176] Martin Fries. Reviews: Harald Morin: Wikner och tidsproblemet (Wikner und das Problem der Zeit). Theoria, 11(1):75–77, April 1945. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1945:BNBa

[177] Anonymous. Biblographical notes: Biblographical notes. Theoria, 11(1): 78–83, April 1945. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1945:BR

[178] Anonymous. Books received. Theoria, 11(1):84, April 1945. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Petzall:1945:ECE

[179] Ake Petzall. Ernst cassier: Ernst cassier. Theoria, 11(2):85–87, August 1945. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kaila:1945:WLL

[180] Eino Kaila. Wenn … so …. (German) [If … then …]. Theoria, 11 (2):88–98, August 1945. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Reenpaa:1945:WDG

[181] Yrjö Reenpää. Wahrnehmen und Denken. (German) [Perception and thought]. Theoria, 11(2):99–125, August 1945. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1945:D

Anonymous:1945:Ra

[183] Anonymous. Reviews. Theoria, 11(2):137–148, August 1945. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1945:BNX

[184] Anonymous. Bibliographical notes (XIII): January 1st to April 30th, 1945. Theoria, 11(2):149–154, August 1945. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Rodhe:1945:KSP

[185] Sven Edvard Rodhe. Kant in der schwedischen Philosophie. (German) [Kant in Swedish philosophy]. Theoria, 11(3):155–171, December 1945. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ross:1945:LNP

[186] Alf Ross. On the logical nature of propositions of value. Theoria, 11(3): 172–210, December 1945. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1945:Rb

[187] Anonymous. Reviews. Theoria, 11(3):211–221, December 1945. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1945:BNBb

[188] Anonymous. Bibliographical notes: Bibliographical notes. Theoria, 11(3):222–226, December 1945. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hagerstrom:1946:EVS

[189] Axel Hägerström. Erkenntnistheoretische Vorausseteungen der speziellen Relativitätstheorie Einsteins. (German) [Theoretical predictions of Einstein’s theory of special relativity]. Theoria, 12(1–2):1–68, April 1946. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Schultzer:1946:CVP

[190] Bent Schultzer. Collectivism viewed in a psychological light. Theoria, 12(1–2):69–90, April 1946. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1946:Da

[191] Anonymous. Discussions. Theoria, 12(1–2):91–109, April 1946. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1946:R

[192] Anonymous. Reviews. Theoria, 12(1–2):110–123, April 1946. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1946:BNXa

[193] Anonymous. Bibliographical notes (XV): September 1st to December 31st, 1945. Theoria, 12(1–2):124–129, April 1946. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1946:BNXb

[194] Anonymous. Bibliographical notes (XVI): January 1st to April 30th, 1946. Theoria, 12(1–2):130–134, April 1946. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1946:BR

[195] Anonymous. Books received. Theoria, 12(1–2):135–136, April 1946. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1946:AV

[196] Anonymous. Allen Vannérus. Theoria, 12(3):137–142, December 1946. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Broad:1946:LLC

[197] C. D. Broad. Leibniz’s last controversy with the Newtonians. Theoria, 12(3):143–168, December 1946. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kinberg:1946:CFP

[198] Olof Kinberg. On the concept of ^{≪}psychopathy^{≫}and the treatment of so called^{≫}psychopaths^{≫}. Theoria, 12(3):169–180, December 1946. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Schmidt:1946:LWH

[199] Folke Schmidt, Leif Gräntze, and Axel Roos. Legal working hours in Swedish agriculture: A summary of a field study. Theoria, 12(3):181–196, December 1946. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1946:Db

vAster:1946:R

[201] E. v.Aster. Reviews. Theoria, 12(3):211–213, December 1946. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1946:BNXc

[202] Anonymous. Bibliographical notes (XVII). Theoria, 12(3):214–217, December 1946. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Petzall:1947:SFP

[203] Âke Petzäll. The social function of punishment. Theoria, 13(1):1–46, January 1947. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sjobring:1947:SOA

[204] H. Sjöbring. Subjective and objective attitude. Theoria, 13(1):47–64, January 1947. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1947:D

[205] Anonymous. Discussions. Theoria, 13(1):65–74, January 1947. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1947:BR

[206] Anonymous. Books received. Theoria, 13(1):74, January 1947. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1947:BNXa

[207] Anonymous. Bibliographical notes (XVIII). Theoria, 13(1):75–80, January 1947. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1947:XCI

[208] Anonymous. Xeme Congrés international de philosophie. (French) [Xth International Congress of Philosophy]. Theoria, 13(1):81–83, January 1947. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kruse:1947:MSS

[209] Víndíng Kruse. The method of social sciences. Theoria, 13(2–3):85–135, August 1947. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Stenius:1947:NIM

[210] Erik Stenius. Natural implication and material implication. Theoria, 13 (2–3):136–156, August 1947. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Szekely:1947:PIV

[211] Lajos Székely. Zur Psychologie des inneren Verhaltens beim Lernen, Denken und Erfahren. (German) [The psychology of inner behavior in learning, thinking and experiencing]. Theoria, 13(2–3):157–182, August 1947. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

MacLeod:1947:WTA

[212] Andries H. D. MacLeod. What is a true assertion? Theoria, 13(2– 3):183–214, August 1947. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ranulf:1947:ICS

[213] Svend Ranulf. IIIe Congrès des Sociétés de Philosophie de Langue Française. (French) [Third Congress of Philosophy Societies for French Language]. Theoria, 13(2–3):215–217, August 1947. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1947:R

[214] Anonymous. Reviews. Theoria, 13(2–3):218–224, August 1947. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1947:BNXb

[215] Anonymous. Bibliographical notes (XIX): (January 1st to April 30th, 1947). Theoria, 13(2–3):225–229, August 1947. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1947:BNXc

[216] Anonymous. Bibliographical notes (XX): (May 1st to August 31st, 1947). Theoria, 13(2–3):230–232, August 1947. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1947:XIC

[217] Anonymous. Xth International Congress of Philosophy: Amsterdam. Theoria, 13(2–3):233–235, August 1947. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Brehier:1948:NPP

[218] Emile Bréhier. La notion de problème en philosophie. (French) [The notion of problems in philosophy]. Theoria, 14(1):1–7, April 1948. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Devaux:1948:SAS

[219] Philippe Devaux. Some aspects of Samuel Alexander’s philosophy. Theoria, 14(1):8–27, April 1948. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hagerstrom:1948:LCS

[220] Axel Hägerström. Lectures on so-called spiritual religion. Theoria, 14 (1):28–67, April 1948. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Saarnio:1948:BGG

[221] Uuno Saarnio. Der Begriff des Guten. (German) [The concept of good]. Theoria, 14(1):68–83, April 1948. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1948:D

[222] Anonymous. Discussions. Theoria, 14(1):84–99, April 1948. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Diderichsen:1948:R

[223] Paul Diderichsen. Reviews. Theoria, 14(1):101–103, April 1948. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1948:BNXa

[224] Anonymous. Bibliographical notes XXI. Theoria, 14(1):104–110, April 1948. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Petzall:1948:RP

[225] Åke Petzäll. “The right to punish”. Theoria, 14(2):113–115, August 1948. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

DeGreeff:1948:DPF

[226] Etienne De Greeff. Le droit de punir. (French) [The right to punish]. Theoria, 14(2):116–117, August 1948. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ewing:1948:PRP

[227] A. C. Ewing. The primary reason for punishment. Theoria, 14(2):118– 123, August 1948. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kinberg:1948:DPF

[228] Olof Kinberg. Le droit de punir. (French) [The right to punish]. Theoria, 14(2):124–143, August 1948. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Mannheim:1948:PCR

[229] Hermann Mannheim. Problems of collective responsibility. Theoria, 14 (2):144–166, August 1948. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Olivecrona:1948:SEL

[230] Karl Olivecrona. Is a sociological explanation of law possible? Theoria, 14(2):167–207, August 1948. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kinberg:1948:MCW

[231] Olof Kinberg. Motive, choice, will. Theoria, 14(3):209–237, December 1948. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Rodhe:1948:EUM

[232] Sven Edvard Rodhe. Is existence a univocal or a multivocal expression? Theoria, 14(3):238–264, December 1948. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hallden:1948:QCL

[233] Sören Halldén. A question concerning a logical calculus related to Lewis’ system of strict implication, which is of special interest for the study of entailment. Theoria, 14(3):265–269, December 1948. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1948:BNXb

[234] Anonymous. Bibliographical notes (XXII): (January 1st to April 30th, 1948). Theoria, 14(3):270–275, December 1948. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1948:BNXc

[235] Anonymous. Bibliographical notes (XXII): (May 1st to August 31st, 1948). Theoria, 14(3):276–278, December 1948. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ahlberg:1949:EMM

[236] Alf Ahlberg. Eschatologische Motive des Marxismus. (German) [Eschatological motives of Marxism]. Theoria, 15(1–3):1–16, April 1949. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Almberg:1949:BFB

[237] Nils Almberg. Zu dem Begriff ^{≪}Breite der Gesundheit^{≪}bei Galen. (German) [On the term “breadth of health’ in Galen]. Theoria, 15(1–3):17–28, April 1949. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Aspelin:1949:LS

[238] Gunnar Aspelin. Locke and Sydenham. Theoria, 15(1–3):29–37, April 1949. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Brandt:1949:GES

[239] Frithiof Brandt. The great earthquake in Søren Kierkegaard’s life. Theoria, 15(1–3):38–53, April 1949. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Broad:1949:LPN

[240] C. D. Broad. Leibniz’s predicate-in-notion principle and some of its alleged consequences. Theoria, 15(1–3):54–70, April 1949. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Destouches:1949:NDN

[241] J. L. Destouches. Notions descriptives et notions construites en physique modern. Theoria, 15(1–3):71–77, April 1949. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Destouches-Fevrier:1949:MSM

[242] Paulette Destouches-Février. Monde sensible et monde atomique. Theoria, 15(1–3):78–89, April 1949. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hartshorne:1949:SIR

[243] Charles Hartshorne. The synthesis of idealism and realism. Theoria, 15 (1–3):90–107, April 1949. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hedenius:1949:VD

[244] Ingemar Hedenius. Values and duties. Theoria, 15(1–3):108–115, April 1949. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Jorgensen:1949:RCC

[245] Jørgen Jørgensen. Remarks concerning the concept of mind and the problem of other people’s minds. Theoria, 15(1–3):116–127, April 1949. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Karitz:1949:SSF

[246] Anders Karitz. Das Symbol und der seelische Funktionszusammenhang: Ein Resümee. (German) [The symbol and the psychological functional relationship: A résumé]. Theoria, 15(1–3):128–140, April 1949. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Katz:1949:BKH

[247] David Katz. Blendungsphänomene und konnektive Hemmung bei Denkprozessen. (German) [Glare phenomena and connective inhibition in thought processes]. Theoria, 15(1–3):141–154, April 1949. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Levi:1949:CPA

[248] Alessandro Levi. Comment peut-on analyser la conscience juridique?. (French) [How can one analyze the legal consciousness?]. Theoria, 15 (1–3):155–163, April 1949. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Lundholm:1949:CPT

[249] Helge Lundholm. Commentary on the physicalistic trend in contemporary psychological science with special reference to the United States. Theoria, 15(1–3):164–179, April 1949. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Lundmark:1949:DPS

[250] Knut Lundmark. Do planetary systems exist? Theoria, 15(1–3):180– 197, April 1949. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Lonborg:1949:SX

[251] Sven Lönborg. Socrates and Xanthippe. Theoria, 15(1–3):198–204, April 1949. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Morin:1949:EGB

[252] Harald Morin. Einige Gesichtspunkte betreffend Natur und Geist in Swedenborgs De commersio animae et corporis. (German) [Some aspects concerning nature and spirit in Swedenborg’s De commersio animae et corporis]. Theoria, 15(1–3):205–219, April 1949. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Naess:1949:TTI

[253] Arne Næss. Toward a theory of interpretation and preciseness. Theoria, 15(1–3):220–241, April 1949. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Petzall:1949:IBL

[254] Åke Petzäll. The individual as the basis of law. Theoria, 15(1–3):242– 252, April 1949. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ruin:1949:PSA

[255] Hans Ruin. La psychologie structurale et l’art moderne. (French) [Structural psychology and modern art]. Theoria, 15(1–3):253–275, April 1949. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Salomaa:1949:ZEG

[256] J. E. Salomaa. Die Zeit als Element der Geschichte. (German) [Time as an element of history]. Theoria, 15(1–3):276–297, April 1949. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Schultzer:1949:ELP

[257] Bent Schultzer. Empiricism as a logical problem. Theoria, 15(1–3):298– 314, April 1949. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Segelberg:1949:SNP

[258] Ivar Segelberg. Some notes on perception. Theoria, 15(1–3):315–322, April 1949. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Segerstedt:1949:RGS

[259] Torgny T. Segerstedt. A research into the general sense of justice. Theoria, 15(1–3):323–338, April 1949. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Siegvald:1949:ZEG

[260] Herman Siegvald. Zahlendiagramme bei Erwachsenen. (German) [Number charts for adults]. Theoria, 15(1–3):339–366, April 1949. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Valen-Sendstad:1949:WZL

[261] Olav Valen-Sendstad. Der Wahrheitsbegriff in der zweiwertigen Logik. (German) [The concept of truth in divalent logic]. Theoria, 15(1–3):367– 383, April 1949. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Wermlund:1949:WMT

[262] Sven Wermlund. William McDougall’s theory of primitive sympathy. Theoria, 15(1–3):384–398, April 1949. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Starfelt:1949:BCL

[263] Eric Starfelt. Bibliographie: Contribution à l‘établissement d’une bibliographie des écrits imprimés de M. le Professeur Alf Nyman de 1899 à 1949. (French) [Bibliography: Contribution to the establishment of a bibliography of M. Professor Alf Nyman from 1899 to 1949]. Theoria, 15 (1–3):399–423, April 1949. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

VonAster:1950:PPP

[264] Ernst Von Aster. Physikalistischer und psychologistischer Positivismus. (German) [Physicalistic and psychological positivism]. Theoria, 16(1):1– 20, April 1950. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bergmann:1950:PEG

[265] Hugo Bergmann. Probleme des Existenzbegriffes. (German) [Problems of the concept of existence]. Theoria, 16(1):21–35, April 1950. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Geiger:1950:SRS

[266] Theodor Geiger. Some reflections on sociometry and its limitations. Theoria, 16(1):36–48, April 1950. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Husen:1950:PSM

[267] Torsten Husén. Problems of scaling in modern psychometrics. Theoria, 16(1):49–62, April 1950. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Marc-Wogau:1950:BBF

[268] Konrad Marc-Wogau. Bemerkungen zum Begriff ^{≪}Sinnesdatum^{≫} in der Diskussion der letzten Jahre. (German) [Remarks on the term “sense datum” in the discussion of the last years]. Theoria, 16(1):63–74, April 1950. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic). See remarks [279].

Anonymous:1950:Da

[269] Anonymous. Discussions. Theoria, 16(1):75–81, April 1950. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1950:BNXa

[270] Anonymous. Bibliographical notes (XXIV). Theoria, 16(1):82–89, April 1950. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bar-Hillel:1950:BDA

[271] Yehoshoua Bar-Hillel. Bolzano’s definition of analytic propositions. Theoria, 16(2):91–117, August 1950. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ofstad:1950:DDCa

[272] Harald Ofstad. The descriptive definition of the concept ‘legal norm’ proposed by Hans kelsen: An elementary analytical and critical investigation. Theoria, 16(2):118–151, August 1950. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1950:Db

[273] Anonymous. Discussion. Theoria, 16(2):152–153, August 1950. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1950:Ra

[274] Anonymous. Reviews. Theoria, 16(2):154–164, August 1950. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1950:BNXb

[275] Anonymous. Bibliographical notes (XXV): (January 1st to December 31st, 1949). Theoria, 16(2):165–178, August 1950. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1950:BNXc

[276] Anonymous. Bibliographical notes (XXVI): (January 1st to June 30th 1950). Theoria, 16(2):179–184, August 1950. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kinberg:1950:SPP

[277] Olof Kinberg. Les situations psychologiques précriminelles révélatrices des caractères de l‘état dangereux. (French) [Pre-sentence psychological situations revealing the characteristics of the dangerous state]. Theoria, 16(3):185–210, December 1950. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ofstad:1950:DDCb

[278] Harald Ofstad. The descriptive definition of the concept ‘legal norm’ proposed by Hans kelsen: An elementary analytical and critical investigation. Theoria, 16(3):211–246, December 1950. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hartnack:1950:NER

[279] Justus Hartnack. A note on existence. Remarks on Professor Konrad Marc-Wogau’s ^{≪}Bemerkungen zum Begriff ‘Sinnesdatum’ in der Diskussion der letzten Jahre^{≪}(Theoria XVI). Theoria, 16(3):247–248, December 1950. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic). See [268, 280].

Marc-Wogau:1950:BAF

[280] Konrad Marc-Wogau. Bemerkung zum Ausdruck ^{≪}existiert^{≪}. Antwort auf die Bemerkungen von Justus Hartnack. (German) [Remark on the expression “exists”. Answer to the comments of Justus Hartnack]. Theoria, 16(3):249–250, December 1950. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic). See [279].

Anonymous:1950:Rb

[281] Anonymous. Reviews. Theoria, 16(3):251–269, December 1950. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1950:BNXd

[282] Anonymous. Bibliographical notes (XXVII): (July 1st to December 31st, 1950). Theoria, 16(3):270–281, December 1950. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Boalt:1951:FNS

[283] Gunnar Boalt. Family name and social class. Theoria, 17(1–3):1–12, April 1951. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Broad:1951:LDS

[284] C. D. Broad. Locke’s doctrine of substantial identity & diversity. Theoria, 17(1–3):13–26, April 1951. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Dahlstrom:1951:SAQ

[285] Edmund Dahlström. Some aspects of quantitative verification in sociology. Theoria, 17(1–3):27–38, April 1951. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Fries:1951:IEG

[286] Martin Fries. Zur Interpretation einiger Gedanken in der Philosophie Hägerströms. (German) [On the interpretation of some thoughts in Hägerström’s philosophy]. Theoria, 17(1–3):39–45, April 1951. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hallden:1951:WW

[287] Sören Halldén. What is a word? Theoria, 17(1–3):46–56, April 1951. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hedenius:1951:NPR

[288] Ingemar Hedenius. A note to Plato’s Republic X, 596 c–d. Theoria, 17 (1–3):57–60, April 1951. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Husen:1951:ISU

[289] Torsten Husen. The influence of schooling upon IQ. Theoria, 17(1–3):61– 88, April 1951. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Katz:1951:ZBP

[290] David Katz. Zwei Beiträge zur Psychologie der Wahrnehmung. (German) [Two articles on the psychology of perception]. Theoria, 17(1–3):89– 102, April 1951. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kinberg:1951:PFC

[291] Olof Kinberg. La prison facteur criminogène. (French) [The prison criminogenic factor]. Theoria, 17(1–3):103–122, April 1951. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Lagerborg:1951:LRW

[292] Rolf Lagerborg. L’ethical relativity de Westermarck. (French) [The ethical relativity of Westermarck]. Theoria, 17(1–3):123–127, April 1951. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Leod:1951:IKN

[293] Andries H. D. Mac Leod. Implikation und Konsequenz nach Gunnar Oxenstierna. (German) [Implication and consequence according to Gunnar Oxenstierna]. Theoria, 17(1–3):128–139, April 1951. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Marc-Wogau:1951:KLA

[294] Konrad Marc-Wogau. Kants Lehre vom analytischen Urteil. (German) [Kant’s doctrine of analytical judgment]. Theoria, 17(1–3):140–154, April 1951. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

vonMorin:1951:WZS

[295] Harald von Morin. Wille und Zeit in Schopenhauers Philosophie. (German) [Will and time in Schopenhauer’s philosophy]. Theoria, 17(1–3): 155–175, April 1951. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Moritz:1951:SHP

[296] Manfred Moritz. Über den Satz die Handlung H ist Pflicht. (German) [On the sentence the action H is duty]. Theoria, 17(1–3):176–182, April 1951. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Petzall:1951:MS

[297] Åke Petzäll. Man and society. Theoria, 17(1–3):183–192, April 1951. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Pfannenstill:1951:CAO

[298] Bertil Pfannenstill. A critical analysis of operational definitions. Theoria, 17(1–3):193–209, April 1951. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Regnell:1951:NNF

[299] Hans Regnell. On negation and negative facts. Theoria, 17(1–3):210– 221, April 1951. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Segelberg:1951:IG

[300] Ivar Segelberg. The intentionality of gladness. Theoria, 17(1–3):222– 225, April 1951. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Segerstedt:1951:SAA

[301] Torgny T. Segerstedt. Some assumptions in attitude research. Theoria, 17(1–3):226–239, April 1951. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Szekely:1951:RAF

[302] Lajos Szekely. Die Realität in der Auffassung Freuds. (German) [Reality in Freud’s view]. Theoria, 17(1–3):240–245, April 1951. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Wedberg:1951:SPL

[303] Anders Wedberg. Some problems in the logical analysis of legal science. Theoria, 17(1–3):246–275, April 1951. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Wermlund:1951:CBP

[304] Sven Wermlund. Culture, behavior, and patterns of behavior. Theoria, 17 (1–3):276–290, April 1951. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hallden:1952:KKO

[305] Sören Halldén. Kants Kritik des ontologischen Gottesbeweises. (German) [Kant’s critique of ontological proof of God]. Theoria, 18(1–2):1–31, April 1952. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Johansen:1952:RTD

[306] Holger Johansen. Die Russellsche Theorie der definiten Deskriptionen vom Standpunkt der Sprachwissenschaft aus betrachtet. (German) [Russell’s theory of definite descriptions from the viewpoint of linguistic science]. Theoria, 18(1–2):32–58, April 1952. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Gotlind:1952:SRH

[307] Erik Götlind. Some remarks on Halldén’s paper ^{≪}book review: What is a word?^{≪}. Theoria, 18(1–2):59–77, April 1952. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1952:R

[308] Anonymous. Reviews. Theoria, 18(1–2):78–86, April 1952. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1952:BNXa

[309] Anonymous. Bibliographical notes (XXVIII). Theoria, 18(1–2):87–110, April 1952. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1952:BR

[310] Anonymous. Books received. Theoria, 18(1–2):111–112, April 1952. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Altavilla:1952:VLT

[311] Enrico Altavilla. Valeur et limites d’une théorie monogénésique du crime fondée sur les troubles affectifs. (French) [The value and limitations of a monogenic theory of crime based on emotional disorders]. Theoria, 18(3): 113–138, December 1952. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Carlsson:1952:SPC

[312] GÖSta Carlsson. Sampling, probability and causal inference. Theoria, 18(3):139–154, December 1952. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Rotenstreich:1952:CPS

[313] Nathan Rotenstreich. Cassirer’s philosophy of symbolic forms and the problem of history. Theoria, 18(3):155–173, December 1952. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Prior:1952:DMD

[314] A. N. Prior. Discussion: Modality de dicto and modality de re. Theoria, 18(3):174–180, December 1952. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1952:BNXb

[315] Anonymous. Bibliographical notes (XXIX): (January 1st to June 30th. 1952). Theoria, 18(3):181–189, December 1952. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ayer:1953:OKO

[316] A. J. Ayer. One’s knowledge of other minds. Theoria, 19(1–2):1–20, April 1953. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Nyman:1953:II

[317] Alf Nyman. Induction et intuition. Theoria, 19(1–2):21–41, April 1953. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kinberg:1953:RCP

[318] Olof Kinberg. Réflexions critiques sur la prévention soi-disant générale. (French) [Critical reflections on so-called general prevention]. Theoria, 19(1–2):42–62, April 1953. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1953:R

[319] Anonymous. Reviews. Theoria, 19(1–2):63–72, April 1953. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1953:BNXa

[320] Anonymous. Bibliographical notes (XXX): (July 1st to December 31st, 1952). Theoria, 19(1–2):73–88, April 1953. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Blanche:1953:OCF

[321] Robert Blanche. Sur l’opposition des concepts. (French) [On the opposition of concepts]. Theoria, 19(3):89–130, December 1953. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1953:VFS

[322] Anonymous. Verpflitchtung und Freiheit : Über den Satz ^{≪}sollen impliziert können^{≫}. (German) [Obligation and freedom: the implication of the sentence “may be implied”]. Theoria, 19(3):131–171, December 1953. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Christensen:1953:NCS

[323] Johnny Christensen. A note concerning the scholastic background of Leibniz’s philosophy. Theoria, 19(3):172–177, December 1953. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Gotlind:1953:NFM

[324] Erik Götlind. Note on a formula in my ^{≪}Bertrand Russell’s Theories of Causation^{≫}. Theoria, 19(3):177–178, December 1953. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Henry:1953:SAV

[325] D. P. Henry. St. Auselm on the varieties of ‘doing’. Theoria, 19(3): 178–183, December 1953. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1953:EUA

[326] Anonymous. Elementare Untersuchungen der Antinomien von Russell, Grelling-Nelson und Eubulides. Von Veli Valpola (Helsingfors). (German) [Elemental investigations of the antinomies of Russell, Grelling-Nelson, and Eubulides. From Veli Valpola (Helsinki)]. Theoria, 19(3):183–188, December 1953. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1953:BNXb

[327] Anonymous. Bibliographical notes (XXXI): (January 1st to June 30th, 1953). Theoria, 19(3):189–201, December 1953. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1953:BR

[328] Anonymous. Books received. Theoria, 19(3):202–205, December 1953. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Devaux:1954:PBP

[329] Philippe Devaux. La place de Berkeley dans la philosophie moderne. (French) [The place of Berkeley in modern philosophy]. Theoria, 20(1–3):

1–22, April 1954. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Pap:1954:PSS

[330] Arthur Pap. Propositions, sentences, and the semantic definition of truth. Theoria, 20(1–3):23–35, April 1954. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kinberg:1954:LFM

[331] Olof Kinberg. L‘étude de la fonction morale dans ses rapports avec l’état dangereux. (French) [The study of the moral function in its relations with the dangerous state]. Theoria, 20(1–3):36–52, April 1954. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Naess:1954:HAE

[332] Arne Næss. Husserl on the apodictic evidence of ideal laws. Theoria, 20(1–3):53–63, April 1954. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Petzall:1954:SAE

[333] Ake Petzäll. La syndérèse de l’aigle d’Ezéchiel à la conscience morale par le Commentaire de Saint Jérôme. (French) [The eagle’s syndication of Ezekiel to the moral conscience by the Commentary of St. Jerome]. Theoria, 20(1–3):64–77, April 1954. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Moritz:1954:PSJ

[334] Manfred Moritz. Der praktische Syllogismus and das juridische Denken. (German) [Practical syllogism and juridical thinking]. Theoria, 20(1–3): 78–127, April 1954. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Marc-Wogau:1954:ZDC

[335] Konrad Marc-Wogau. Der Zweifel Descartes’ and das Cogito ergo sum. (German) [The doubt of Descartes’ and the I think, therefore I am]. Theoria, 20(1–3):128–152, April 1954. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Stenius:1954:LSS

[336] Erik Stenius. Linguistic structure and the structure of experience. Theoria, 20(1–3):153–172, April 1954. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Rescher:1954:NSD

[337] Nicholas Rescher. A note on a species of definition. Theoria, 20(1–3):173– 180, April 1954. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1954:BNXa

[338] Anonymous. Bibliographical notes (XXXII). Theoria, 20(1–3):181–193, April 1954. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1954:BNXb

[339] Anonymous. Bibliographical notes (XXXIII). Theoria, 20(1–3):194–204, April 1954. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1954:BR

[340] Anonymous. Books received. Theoria, 20(1–3):205–206, April 1954. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Gotlind:1955:VVT

[341] Erik Götlind. Vacuous variants and truth by convention. Theoria, 21(1): 1–24, April 1955. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Regnell:1955:RS

[342] Hans Regnell. Regularity and space. Theoria, 21(1):25–41, April 1955. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tornebohm:1955:DA

[343] Hakan Törnebohm. Discourse analysis. Theoria, 21(1):42–54, April 1955. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1955:Da

[344] Anonymous. Discussion. Theoria, 21(1):55–62, April 1955. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1955:RWS

[345] Anonymous. Reviews: Wolfgang Stegmüller: Hauptströmungen der Gegenwartsphilosophie. Theoria, 21(1):63–64, April 1955. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Blom:1955:CCM

[346] Siri Blom. Concerning a controversy on the meaning of ‘ probability ’. Theoria, 21(2–3):65–98, August 1955. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kanger:1955:NPP

[347] Stig Kanger. A note on partial postulate sets for propositional logic. Theoria, 21(2–3):99–104, August 1955. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ofstad:1955:BOC

[348] Harald Ofstad. Broad on ought and can: A critical discussion of his essay: Determinism, indeterminism and libertarianism. Theoria, 21(2– 3):105–116, August 1955. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Prior:1955:BLF

[349] A. N. Prior. Berkeley in logical form. Theoria, 21(2–3):117–122, August 1955. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ryding:1955:NDM

[350] Erik Ryding. La notion du moi chez Condillac. (French) [The notion of the ego in Condillac]. Theoria, 21(2–3):123–130, August 1955. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Stenius:1955:DCV

[351] Erik Stenius. Definitions of the concept “value-judgement”. Theoria, 21 (2–3):131–145, August 1955. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tornebohm:1955:CI

[352] Håkan Törnebohm. Content of information. Theoria, 21(2–3):146–157, August 1955. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Wang:1955:NAS

[353] Hao Wang. Notes on the analytic-synthetic distinction. Theoria, 21(2– 3):158–178, August 1955. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1955:Db

[354] Anonymous. Discussion. Theoria, 21(2–3):179–181, August 1955. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1955:E

[355] Anonymous. Erratum. Theoria, 21(2–3):181, August 1955. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1955:NN

[356] Anonymous. Notes: NOTES. Theoria, 21(2–3):182–190, August 1955. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1955:RWE

[357] Anonymous. Reviews: William Emerson Ritter: Charles Darwin and the Golden Rule. Theoria, 21(2–3):191–195, August 1955. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1955:BN

[358] Anonymous. Bibliographical notes. Theoria, 21(2–3):196–209, August 1955. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1955:BR

[359] Anonymous. Books received. Theoria, 21(2–3):210–212, August 1955. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Apostel:1956:IWT

[360] Leo Apostel. The interpretation of written texts. Theoria, 22(1):1– 32, April 1956. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Nyman:1956:ARS

[361] Alf Nyman. Antonio Rosmini et la Scandinavie. (French) [Antonio Rosmini and Scandinavia]. Theoria, 22(1):33–42, April 1956. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Palmieri:1956:CV

[362] L. E. Palmieri. Comments on verification. Theoria, 22(1):43–48, April 1956. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Xenakis:1956:FMN

[363] Jason Xenakis. Function and meaning of names. Theoria, 22(1):49– 60, April 1956. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1956:R

[364] Anonymous. Reviews. Theoria, 22(1):61–71, April 1956. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1956:BRa

[365] Anonymous. Books received. Theoria, 22(1):72–74, April 1956. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Eklund:1956:LC

[366] Harald Eklund. On the logic of creeds. Theoria, 22(2):75–84, August 1956. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Khatchadourian:1956:FC

[367] Haig Khatchadourian. Frege on concepts. Theoria, 22(2):85–100, August 1956. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1956:RLG

[368] Anonymous. Reviews: Lars Gurmund. Theoria, 22(2):101–105, August 1956. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1956:SBX

[369] Anonymous. Scandinavian bibliography (xxxv): Scandinavian bibliography (xxxv). Theoria, 22(2):106–143, August 1956. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Issman:1956:SU

[370] S. Issman. Signification et usage. Theoria, 22(3):145–152, December 1956. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Mercier:1956:DPF

[371] André Mercier. Le dilemme de la philosophie. (French) [The dilemma of philosophy]. Theoria, 22(3):153–166, December 1956. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Michaud-Quantin:1956:DVA

[372] Pierre Michaud-Quantin. La double-vérité des Averroïstes: Un texte nouveau de Boèce de Dacie. (French) [The double truth of the Averroistes: A new text of Boethius of Dacia]. Theoria, 22(3):167–184, December 1956. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tornebohm:1956:TIT

[373] Håkan Törnebohm. On truth, implication, and three-valued logic. Theoria, 22(3):185–198, December 1956. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1956:DD

[374] Anonymous. Discussion: Discussion. Theoria, 22(3):199–203, December

1956. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1956:BRb

[375] Anonymous. Books received. Theoria, 22(3):204–205, December 1956.

CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kanger:1957:MSP

[376] Stig Kanger. The morning star paradox. Theoria, 23(1):1–11, April 1957.

CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Marc-Wogau:1957:BSE

[377] Konrad Marc-Wogau. Berkeley’s sensationalism and the Esse est percipiprinciple. Theoria, 23(1):12–36, April 1957. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tranoy:1957:IAH

[378] Knut Erik Tranöy. An important aspect of humanism. Theoria, 23(1):37– 52, April 1957. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1957:SB

[379] Anonymous. Scandinavian bibliography. Theoria, 23(1):53–67, April

1957. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Marc-Wogau:1957:AP

[380] Konrad Marc-Wogau. Åke Petzäll. Theoria, 23(2):69–70, August 1957.

CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Stromberg:1957:SRC

[381] Tore Strömberg. Some reflections on the concept of punishment. Theoria, 23(2):71–83, August 1957. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tornebohm:1957:TLS

[382] Håkan Törnebohm. On two logical systems proposed in the philosophy of quantum-mechanics. Theoria, 23(2):84–101, August 1957. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Weinberger:1957:NSG

[383] Ota Weinberger. Über die Negation von Sollsätzen. (German) [On the negation of setpoints]. Theoria, 23(2):102–132, August 1957. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1957:Da

[384] Anonymous. Discussion. Theoria, 23(2):133–134, August 1957. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1957:BR

[385] Anonymous. Book received. Theoria, 23(2):135–136, August 1957. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Eklund:1957:EAE

[386] Harald Eklund. Über Evidenz und Anschaulichkeit: Einige Reflexionen zu Hans Reichenbachs Fundamentalsätzen. (German) [On evidence and vividness: some reflections on Hans Reichenbach’s Fundamental Principles]. Theoria, 23(3):137–151, December 1957. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kanger:1957:CM

[387] Stig Kanger. On the characterization of modalities. Theoria, 23(3): 152–155, December 1957. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Nyman:1957:FFL

[388] Alf Nyman. Der Fiktionalismus und die Fiktionen in der Lyrik: Eine Übersicht und eine Auseinandersetzung über die Metaphern. (German) [Fictionalism and fictions in lyricism: A survey and discussion of metaphors]. Theoria, 23(3):156–179, December 1957. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Steenberg:1957:SA

[389] Elisa Steenberg. A study of aesthetics. Theoria, 23(3):180–192, December 1957. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1957:Db

[390] Anonymous. Discussions. Theoria, 23(3):193–197, December 1957. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Berg:1958:NRS

[391] Jan Berg. A note on reduction sentences. Theoria, 24(1):1–8, April 1958. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Nielsen:1958:GRE

[392] Kai Nielsen. Good reasons in ethics: An examination of the Toulmin– Hare controversy. Theoria, 24(1):9–28, April 1958. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Saarnio:1958:GIE

[393] Uuno Saarnio. Über die Gesamtheit und ihre Einteilung. (German) [On the whole and its division]. Theoria, 24(1):29–38, April 1958. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tornebohm:1958:OBT

[394] Håkan Törnebohm. Outlines of a Boolean tensor algebra with applications to the lower functional calculus. Theoria, 24(1):39–47, April 1958. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ryding:1958:DDB

[395] Erik Ryding. Discussion:the distinction between one-one relation and one-many relation. Theoria, 24(1):48–50, April 1958. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1958:SBXa

[396] Anonymous. Scandinavian bibliography (XXXVII). Theoria, 24(1):51– 69, April 1958. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bergman:1958:A

[397] Gustav Bergman. Analyticity. Theoria, 24(2):71–93, August 1958. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Marc-Wogau:1958:AIB

[398] Konrad Marc-Wogau. The argument from illusion and Berkeley’s idealism. Theoria, 24(2):94–106, August 1958. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Black:1958:NMRa

[399] Max Black. Notes on the meaning of ‘rule’. Theoria, 24(2):107–126, August 1958. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1958:Da

[400] Anonymous. Discussions. Theoria, 24(2):127–135, August 1958. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

vonWright:1958:EK

[401] Georg Henrik von Wright. Eino Kaila. Theoria, 24(3):137–138, December 1958. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Black:1958:NMRb

[402] Max Black. Notes on the meaning of ‘rule’. Theoria, 24(3):139–161, December 1958. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hartnack:1958:D

[403] Justus Hartnack. Dreams. Theoria, 24(3):162–171, December 1958. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Khatchadourian:1958:WR

[404] Haig Khatchadourian. What is rationality? Theoria, 24(3):172–187, December 1958. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1958:Db

[405] Anonymous. Discussions. Theoria, 24(3):188–191, December 1958. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1958:SBXb

[406] Anonymous. Scandinavian bibliography (XXXVIII). Theoria, 24(3): 192–203, December 1958. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Curry:1959:IFI

[407] Haskell B. Curry. The interpretation of formalized implication. Theoria, 25(1):1–26, April 1959. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hjorth:1959:MPS

[408] Sune Hjorth. The meanings of probability statements. Theoria, 25(1):27– 30, April 1959. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Naess:1959:DWK

[409] Arne Næss. Do we know that basic norms cannot be true or false? Theoria, 25(1):31–53, April 1959. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic). See critical comments [473].

Palmieri:1959:DPP

[410] L. E. Palmieri. Discussion: Past and present. some remarks on A. I. Ayer’s theory on past and present in his “The Problem of Knowledge”. Theoria, 25(1):54–55, April 1959. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Zinkernagel:1959:DGR

[411] Peter Zinkernagel. Discussion: General rules of language. Theoria, 25 (1):56–64, April 1959. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Fain:1959:MEP

[412] Haskell Fain. More on the esse is percipi principle. Theoria, 25(2):65–81, August 1959. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Svenonius:1959:NCF

[413] Lars Svenonius. No-categoricity in first-order predicate calculus. Theoria, 25(2):82–94, August 1959. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Turnbull:1959:ADL

[414] Robert G. Turnbull. Aseity and dependence in Leibniz’s metaphysics. Theoria, 25(2):95–114, August 1959. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Marc-Wogau:1959:DSR

[415] Konrad Marc-Wogau. Discussion: Some remarks on Haskell Fain’s paper “More on the esse is percipi principle”. Theoria, 25(2):115–117, August 1959. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hovstad:1959:RSC

[416] Johan Hovstad. Review: Stephen C. Pepper: The Sources of Value. Theoria, 25(2):118–121, August 1959. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Jorgensen:1959:TTI

[417] Jørgen Jørgensen. Towards a theory of inference. Theoria, 25(3):123–147, December 1959. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kinberg:1959:AAD

[418] Olof Kinberg. Abus alcoolique et délinquance: Étude fondée essentiellement sur des experéences suédoises. (French) [Alcohol abuse and delinquency: A study based essentially on Swedish experiments]. Theoria, 25(3):148–157, December 1959. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Phillips:1959:ILF

[419] Jean A. Phillips. An interpretation of logical formulas. Theoria, 25(3): 158–172, December 1959. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Svenonius:1959:TPM

[420] Lars Svenonius. A theorem on permutations in models. Theoria, 25(3): 173–178, December 1959. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Szekely:1959:DPE

[421] Lajos Székely. Discussion: The problem of experience in the gestalt psychology. Theoria, 25(3):179–186, December 1959. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Storheim:1959:RNA

[422] Eivind Storheim. Review: Naess, Arne: Wie färdert man heute die empirische Bewegung? Eine Auseinandersetzung mit dem Empirismus von Otto Neurath und Rudolph Carnap. (German) [How do you transform the empirical movement today? A Confrontation with the Empiricism of Otto Neurath and Rudolph Carnap]. Theoria, 25(3):187–191, December 1959. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1959:SBX

[423] Anonymous. Scandinavian bibliography (XXXIX): (January 1st to December 31st, 1958). Theoria, 25(3):192–208, December 1959. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1959:BR

[424] Anonymous. Books received. Theoria, 25(3):209–212, December 1959. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hedenius:1960:RL

[425] Ingemar Hedenius. Rolf Lagerborg. Theoria, 26(1):1–2, April 1960. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Berg:1960:SPC

[426] Jan Berg. Some problems concerning disposition concepts. Theoria, 26 (1):3–16, April 1960. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Grossmann:1960:DBN

[427] Reinhardt Grossmann. Digby and Berkeley on notions. Theoria, 26(1): 17–30, April 1960. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Shwayder:1960:ULU

[428] D. S. Shwayder. Uses of language and uses of words: With application to a problem of Frege. Theoria, 26(1):31–43, April 1960. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Svenonius:1960:MMF

[429] Lars Svenonius. On minimal models of first-order systems. Theoria, 26(1):44–52, April 1960. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tornebohm:1960:SHT

[430] Håkan Törnebohm. A study in Hobbes’ theory of denotation and truth. Theoria, 26(1):53–67, April 1960. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Smith:1960:IC

[431] P. H. Nowell Smith. Ifs and cans. Theoria, 26(2):85–101, August 1960. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Prawitz:1960:IPP

[432] Dag Prawitz. An improved proof procedure. Theoria, 26(2):102–139, August 1960. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Jorgensen:1960:BMA

[433] Jørgen Jørgensen and Harry M. Bracken. Berkeley and mental acts. Theoria, 26(2):140–146, August 1960. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Richman:1960:TA

[434] Robert J. Richman. On a type of “ambiguity”. Theoria, 26(2):146–150, August 1960. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Jorgensen:1960:E

[435] Jørgen Jørgensen. Erratum. Theoria, 26(2):150, August 1960. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Castaneda:1960:OTG

[436] Hector Neri Castañeda. Outline of a theory on the general logical structure of the language of action. Theoria, 26(3):151–182, December 1960. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Morris:1960:IO

[437] Herbert Morris. Imperatives and orders. Theoria, 26(3):183–209, December 1960. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ossowska:1960:ESM

[438] Marie Ossowska. Existe-t-il un système de morale qui mérite le nom de système?. (French) [Is there a system of morality that deserves the name system?]. Theoria, 26(3):210–223, December 1960. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1960:D

Anonymous:1960:SBX

[440] Anonymous. Scandinavian bibliography (XL). Theoria, 26(3):241–256, December 1960. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1960:BR

[441] Anonymous. Books received. Theoria, 26(3):257–258, December 1960. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Rescher:1961:CR

[442] Nicholas. Rescher. The concept of randomness. Theoria, 27(1):1–11, April 1961. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Zabeeh:1961:HML

[443] Farhang. Zabeeh. Hume on metaphysics and the limits of human knowledge. Theoria, 27(1):12–25, April 1961. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Zich:1961:JBL

[444] Otakar. Zich. Über den Jevonsschen Begriff des logischen Sinnes und verwandte Begriffe. (German) [On Jevons’ concept of logical sense and related terms]. Theoria, 27(1):26–46, April 1961. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1961:Da

[445] Anonymous. Discussion. Theoria, 27(1):47–48, April 1961. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Dawson:1961:SEP

[446] E. E. Dawson. Sense experience and physical objects. Theoria, 27(2):49– 57, August 1961. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Gotlind:1961:TVA

[447] Erik. Götlind. Two views about the function of models in empirical theories. Theoria, 27(2):58–69, August 1961. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Louch:1961:UML

[448] A. R. Louch. “The use” of moral language. Theoria, 27(2):70–87, August 1961. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1961:E

[449] Anonymous. Erratum. Theoria, 27(2):87, August 1961. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1961:Db

[450] Anonymous. Discussions. Theoria, 27(2):88–94, August 1961. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1961:BRa

[451] Anonymous. Books received. Theoria, 27(2):95–97, August 1961. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Berg:1961:EOP

[452] Jan. Berg. An examination of the ontological proof. Theoria, 27(3): 99–106, December 1961. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Fisher:1961:TVC

[453] Mark. Fisher. A three-valued calculus for deontic logic. Theoria, 27(3): 107–118, December 1961. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hintikka:1961:MQ

[454] Jaakko. Hintikka. Modality and quantification. Theoria, 27(3):119–128, December 1961. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Opalek:1961:LPT

[455] Kazimierz. Opalek. The Leon petrazycki theory of law. Theoria, 27(3): 129–150, December 1961. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tornebohm:1961:SML

[456] Håkan. Törnebohm. A study in modal logic. Theoria, 27(3):151–164, December 1961. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1961:Dc

Anonymous:1961:BRb

[458] Anonymous. Books received. Theoria, 27(3):188–189, December 1961. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bergmann:1962:GE

[459] Gustav Bergmann. Generality and existence. Theoria, 28(1):1–26, April 1962. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Copeland:1962:SIFa

[460] Arthur H. Copeland, Sr. Statistical induction and the foundations of probability (I). Theoria, 28(1):27–44, April 1962. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hillman:1962:QC

[461] Donald J. Hillman. On quality classes. Theoria, 28(1):45–52, April 1962. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Jorgensen:1962:ROH

[462] Carl Jørgensen. The relation is/ought: Hume’s problem. Theoria, 28(1): 53–69, April 1962. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bergstrom:1962:DCC

[463] Lars Bergström. Discussion: Comments on Castañeda’s semantics of prescriptive discourse. Theoria, 28(1):70–72, April 1962. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Castaneda:1962:DSP

[464] H. N. Castañeda. Discussion: The semantics of prescriptive discourse. (A reply to Lars Bergstrom). Theoria, 28(1):72–78, April 1962. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1962:SBX

[465] Anonymous. Scandinavian bibliography (XL): (January 1st to December 1st, 1959). Theoria, 28(1):79–84, April 1962. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1962:BRa

[466] Anonymous. Books received. Theoria, 28(1):85–86, April 1962. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Copeland:1962:SIFb

[467] Arthur H. Copeland. Statistical induction and the foundations of probability. Theoria, 28(2):87–109, August 1962. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Nielsen:1962:SG

[468] Kai Nielsen. On speaking of God. Theoria, 28(2):110–137, August 1962. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Aaqvist:1962:RLN

[469] Lennart Åqvist. Reflections on the logic of nonsense. Theoria, 28(2): 138–157, August 1962. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Llewelyn:1962:PAP

[470] J. E. Llewelyn. Presuppositions, assumptions and presumptions. Theoria, 28(2):158–172, August 1962. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ross:1962:DXP

[471] James F. Ross. Does ‘x is possible’ ever yield ‘x exists? Theoria, 28 (2):173–195, August 1962. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Fisher:1962:SWN

[472] Mark Fisher. Strong, and weak negation of imperatives. Theoria, 28 (2):196–200, August 1962. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Osterberg:1962:WKN

[473] Dag Østerberg. We know that norms cannot be true or false. Critical comments on Arne Næss: Do we know that basic norms cannot be true or false? Theoria, 28(2):200–209, August 1962. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic). See [409].

Anonymous:1962:NN

[474] Anonymous. Note: NOTE. Theoria, 28(2):209, August 1962. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1962:BRb

[475] Anonymous. Books received. Theoria, 28(2):210–211, August 1962. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Wogau:1962:HE

[476] K. Marc Wogau. On historical explanation. Theoria, 28(3):213–233, December 1962. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tranoy:1962:HEC

[477] Knut Erik Tranøy. Historical explanation: causes and conditions. Theoria, 28(3):234–249, December 1962. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Schock:1962:DES

[478] Rolf Schock. A definition of event and some of its applications. Theoria, 28(3):250–268, December 1962. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Suszko:1962:NCB

[479] Roman Suszko. A note concerning the binary quantifiers. Theoria, 28(3): 269–276, December 1962. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

VonWright:1962:P

[480] Georg Henrik Von Wright. On promises. Theoria, 28(3):277–297, December 1962. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Blackstone:1962:LSM

[481] William T. Blackstone. On the logical status of meta-ethical theories. Theoria, 28(3):298–303, December 1962. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Margolis:1962:FPS

[482] Joseph Margolis. Fourteen points on the senses and their objects. Theoria, 28(3):303–308, December 1962. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Mosher:1962:RMW

[483] David Mosher. A reply to Mr. Wheatley. Theoria, 28(3):308–312, December 1962. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Wheatley:1962:TV

[484] Jon Wheatley. Tomatoes and vegetables. Theoria, 28(3):312–315, December 1962. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1962:BRc

[485] Anonymous. Books received. Theoria, 28(3):316, December 1962. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1962:EE

[486] Anonymous. Errata: ERRATA. Theoria, 28(3):316, December 1962. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Armour:1963:RGI

[487] Leslie Armour. Rationality, goodness and immortality. Theoria, 29(1): 1–11, April 1963. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

King-Farlow:1963:TAP

[488] John King-Farlow. Toulmin’s analysis of probability. Theoria, 29(1):12– 26, April 1963. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kyburg:1963:PR

[489] Henry E. Kyburg, Jr. Probability and randomness. Theoria, 29(1):27– 55, April 1963. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Lemos:1963:IIS

[490] Ramon M. Lemos. Ideas, images, and sensations. Theoria, 29(1):56– 69, April 1963. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sluga:1963:SRD

[491] H. D. Sluga. Some remarks on deontics. Theoria, 29(1):70–78, April 1963. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tornebohm:1963:CPN

[492] Håkan Törnebohm. The clock paradox and the notion of clock retardation in the special theory of relativity. Theoria, 29(1):79–90, April 1963. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1963:BR

[493] Anonymous. Books received. Theoria, 29(1):91–92, April 1963. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Golind:1963:SMS

[494] Erik Gölind. Stimulus meaning: Some problems in connection with a definition by Quine. Theoria, 29(2):93–114, August 1963. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hedenius:1963:P

[495] Ingemar Hedenius. Performatives. Theoria, 29(2):115–136, August 1963. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hartnack:1963:PUS

[496] Justus Hartnack. The performatory use of sentences. Theoria, 29(2): 137–146, August 1963. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tornebohm:1963:TCS

[497] Håkan Törnebohm. Two concepts of simultaneity in the special theory of relativity. Theoria, 29(2):147–153, August 1963. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

aaQvist:1963:PSD

[498] Lennart åQvist. Postulate sets and decision procedures for some systems of deontic logic. Theoria, 29(2):154–175, August 1963. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1963:DD

[499] Anonymous. Discussion: Discussion. Theoria, 29(2):176–204, August 1963. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1963:SBX

[500] Anonymous. Scandinavian bibliography (xli): Scandinavian bibliography (xli). Theoria, 29(2):205–228, August 1963. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Allaire:1963:BI

[501] Edwin B. Allaire. Berkeley’s idealism. Theoria, 29(3):229–244, December 1963. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Harre:1963:CI

[502] R. Harré. Counter-induction. Theoria, 29(3):245–264, December 1963. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Houlgate:1963:CRT

[503] Laurence D. Houlgate. Causation, recipes and theory. Theoria, 29(3): 265–276, December 1963. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Mercier:1963:PPE

[504] André Mercier. Petits prolégomènes à une étude sur le temps. (French) [Small prolegomena to a study on time]. Theoria, 29(3):277–282, December 1963. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tornebohm:1963:CDL

[505] Håkan. Törnebohm. On the concepts of distance and length in the special theory of relativity. Theoria, 29(3):283–289, December 1963. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Zabeeh:1963:VH

[506] Farhang. Zabeeh. Vindication of Hume. Theoria, 29(3):290–303, December 1963. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1963:D

[507] Anonymous. Discussions. Theoria, 29(3):304–315, December 1963. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Johnstone:1964:AIC

[508] Henry W. Johnstone and Robert Price. Axioms for the implicational calculus with one variable. Theoria, 30(1):1–4, April 1964. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Richman:1964:WSP

[509] Robert Richman. Why are synthetic a priori judgments necessary? Theoria, 30(1):5–20, April 1964. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Thompson:1964:PMV

[510] John W. Thompson. Polarity and the measurement of values. Theoria, 30(1):21–30, April 1964. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Gustafson:1964:DHT

[511] Don F. Gustafson. Discussions: Hampshire on trying. Theoria, 30(1):31– 38, April 1964. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hare:1964:R

[512] R. M. Hare and Lars Bergström. Review. Theoria, 30(1):39–49, April 1964. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1964:SBXa

[513] Anonymous. Scandinavian bibliography (XLII): (January 1st to December 31st, 1961). Theoria, 30(1):50–67, April 1964. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1964:BRa

[514] Anonymous. Books received. Theoria, 30(1):68–71, April 1964. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hartnack:1964:MMB

[515] Justus Hartnack. Movement, moments and beginning. Theoria, 30(2): 73–79, August 1964. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sinisi:1964:KTG

[516] Vito F. Sinisi. Kotarbiński’s theory of genuine names. Theoria, 30(2):80– 95, August 1964. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Strawson:1964:IRT

[517] P. F. Strawson. Identifying reference and truth-values. Theoria, 30(2): 96–118, August 1964. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Wheatley:1964:HGW

[518] Jon Wheatley. How to give a word a meaning. Theoria, 30(2):119–136, August 1964. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1964:DAQ

[519] Anonymous. Discussions: Ansätze zu einer Quantenlogik. (German) [Approaches to quantum logic]. Theoria, 30(2):137–143, August 1964. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Cornman:1964:SP

[520] James W. Cornman. Strawson’s “person”. Theoria, 30(3):145–156, December 1964. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Davis:1964:KEP

[521] Steven Davis. ‘i know’ as an explicit performative. Theoria, 30(3):157– 165, December 1964. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Khatchadourian:1964:A

[522] Haig Khatchadourian. Ambiguity. Theoria, 30(3):166–182, December 1964. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Lindstrom:1964:MC

[523] Per Lindström. On model-completeness. Theoria, 30(3):183–196, December 1964. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1964:D

[524] Anonymous. Discussion. Theoria, 30(3):197–205, December 1964. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1964:SBXb

[525] Anonymous. Scandinavian bibliography (XLIII): (January 1st to December 31st, 1962). Theoria, 30(3):206–222, December 1964. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1964:BRb

[526] Anonymous. Books received. Theoria, 30(3):223–225, December 1964. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ammerman:1965:OKS

[527] Robert Ammerman. Our knowledge of substance according to Locke. Theoria, 31(1):1–8, April 1965. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Berger:1965:EL

[528] Fred R. Berger. Excuses and the law. Theoria, 31(1):9–19, April 1965. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Danielsson:1965:DP

[529] Sven Danielsson. Definitions of ‘performative’. Theoria, 31(1):20–31, April 1965. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sobel:1965:GA

[530] J. Howard Sobel. Generalization arguments. Theoria, 31(1):32–60, April 1965. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Nelson:1965:EST

[531] John O. Nelson. An examination of Sommers’ truth–functional counterfactuals. Theoria, 31(1):61–63, April 1965. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1965:BR

[532] Anonymous. Books received. Theoria, 31(1):64, April 1965. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Richman:1965:CC

[533] Robert J. Richman. Concepts without criteria. Theoria, 31(2):65–85, August 1965. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ringbom:1965:PI

[534] Sixten Ringbom. Plato on images. Theoria, 31(2):86–109, August 1965. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Todd:1965:CLA

[535] William Todd. Causal laws and accidents. Theoria, 31(2):110–124, August 1965. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kerner:1965:D

[536] George C. Kerner. Discussions. Theoria, 31(2):125–132, August 1965. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Margolis:1965:BPS

[537] Joseph Margolis. Brain processes and sensations. Theoria, 31(2):133–138, August 1965. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Odegard:1965:EML

[538] Douglas Odegard. Excluding the middle from loose concepts. Theoria, 31 (2):138–144, August 1965. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Chihara:1965:WDM

[539] Charles S. Chihara. What dreams are made on. Theoria, 31(3):145–158, December 1965. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Grimm:1965:CES

[540] Robert Grimm. Cogito, ergo sum. Theoria, 31(3):159–173, December 1965. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Henze:1965:BSQ

[541] Donald F. Henze. Berkeley on sensations and qualities. Theoria, 31(3): 174–180, December 1965. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Mullane:1965:UE

[542] Harvey Mullane. Unconscious emotion. Theoria, 31(3):181–190, December 1965. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Rosenfield:1965:DMA

[543] Lawrence W. Rosenfield. The doctrine of the mean in Aristotle’s rhetoric. Theoria, 31(3):191–198, December 1965. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Segerberg:1965:CNL

[544] Krister Segerberg. A contribution to nonsense-logics. Theoria, 31(3): 199–217, December 1965. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sinisi:1965:KTP

[545] Vito F. Sinisi. Kotarbiński’s theory of pseudo-names. Theoria, 31(3): 218–241, December 1965. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Thalberg:1965:SPR

[546] Irving Thalberg. The Socratic paradox and reasons for action. Theoria, 31(3):242–254, December 1965. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

VonWright:1965:PC

[547] Georg Henrik Von Wright. The paradoxes of confirmation. Theoria, 31(3):255–274, December 1965. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1965:Da

[548] Anonymous. Discussions. Theoria, 31(3):274–288, December 1965. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1965:Db

[549] Anonymous. Discussions. Theoria, 31(3):275–288, December 1965. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hintikka:1966:KOO

[550] Jaakko Hintikka. “Knowing oneself” and other problems in epistemic logic. Theoria, 32(1):1–13, April 1966. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Lehman:1966:FEE

[551] Hugh Lehman. On the form of explanation in evolutionary theory. Theoria, 32(1):14–24, April 1966. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Luce:1966:CB

[552] David Randall Luce. A calculus of ‘before’. Theoria, 32(1):25–44, April 1966. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Newman:1966:TAP

[553] Fred Newman. Two analyses of prediction^{,}. Theoria, 32(1):45–55, April 1966. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Stangenberg:1966:NDS

[554] Christer Stangenberg. New definitions of scale-types. Theoria, 32(1):56– 61, April 1966. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Wheatley:1966:DC

[555] Jon Wheatley. Definition via categories. Theoria, 32(1):62–66, April 1966. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1966:Da

[556] Anonymous. Discussions. Theoria, 32(1):67–78, April 1966. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1966:BRa

[557] Anonymous. Books received. Theoria, 32(1):79–83, April 1966. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1966:E

[558] Anonymous. Erratum. Theoria, 32(1):83, April 1966. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kanger:1966:RP

[559] Stig Kanger and Helle Kanger. Rights and parliamentarism. Theoria, 32 (2):85–115, August 1966. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tillman:1966:FET

[560] Frank A. Tillman. Facts, events, and true statements. Theoria, 32(2): 116–129, August 1966. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Wheatley:1966:TTL

[561] Jon Wheatley. Two theories of language. Theoria, 32(2):130–143, August 1966. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1966:Db

Lindstrom:1966:C

[563] Per Lindström. On characterizability in L_{ω1ω0}. Theoria, 32(3):165–171, December 1966. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Lindstrom:1966:RBS

[564] Per Lindström. On relations between structures. Theoria, 32(3):172–185, December 1966. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1966:FOP

[565] Anonymous. First order predicate logic with generalized quantifiers. Theoria, 32(3):186–195, December 1966. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Maccallum:1966:AR

[566] Gerald C. Maccallum, Jr. On applying rules. Theoria, 32(3):196–210, December 1966. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sosa:1966:PIL

[567] Ernest Sosa. On practical inference and the logic of imperatives. Theoria, 32(3):211–223, December 1966. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sosa:1966:LI

[568] Ernest Sosa. The logic of imperatives. Theoria, 32(3):224–235, December 1966. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1966:Dc

[569] Anonymous. Discussion. Theoria, 32(3):237–251, December 1966. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1966:BRb

[570] Anonymous. Book received. Theoria, 32(3):252, December 1966. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Follesdal:1967:KIE

[571] Dagfinn Føllesdal. Knowledge, identity, and existence. Theoria, 33(1): 1–27, April 1967. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Prior:1967:SMT

[572] A. N. Prior. Stratified metric tense logic. Theoria, 33(1):28–38, April 1967. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Rescher:1967:NCL

[573] Nicholas Rescher and James W. Garson. A note on chronological logic. Theoria, 33(1):39–44, April 1967. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Segerberg:1967:LM

[574] Krister Segerberg. On the logic of “to-morrow”^{1}. Theoria, 33(1):45– 52, April 1967. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Segerberg:1967:SML

[575] Krister Segerberg. Some modal logics based on a three-valued logic. Theoria, 33(1):53–71, April 1967. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1967:Da

[576] Anonymous. Discussion. Theoria, 33(1):72–77, April 1967. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Binkley:1967:CAE

[577] Robert Binkley and Romane Clark. A cancellation algorithm for elementary logic. Theoria, 33(2):79–97, August 1967. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Greenlee:1967:LII

[578] Douglas Greenlee. Locke’s idea of “idea”. Theoria, 33(2):98–106, August 1967. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Wennerberg:1967:CFR

[579] Hjalmar Wennerberg. The concept of family resemblance in Wittgenstein’s later philosophy. Theoria, 33(2):107–132, August 1967. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1967:Db

Aspelin:1967:HKG

[581] Gunnar Aspelin. Historische Kausalsätze. (German) [Historical causal sentences]. Theoria, 33(3):157–175, December 1967. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Daniels:1967:IKA

[582] Charles B. Daniels. Immediate knowledge: Ayer, Strawson, and Shoemaker. Theoria, 33(3):176–188, December 1967. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Danielsson:1967:MLB

[583] Sven Danielsson. Modal logic based on probability theory. Theoria, 33(3): 189–197, December 1967. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kim:1967:PLT

[584] Jaegwon Kim. Psychophysical laws and theories of mind. Theoria, 33(3): 198–210, December 1967. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Mann:1967:DDO

[585] William E. Mann. Definite descriptions and the ontological argument. Theoria, 33(3):211–229, December 1967. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Odegard:1967:APP

[586] Douglas Odegard. Arthur pap and the paradox of analysis. Theoria, 33(3):230–245, December 1967. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Prawitz:1967:CHS

[587] Dag Prawitz. Completeness and hauptsatz for second order logic. Theoria, 33(3):246–258, December 1967. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Thalberg:1967:VDW

[588] Irving Thalberg. Verbs, deeds and what happens to us. Theoria, 33(3): 259–277, December 1967. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1967:Dc

[589] Anonymous. Discussions. Theoria, 33(3):278–289, December 1967. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1967:C

[590] Anonymous. Corrections. Theoria, 33(3):289, December 1967. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1967:BR

[591] Anonymous. Books received. Theoria, 33(3):290–292, December 1967. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kanger:1968:ET

[592] Stig Kanger. Equivalent theories. Theoria, 34(1):1–6, April 1968. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Segerberg:1968:DS

[593] Krister Segerberg. Decidability of S4.1. Theoria, 34(1):7–20, April 1968. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Segerberg:1968:DFM

[594] Krister Segerberg. Decidability of four modal logics. Theoria, 34(1):21– 25, April 1968. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Segerberg:1968:PLR

[595] Krister Segerberg. Propositional logics related to Heyting’s and Johansson’s. Theoria, 34(1):26–61, April 1968. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Wolfe:1968:CS

[596] Julian Wolfe. The criteria of sleep. Theoria, 34(1):62–65, April 1968. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Lyon:1968:FRV

[597] Ardon Lyon. Family resemblance, vagueness, and change of meaning. Theoria, 34(1):66–75, April 1968. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Prawitz:1968:DNU

[598] Dag Prawitz. A discussion note on utilitarianism. Theoria, 34(1):76– 84, April 1968. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Binkley:1968:CAC

[599] Robert Binkley and Romane Clark. A cancellation algorithm corrected. Theoria, 34(1):85, April 1968. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1968:BR

[600] Anonymous. Book received. Theoria, 34(1):86–87, April 1968. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Forguson:1968:RDB

[601] L. W. Forguson. On “it’s raining, but I don’t believe it”. Theoria, 34 (2):89–101, August 1968. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Margolis:1968:DE

[602] Joseph Margolis. “Directly evident”. Theoria, 34(2):102–116, August 1968. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Peczenik:1968:NR

[603] Aleksander Peczenik. Norms and reality. Theoria, 34(2):117–133, August 1968. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Prawitz:1968:P

[604] Dag Prawitz. Propositions. Theoria, 34(2):134–146, August 1968. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Snyder:1968:ME

[605] Aaron Snyder. Misusing expressions. Theoria, 34(2):147–162, August 1968. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1968:Da

[606] Anonymous. Discussion. Theoria, 34(2):163–170, August 1968. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bull:1968:PWP

[607] R. A. Bull. On possible worlds in propositional calculi. Theoria, 34(3): 171–182, December 1968. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Prior:1968:MLL

[608] A. N. Prior. Modal logic and the logic of applicability. Theoria, 34(3): 183–202, December 1968. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Stalnaker:1968:AFO

[609] Robert C. Stalnaker and Richmond H. Thomason. Abstraction in firstorder modal logic. Theoria, 34(3):203–207, December 1968. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

vonWright:1968:A

[610] G. H. von Wright. “Always”. Theoria, 34(3):208–221, December 1968. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Gram:1968:SSP

[611] Moltke S. Gram. Spinoza, substance, and predication. Theoria, 34(3): 222–244, December 1968. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Srzednicki:1968:JWA

[612] Jan Srzednicki. On judging works of art. Theoria, 34(3):245–255, December 1968. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1968:Db

[613] Anonymous. Discussions. Theoria, 34(3):256–258, December 1968. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Lindstrom:1969:EEL

[614] Per Lindström. On extensions of elementary logic. Theoria, 35(1):1– 11, April 1969. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Martin-Lof:1969:LMK

[615] Per Martin-Löf. The literature on von Mises’ Kollektivs Revisited. Theoria, 35(1):12–37, April 1969. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

OHair:1969:IM

[616] S. G. O’Hair. Implications and meaning. Theoria, 35(1):38–54, April 1969. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ryding:1969:BS

[617] Erik Ryding. Biased selection. Theoria, 35(1):55–63, April 1969. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1969:Da

[618] Anonymous. Discussions. Theoria, 35(1):64–79, April 1969. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Cocchiarella:1969:CTS

[619] Nino B. Cocchiarella. A completeness theorem in second order modal logic. Theoria, 35(2):81–103, August 1969. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Eberle:1969:DTP

[620] Rolf A. Eberle. Denotationless terms and predicates expressive of positive qualities. Theoria, 35(2):104–123, August 1969. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Lewis:1969:ESL

[621] Douglas Lewis. The existence of substances and Locke’s way of ideas. Theoria, 35(2):124–146, August 1969. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Odegard:1969:LUR

[622] Douglas Odegard. Locke and the unreality of relations. Theoria, 35 (2):147–152, August 1969. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Woolhouse:1969:SSL

[623] Roger Woolhouse. Substance and substances in Locke’s essay. Theoria, 35(2):153–167, August 1969. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Moline:1969:D

[624] Jon Moline. Discussion. Theoria, 35(2):168–182, August 1969. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1969:BR

[625] Anonymous. Books received. Theoria, 35(2):183–184, August 1969. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Glasgow:1969:AP

[626] W. D. Glasgow. The act of praising. Theoria, 35(3):185–203, December 1969. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Griffiths:1969:LAL

[627] Leslie Griffiths. The logic of Austin’s locutionary subdivision. Theoria, 35(3):204–214, December 1969. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Routley:1969:CET

[628] R. Routley and V. Routley. Categories — expressions or things? Theoria, 35(3):215–238, December 1969. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Resnik:1969:STA

[629] Michael D. Resnik. A set theoretic approach to the simple theory of types. Theoria, 35(3):239–258, December 1969. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Fitting:1969:LSM

[630] Melvin Fitting. Logics with several modal operators. Theoria, 35(3): 259–266, December 1969. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1969:Db

[631] Anonymous. Discussion. Theoria, 35(3):267–272, December 1969. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ruse:1970:RB

[632] Michael E. Ruse. The revolution in biology. Theoria, 36(1):1–22, April 1970. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Stalnaker:1970:SAC

[633] Robert C. Stalnaker and Richmond H. Thomason. A semantic analysis of conditional logic. Theoria, 36(1):23–42, April 1970. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Parsons:1970:AAC

[634] Charles Parsons. Axiomatization of Åqvist’s CS — logics. Theoria, 36 (1):43–64, April 1970. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Nilson:1970:LHE

[635] Sten Sparre Nilson. On the logic of historical explanation. Theoria, 36 (2):65–81, August 1970. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Martin:1970:CAC

[636] Raymond Martin. Causes and alternate causes. Theoria, 36(2):82–92, August 1970. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Dietl:1970:A

[637] Paul J. Dietl. Abnormalism. Theoria, 36(2):93–99, August 1970. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Mitchell:1970:MTV

[638] Dorothy Mitchell. Mill’s theory of value. Theoria, 36(2):100–115, August 1970. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Prawitz:1970:AA

[639] Dag Prawitz. The alternatives to an action. Theoria, 36(2):116–126, August 1970. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Srzednicki:1970:RD

[640] Jan Srzednicki. Reference and description. Theoria, 36(2):127–141, August 1970. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Stenius:1970:SAT

[641] Erik Stenius. Semantic antinomies and the theory of well-formed rules. Theoria, 36(2):142–160, August 1970. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Wolniewicz:1970:FNI

[642] Boguslaw Wolniewicz. Four notions of independence. Theoria, 36(2): 161–164, August 1970. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kung:1970:SQL

[643] Guido Küng and John Thomas Canty. Substitutional quantification and Leśniewskian quantifiers. Theoria, 36(2):165–182, August 1970. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Woodruff:1970:NJ

[644] Peter W. Woodruff. A note on JP’. Theoria, 36(2):183–184, August 1970. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Geach:1970:EAP

[645] P. T. Geach. Editorial: Arthur prior: A personal impression. Theoria, 36(3):185–188, December 1970. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Flo:1970:BPW

[646] Olav Flo. Bibliography of the philosophical writings of A. N. Prior. Theoria, 36(3):189–213, December 1970. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Stenius:1970:SSL

[647] Erik Stenius. The semantic status of laws of logic. Theoria, 36(3):214– 220, December 1970. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tranoy:1970:DLD

[648] K. E. Tranøy. Deontic logic and deontically perfect worlds. Theoria, 36(3):221–231, December 1970. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hintikka:1970:PFO

[649] Jaakko Hintikka. ‘prima facie’ obligations and iterated modalities. Theoria, 36(3):232–240, December 1970. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:1970:DLD

[650] Bengt Hansson. Deontic logic and different levels of generality. Theoria, 36(3):241–248, December 1970. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hamblin:1970:EWS

[651] C. L. Hamblin. The effect of when it’s said. Theoria, 36(3):249–263, December 1970. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Thomason:1970:ITT

[652] Richmond H. Thomason. Indeterminist time and truth-value gaps. Theoria, 36(3):264–281, December 1970. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bull:1970:ATL

[653] R. A. Bull. An approach to tense logic. Theoria, 36(3):282–300, December 1970. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Segerberg:1970:MLL

[654] Krister Segerberg. Modal logics with linear alternative relations. Theoria, 36(3):301–322, December 1970. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Gabbay:1970:SFM

[655] Dovm. Gabbay. Selective filtration in modal logic. Part A. Semantic tableaux method. Theoria, 36(3):323–330, December 1970. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Makinson:1970:GCR

[656] David Makinson. A generalisation of the concept of a relational model for modal logic. Theoria, 36(3):331–335, December 1970. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Fine:1970:PQM

[657] Kit Fine. Propositional quantifiers in modal logic. Theoria, 36(3):336– 346, December 1970. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Cresswell:1970:CIL

[658] M. J. Cresswell. Classical intensional logics. Theoria, 36(3):347–372, December 1970. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Montague:1970:UG

[659] Richard Montague. Universal grammar. Theoria, 36(3):373–398, December 1970. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Martin:1971:BMP

[660] Michael Martin. The body-mind problem and neurophysiological reduction. Theoria, 37(1):1–14, April 1971. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Marc-Wogau:1971:RNP

[661] K. Marc-Wogau. Remarks on Nilson’s paper “On the logic of historical explanation”. Theoria, 37(1):15–20, April 1971. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hochberg:1971:FCF

[662] Herbert Hochberg. Frege on concepts as functions: a fundamental ambiguity. Theoria, 37(1):21–32, April 1971. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Aqvist:1971:RFI

[663] Lennart Aqvist. Revised foundations for imperative-epistemic and interrogative logic. Theoria, 37(1):33–73, April 1971. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Lewis:1971:CDT

[664] David Lewis. Completeness and decidability of three logics of counterfactual conditionals. Theoria, 37(1):74–85, April 1971. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Germano:1971:ITD

[665] G. Germano. Incompleteness and truth definitions. Theoria, 37(1):86– 90, April 1971. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Cresswell:1971:EEP

[666] M. J. Cresswell. Essence and existence in Plato and Aristotle. Theoria, 37(2):91–113, August 1971. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Chellas:1971:I

[667] Brian F. Chellas. Imperatives. Theoria, 37(2):114–129, August 1971. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hamblin:1971:MMD

[668] C. L. Hamblin. Mathematical models of dialogue. Theoria, 37(2):130– 155, August 1971. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1971:Ra

[669] Anonymous. Reviews. Theoria, 37(2):156–183, August 1971. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1971:BR

[670] Anonymous. Books received. Theoria, 37(2):184, August 1971. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Malmgren:1971:MCI

[671] Helge Malmgren. Moore’s concept “indirect apprehension”. Theoria, 37(3):185–206, December 1971. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1971:AA

[672] Anonymous. Addendum: Addendum. Theoria, 37(3):207–208, December 1971. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tornebohm:1971:DDT

[673] Hakan Törnebohm. Durations and distances in time. Theoria, 37(3): 209–226, December 1971. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kamp:1971:MAP

[674] Hans Kamp. To the memory of Arthur prior formal properties of ‘now’. Theoria, 37(3):227–273, December 1971. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Urquhart:1971:CWI

[675] Alasdair I. F. Urquhart. Completeness of weak implication. Theoria, 37(3):274–282, December 1971. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1971:Rb

[676] Anonymous. Reviews. Theoria, 37(3):283–302, December 1971. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kanger:1972:MEP

[677] Stig Kanger. Measurement: An essay in philosophy of science. Theoria, 38(1–2):1–44, April 1972. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Suppes:1972:FEI

[678] Patrick Suppes. Finite equal-interval measurement structures. Theoria, 38(1–2):45–63, April 1972. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Domotor:1972:SMS

[679] Zoltan Domotor. Species of measurement structures. Theoria, 38(1–2): 64–81, April 1972. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Titiev:1972:MMU

[680] Robert J. Titiev. Multidimensional measurement and universal axiomatizability. Theoria, 38(1–2):82–88, April 1972. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1972:BR

[681] Anonymous. Books received. Theoria, 38(1–2):89, April 1972. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Gabbay:1972:GTC

[682] Dov M. Gabbay. A general theory of the conditional in terms of a ternary operator. Theoria, 38(3):97–104, December 1972. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kanger:1972:LL

[683] Stig Kanger. Law and logic. Theoria, 38(3):105–132, December 1972. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Fitting:1972:NCL

[684] Melvin Fitting. Non-classical logics and the independence results of set theory. Theoria, 38(3):133–142, December 1972. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Barnette:1972:CCN

[685] R. L. Barnette. Comment: Comments on neurophysiological reduction. Theoria, 38(3):143–144, December 1972. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bronaugh:1972:BUO

[686] Richard N. Bronaugh. Bergström’s utilitarian objection to T_{1}. Theoria, 38(3):145–147, December 1972. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1972:RBR

[687] Anonymous. Reply to bronaugh: Reply to bronaugh. Theoria, 38(3): 148–149, December 1972. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1972:BRB

[688] Anonymous. Books received: Books received. Theoria, 38(3):150, December 1972. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Danto:1973:ART

[689] Arthur C. Danto. Artworks and real things. Theoria, 39(1–3):1–17, April 1973. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sclafani:1973:AAT

[690] Richard J. Sclafani. Artworks, art theory, and the artworld. Theoria, 39(1–3):18–34, April 1973. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tormey:1973:CJ

[691] Alan Tormey. Critical judgments. Theoria, 39(1–3):35–49, April 1973. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Beardsley:1973:WAQ

[692] Monroe C. Beardsley. What is an aesthetic quality? Theoria, 39(1–3): 50–70, April 1973. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hermeren:1973:AQV

[693] Göran Hermerén. Aesthetic qualities, value and emotive meaning. Theoria, 39(1–3):71–100, April 1973. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Weitz:1973:IVA

[694] Morris Weitz. Interpretation and the visual arts. Theoria, 39(1–3):101– 112, April 1973. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Cohen:1973:ANA

[695] Ted Cohen. Aesthetic/non-aesthetic and the concept of taste: a critique of Sibley’s position. Theoria, 39(1–3):113–152, April 1973. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Dickie:1973:TAO

[696] George Dickie. Taste and attitude: the origin of the aesthetic. Theoria, 39 (1–3):153–170, April 1973. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Schaper:1973:RMA

[697] Eva Schaper. Review: Mimesis and art. Studies in the origin and early development of an aesthetic vocabulary. Theoria, 39(1–3):171–173, April 1973. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1973:BR

[698] Anonymous. Books received. Theoria, 39(1–3):174–175, April 1973. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Nilson:1974:TAR

[699] Sten Sparre Nilson. Two additional remarks on the logic of historical explanation. Theoria, 40(1):1–8, April 1974. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

VanInwagen:1974:FAP

[700] Peter Van Inwagen. A formal approach to the problem of free will and determinism. Theoria, 40(1):9–22, April 1974. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Fine:1974:ILC

[701] Kit Fine. An incomplete logic containing S4. Theoria, 40(1):23–29, April 1974. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Thomason:1974:ITM

[702] S. K. Thomason. An incompleteness theorem in modal logic. Theoria, 40(1):30–34, April 1974. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1974:R

[703] Anonymous. Reviews. Theoria, 40(1):35–61, April 1974. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Black:1974:T

[704] Carolyn Black. Taking. Theoria, 40(2):66–75, August 1974. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Howell:1974:LSP

[705] Robert Howell. The logical structure of pictorial representation. Theoria, 40(2):76–109, August 1974. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Fine:1974:ACS

[706] Kit Fine. An ascending chain of S4 logics. Theoria, 40(2):110–116, August 1974. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Lachlan:1974:NTR

[707] A. H. Lachlan. A note on Thomason’s refined structures for tense logics. Theoria, 40(2):117–120, August 1974. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kavka:1974:EEU

[708] Gregory S. Kavka. Extensional equivalence and utilitarian generalization. Theoria, 40(3):125–147, December 1974. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Gabbay:1974:NLC

[709] Dov M. Gabbay. A normal logic that is complete for neighborhood frames but not for Kripke frames. Theoria, 40(3):148–153, December 1974. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Thomason:1974:RTL

[710] S. K. Thomason. Reduction of tense logic to modal logic II. Theoria, 40(3):154–169, December 1974. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Gardenfors:1974:R

[711] Peter Gärdenfors. Review. Theoria, 40(3):170–175, December 1974. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Titiev:1975:SSS

[712] Robert J. Titiev. Sentences, semantics, stimuli, and Quine. Theoria, 41(1):1–10, April 1975. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hughes:1975:OLC

[713] G. E. Hughes and M. J. Cresswell. Omnitemporal logic and converging time. Theoria, 41(1):11–34, April 1975. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hintikka:1975:CPB

[714] Jaakko Hintikka. Comment on Professor Bergström. Theoria, 41(1):35– 38, April 1975. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic). See reply [717].

Anonymous:1975:R

[715] Anonymous. Reviews. Theoria, 41(1):39–60, April 1975. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Malmgren:1975:IRA

[716] H. Malmgren. Internal relations in the analysis of consciousness. Theoria, 41(2):61–83, August 1975. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bergstrom:1975:RPH

[717] Lars Bergström. Reply to Professor Hintikka. Theoria, 41(2):84, August 1975. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic). See [714].

Hughes:1975:BSS

[718] G. E. Hughes. B(S4.3, S4) unveiled. Theoria, 41(2):85–88, August 1975. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kessler:1975:NFB

[719] Glenn Kessler. A note on future branching time. Theoria, 41(2):89–95, August 1975. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Aaberg:1975:GEE

[720] Claes Åberg. Generic extensions and elementary embeddings. Theoria, 41(2):96–104, August 1975. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Carter:1975:PEP

[721] W. R. Carter. Plato on essence: Phaedo. Theoria, 41(3):103–111, December 1975. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Cummins:1975:ECC

[722] Robert Cummins. Epistemology and the Cartesian circle. Theoria, 41(3): 112–124, December 1975. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kavka:1975:EEU

[723] Gregory S. Kavka. Extensional equivalence and utilitarian generalization. Theoria, 41(3):125–147, December 1975. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Gabbay:1975:NLC

[724] Dov M. Gabbay. A normal logic that is complete for neighborhood frames but not for Kripke frames. Theoria, 41(3):148–153, December 1975. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Thomason:1975:RTL

[725] S. K. Thomason. Reduction of tense logic to modal logic II. Theoria, 41(3):154–169, December 1975. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Gardenfors:1975:R

[726] Peter Gärdenfors. Review. Theoria, 41(3):170–175, December 1975. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Donagan:1976:RFM

[727] Alan Donagan. Realism and freethinking in metaphysics. Theoria, 42 (1–3):1–19, April 1976. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hochberg:1976:SSM

[728] Herbert Hochberg. Strawson, scepticism, and metaphysics. Theoria, 42 (1–3):20–43, April 1976. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Castaneda:1976:OGR

[729] Hector-Neri Castañeda. Ontology and grammar: I. Russell’s paradox and the general theory of properties in natural language. Theoria, 42 (1–3):44–92, April 1976. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Allaire:1976:OST

[730] Edwin B. Allaire. Ontologically speaking, things are …. Theoria, 42 (1–3):93–114, April 1976. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Brownstein:1976:DCF

[731] Donald Brownstein. Denoting, corresponding and facts. Theoria, 42 (1–3):115–138, April 1976. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Plantinga:1976:APW

[732] Alvin Plantinga. Actualism and possible worlds. Theoria, 42(1–3):139– 160, April 1976. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Rosen:1976:SD

[733] Stanley Rosen. Socrates’ dream. Theoria, 42(1–3):161–188, April 1976. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Malmgren:1976:IKO

[734] Helge Malmgren. Immediate knowledge of other minds. Theoria, 42 (1–3):189–205, April 1976. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1977:A

[735] Anonymous. Announcement. Theoria, 43(1):1, April 1977. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Prawitz:1977:MPC

[736] Dag Prawitz. Meaning and proofs: on the conflict between classical and intuitionistic logic. Theoria, 43(1):2–40, April 1977. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Margolis:1977:SOB

[737] Joseph Margolis. The stubborn opacity of belief contexts. Theoria, 43 (1):41–46, April 1977. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sundholm:1977:CPI

[738] GÖRan Sundholm. A completeness proof for an infinitary tense-logic. Theoria, 43(1):47–51, April 1977. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Esakia:1977:FCM

[739] L. Esakia and V. Meskhi. Five critical modal systems. Theoria, 43(1):52– 60, April 1977. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Lovtrup:1977:VSI

[740] Søren Løvtrup. Variation, selection, isolation, environment: An analysis of Darwin’s theory. Theoria, 43(2):65–83, August 1977. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bergstrom:1977:UFM

[741] Lars Bergström. Utilitarianism and future mistakes. Theoria, 43(2):84– 102, August 1977. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hammond:1977:WCH

[742] Michael Hammond. Weighting causes in historical explanation. Theoria, 43(2):103–128, August 1977. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Haack:1977:ALT

[743] Susan Haack. Analyticity and logical truth in the roots of reference. Theoria, 43(2):129–143, August 1977. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ericsson:1977:SRE

[744] Lars O. Ericsson. Some reflections on extended utilitarianism. Theoria, 43(3):145–156, December 1977. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Lubow:1977:DMS

[745] Neil Lubow. On defining mental state names. Theoria, 43(3):157–173, December 1977. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Miller:1977:UAF

[746] David Miller. The uniqueness of atomic facts in Wittgenstein’s Tractatus. Theoria, 43(3):174–185, December 1977. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Adams:1977:NCP

[747] Ernest W. Adams. A note on comparing probabilistic and modal logics of conditionals. Theoria, 43(3):186–194, December 1977. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Isard:1977:FAU

[748] Stephen Isard. A finitely axiomatizable undecidable extension of K. Theoria, 43(3):195–202, December 1977. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tannsjo:1978:MAE

[749] Torbjörn Tännsjö. The morality of abstract entities. Theoria, 44(1):1– 18, April 1978. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Rabinowicz:1978:UCO

[750] Wlodzimierz Rabinowicz. Utilitarianism and conflicting obligations. Theoria, 44(1):19–24, April 1978. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

VanBenthem:1978:TSI

[751] J. F. A. K. Van Benthem. Two simple incomplete modal logics. Theoria, 44(1):25–37, April 1978. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Goldblatt:1978:ANP

[752] Rob Goldblatt. Arithmetical necessity, provability and intuitionistic logic. Theoria, 44(1):38–46, April 1978. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1978:R

[753] Anonymous. Reviews. Theoria, 44(1):47–51, April 1978. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Gardenfors:1978:FIC

[754] Peter Gärdenfors. Fairness without interpersonal comparisons. Theoria, 44(2):57–74, August 1978. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Gullvaag:1978:LA

[755] Ingemund Gullvåg. The logic of assertion. Theoria, 44(2):75–116, August 1978. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

VanBenthem:1978:TFD

[756] J. F. A. K. Van Benthem and W. J. Blok. Transitivity follows from Dummett’s axiom. Theoria, 44(2):117–118, August 1978. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Marc-Wogau:1978:MAW

[757] Konrad Marc-Wogau. In memoriam: Anders wedberg. Theoria, 44(3): 122–123, December 1978. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1978:PWA

[758] Anonymous. Philosophical works by Anders Wedberg. Theoria, 44(3): 124–130, December 1978. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hill:1978:KAM

[759] Thomas E. Hill, Jr. Kant’s anti-moralistic strain. Theoria, 44(3):131–151, December 1978. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Lindstrom:1978:OUT

[760] Per Lindström. Omitting uncountable types and extensions of elementary logic. Theoria, 44(3):152–156, December 1978. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Burgess:1978:UF

[761] John P. Burgess. The unreal future. Theoria, 44(3):157–179, December 1978. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Smokler:1979:CMD

[762] Howard Smokler. The collapse of modal distinctions in probabilistic contexts. Theoria, 45(1):1–7, April 1979. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Cole:1979:PM

[763] Richard Cole. Possibility matrices. Theoria, 45(1):8–39, April 1979. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Nilson:1979:WHC

[764] Sten Sparre Nilson. Weighting of historical causes. Theoria, 45(1):40– 43, April 1979. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:1979:RLL

[765] Bengt Hansson. Review: Lindahl, Lars. Position and change: A study in law and logic. Theoria, 45(1):44–48, April 1979. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Fischer:1979:LNM

[766] John Martin Fischer. Lehrer’s new move: ‘can’ in theory and practice. Theoria, 45(2):49–62, August 1979. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

VanBenthem:1979:SAM

[767] J. F. A. K. Van Benthem. Syntactic aspects of modal incompleteness theorems. Theoria, 45(2):63–77, August 1979. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Mckinsey:1979:ADD

[768] Michael Mckinsey. The ambiguity of definite descriptions. Theoria, 45 (2):78–89, August 1979. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sundholm:1979:RDM

[769] Göran Sundholm. Review: Dummett, Michael. The elements of intuitionism. Theoria, 45(2):90–95, August 1979. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Fitting:1979:AAC

[770] Melvin C. Fitting. An axiomatic approach to computers. Theoria, 45 (3):97–113, December 1979. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tuomela:1979:PR

[771] Raimo Tuomela. Putnam’s realisms. Theoria, 45(3):114–126, December 1979. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hall:1979:SPD

[772] Richard J. Hall. Seeing perfectly dark things and the causal conditions of seeing. Theoria, 45(3):127–134, December 1979. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sutton:1979:KCE

[773] Agneta Sutton. The Kantian and the consequentialist elements in Rawls’s theory of justice. Theoria, 45(3):135–140, December 1979. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Dahlquist:1980:AML

[774] Thorild Dahlquist. Andries Mac leod (1891–1977) in memoriam. Theoria, 46(1):1–4, April 1980. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1980:WAM

[775] Anonymous. Writings of andries Mac leod. Theoria, 46(1):5–6, April 1980. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Boolos:1980:SML

[776] George Boolos. On systems of modal logic with provability interpretations. Theoria, 46(1):7–18, April 1980. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sainsbury:1980:RCF

[777] R. M. Sainsbury. Russell on constructions and fictions. Theoria, 46(1): 19–36, April 1980. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bar-Elli:1980:CDR

[778] Gilead Bar-Elli. Constituents and denotation in Russell. Theoria, 46(1): 37–51, April 1980. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1980:R

[779] Anonymous. Reviews. Theoria, 46(1):52–58, April 1980. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bender:1980:PDS

[780] Gloria Bender. Philosophical dissertations in Sweden 1970–1979: Philosophical dissertations in Sweden 1970–1979. Theoria, 46(1):59–62, April 1980. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Musgrave:1980:WI

[781] Alan Musgrave. Wittgensteinian instrumentalism. Theoria, 46(2–3):65– 105, August 1980. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tweedale:1980:AOC

[782] Martin M. Tweedale. Abailard and Ockham: Contrasting defences of nominalism. Theoria, 46(2–3):106–122, August 1980. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Cresswell:1980:JP

[783] M. J. Cresswell. Jackson on perception. Theoria, 46(2–3):123–147, August 1980. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Currie:1980:CN

[784] Gregory Currie and Graham Oddie. Changing numbers. Theoria, 46(2– 3):148–164, August 1980. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tichy:1980:TT

[785] Pavel Tichý. The transiency of truth. Theoria, 46(2–3):165–182, August 1980. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Segerberg:1980:NLE

[786] Krister Segerberg. A note on the logic of elsewhere. Theoria, 46(2– 3):183–187, August 1980. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Malcolm:1980:RTC

[787] Wilf Malcolm. Review: Topot, The categorial analysis of logic. Theoria, 46(2–3):188–192, August 1980. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Vision:1981:EE

[788] Gerald Vision. Existentials and existents. Theoria, 47(1):1–30, April 1981. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Denyer:1981:TMD

[789] Nicholas Denyer. Time and modality in Diodorus Cronus. Theoria, 47 (1):31–53, April 1981. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Rosenbluth:1981:RIM

[790] Richard Rosenbluth. The role of intention in moral conventions. Theoria, 47(1):54–62, April 1981. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Cohen:1981:WSW

[791] L. Jonathan Cohen. Who is starving whom? Theoria, 47(2):65–81, August 1981. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Belliotti:1981:NPD

[792] Raymond A. Belliotti. Negative and positive duties. Theoria, 47(2):82– 92, August 1981. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Stevens:1981:ITK

[793] Rex Patrick Stevens. The impact of theodicy on Kant’s conception of moral practice. Theoria, 47(2):93–108, August 1981. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Lewis:1981:WFB

[794] David Lewis. Are we free to break the laws? Theoria, 47(3):113–121, December 1981. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hedenius:1981:RE

[795] Ingemar Hedenius. On relativism in ethics. Theoria, 47(3):122–133, December 1981. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bacon:1981:PPM

[796] John Bacon. Purely physical modalities. Theoria, 47(3):134–141, December 1981. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Flage:1981:LRI

[797] Daniel E. Flage. Locke’s relative ideas. Theoria, 47(3):142–159, December 1981. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Persson:1981:RRN

[798] Zngmar Persson. Review: Robert Nozick. Philosophical Explanations. Theoria, 47(3):160–166, December 1981. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hallden:1982:IHM

[799] Soren Hallden. Ingemar Hedenius: In memoriam. Theoria, 48(1):1– 3, April 1982. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ryding:1982:TAP

[800] Erik Ryding. Are there “archetypical performatives”? Theoria, 48(1): 4–13, April 1982. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Alchourron:1982:LTC

[801] Carlos E. Alchourron and David Makinson. On the logic of theory change: Contraction functions and their associated revision functions. Theoria, 48(1):14–37, April 1982. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Williams:1982:AMP

[802] John N. Williams. The absurdities of Moore’s paradoxes. Theoria, 48(1): 38–46, April 1982. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Stenius:1982:RPW

[803] Erik Stenius. Ross’ paradox and well-formed codices. Theoria, 48(2):49– 77, August 1982. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Polakow:1982:RDR

[804] Avron Polakow. Radical defeasability and rigid designation. Theoria, 48(2):78–89, August 1982. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bertolet:1982:RSI

[805] Rod Bertolet. Russell and Strawson, indexical and improper descriptions. Theoria, 48(2):90–98, August 1982. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1982:R

[806] Anonymous. Reviews. Theoria, 48(2):99–112, August 1982. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sobel:1982:UPI

[807] Jordan Howard Sobel. Utilitarian principles for imperfect agents. Theoria, 48(3):113–126, December 1982. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hart:1982:PB

[808] W. D. Hart. Prior and Belnap. Theoria, 48(3):127–138, December 1982. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Callen:1982:MML

[809] Donald M. Callen. Making music live. Theoria, 48(3):139–168, December 1982. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Olding:1983:BK

[810] A. Olding. Biology and knowledge. Theoria, 49(1):1–22, April 1983. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Clark:1983:MEE

[811] Andrew J. Clark. Meaning and evolutionary epistemology. Theoria, 49 (1):23–31, April 1983. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tennant:1983:DEE

[812] Neil Tennant. In defence of evolutionary epistemology. Theoria, 49(1): 32–48, April 1983. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Carruthers:1983:CO

[813] Peter Carruthers. On concept and object. Theoria, 49(2):49–86, August 1983. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hesslow:1983:EDW

[814] Germund Hesslow. Explaining differences and weighting causes. Theoria, 49(2):87–111, August 1983. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Shenoy:1983:TID

[815] Prakash P. Shenoy and Rex Martin. Two interpretations of the difference principle in Rawls’s theory of justice. Theoria, 49(3):113–141, December 1983. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Carson:1983:SCR

[816] Thomas L. Carson. Strict compliance and Rawls’s critique of utilitarianism. Theoria, 49(3):142–158, December 1983. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sobel:1983:EUR

[817] Jordan Howard Sobel. Expected utilities and rational actions and choices. Theoria, 49(3):159–183, December 1983. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Berger:1983:BRH

[818] George Berger. Book review: Husserl and intentionality: A study of mind, meaning, and language. Theoria, 49(3):184–188, December 1983. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kvart:1984:HPC

[819] Igal Kvart. The Hesperus–phosphorus case. Theoria, 50(1):1–35, April 1984. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Pettersson:1984:OLW

[820] Anders Pettersson. The ontology of literary works. Theoria, 50(1):36– 51, April 1984. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1984:R

[821] Anonymous. Review. Theoria, 50(1):52–63, April 1984. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Stenius:1984:ANA

[822] Erik Stenius. Autobiographical note: Autobiographical note. Theoria, 50 (2–3):67–72, August 1984. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Young:1984:WKS

[823] Julian Young. Wittgenstein, Kant, Schopenhauer, and critical philosophy. Theoria, 50(2–3):73–105, August 1984. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

VanBenthem:1984:ASS

[824] Johan Van Benthem. Analytic/synthetic: Sharpening a philosophical tool. Theoria, 50(2–3):106–137, August 1984. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Makinson:1984:SAD

[825] David Makinson. Stenius’ approach to disjunctive permission. Theoria, 50(2–3):138–147, August 1984. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Copeland:1984:HHR

[826] B. I. Copeland. Horseshoe, hook, and relevance. Theoria, 50(2–3):148– 164, August 1984. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kuhn:1984:SM

[827] Steven T. Kuhn. Stenius on meaning. Theoria, 50(2–3):165–177, August 1984. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kroon:1984:SP

[828] Fred Kroon. Stenius on the paradoxes. Theoria, 50(2–3):178–211, August 1984. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kivinen:1984:SPG

[829] S. Albert Kivinen. Stenius on the philosophy of geometry. Theoria, 50(2– 3):212–240, August 1984. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Porn:1984:ASP

[830] Ingmar Pörn. Analysis of signalling: A problem and its solution. Theoria, 50(2–3):241–255, August 1984. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tengstrom:1984:GP

[831] Leif Tengström. Get the picture? Theoria, 50(2–3):256–266, August 1984. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Stenius:1984:CC

[832] Erik Stenius. Comments: Comments. Theoria, 50(2–3):267–309, August 1984. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1984:BPW

[833] Anonymous. Bibliography of the philosophical writings of Erik Stenius 1935–1984. Theoria, 50(2–3):310–320, August 1984. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tiles:1985:RAD

[834] J. E. Tiles. A ‘rationalist’ approach to dispositional concepts. Theoria, 51(1):1–15, April 1985. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Brown:1985:PAP

[835] Geoffrey Brown. On pleasing all of the people all the time. Theoria, 51(1):16–23, April 1985. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Peczenik:1985:FTT

[836] Aleksander Peczenik and Jerzy Wróblewski. Fuzziness and transformation: Towards explaining legal reasoning. Theoria, 51(1):24–44, April 1985. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1985:R

[837] Anonymous. Reviews. Theoria, 51(1):45–60, April 1985. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Jackson:1985:COE

[838] Frank Jackson and Robert Pargetter. Causal origin and evidence. Theoria, 51(2):65–76, August 1985. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sajama:1985:SHV

[839] Seppo Sajama. Supererogation and high values. Theoria, 51(2):77–88, August 1985. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bellin:1985:SND

[840] Gianluigi Bellin. A system of natural deduction for GL. Theoria, 51 (2):89–114, August 1985. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1985:RHR

[841] Anonymous. Review: S Heikki. Re-enactment: A study of R. G. Collingwood’s philosophy of history. Acta Academiae Aboensis, Series A (Humaniora). Theoria, 51(2):115–124, August 1985. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Odegard:1985:MRA

[842] Douglas Odegard. Moral responsibility and alternatives. Theoria, 51(3): 125–136, December 1985. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Person:1985:UBE

[843] Ingmar Person. The universal basis of egoism. Theoria, 51(3):137–158, December 1985. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Dosen:1985:CU

[844] Kosta Došen and Peter Schroeder-Heister. Conservativeness and uniqueness. Theoria, 51(3):159–173, December 1985. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Brownstein:1985:TPA

[845] Donald Brownstein. Troubles with Plantinga’s actualism. Theoria, 51(3): 174–189, December 1985. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1985:RIP

[846] Anonymous. Review: Ingmar Persson, The Primacy of Perception: Towards a Neutral Monism. Library of Theoria, Lund 1985. Theoria, 51(3): 190–194, December 1985. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Johansson:1986:LIT

[847] Ingvar Johansson. Levels of intension and theories of reference. Theoria, 52(1–2):1–15, April 1986. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Wreen:1986:PP

[848] Michael J. Wreen. The power of potentiality. Theoria, 52(1–2):16–40, April 1986. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Robins:1986:IC

[849] Michael H. Robins. Intention and control. Theoria, 52(1–2):41–56, April 1986. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Schlesinger:1986:R

[850] George N. Schlesinger. Relevance. Theoria, 52(1–2):57–67, April 1986. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bar-elli:1986:CRJ

[851] Gilead Bar-elli and David Heyd. Can revenge be just or otherwise justified? Theoria, 52(1–2):68–86, April 1986. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:1986:ICA

[852] Sven Ove Hansson. Individuals and collective actions. Theoria, 52(1–2): 87–97, April 1986. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bishop:1986:RSS

[853] John Bishop. Review: Sajama, Seppo. Idea, Judgement and Will, University of Turku, Finland, 1983, Reports from the Department of Theoretical Philosophy. Theoria, 52(1–2):98–117, April 1986. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Walton:1986:RJT

[854] Douglas N. Walton. Review: J. Talja. Studies in Possibility, Ability and Not-Doing, Reports from the Department of Philosophy, No. 11, University of Turku, Turku, Finland. Theoria, 52(1–2):117–123, April 1986. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Cerezo:1986:RTR

[855] José A. López Cerezo. Review: Tuomela, Raimo. A Theory of Social Action. Dordrecht: Reidel, 1984. Theoria, 52(1–2):123–126, April 1986. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1986:E

[856] Anonymous. Errata. Theoria, 52(1–2):127, April 1986. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kvart:1986:BB

[857] Igal Kvart. Beliefs and believing. Theoria, 52(3):129–145, December 1986. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Davies:1986:ETA

[858] Stephen Davies. The expression theory again. Theoria, 52(3):146–167, December 1986. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Dudman:1986:AC

[859] V. H. Dudman. Antecedents and consequents. Theoria, 52(3):168–199, December 1986. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1986:RPB

[860] Anonymous. Review: Patrice Bailhache: Les Normes Dans Le Temps et Sur l’action. Université de Nantes, 1986. Theoria, 52(3):200–208, December 1986. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sobel:1987:LSH

[861] Jordan Howard Sobel. The law student and his teacher. Theoria, 53(1): 1–18, April 1987. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Clark:1987:MPE

[862] Andy Clark. Meaning, publicity and epistemology. Theoria, 53(1):19– 30, April 1987. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Svenonius:1987:TWC

[863] Lars Svenonius. Three ways to conceive of functions and relations. Theoria, 53(1):31–58, April 1987. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1987:R

[864] Anonymous. Review. Theoria, 53(1):59–71, April 1987. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Odegard:1987:DMD

[865] Douglas Odegard. Deep moral dilemmas. Theoria, 53(2–3):73–86, August 1987. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tannsjo:1987:MIM

[866] Torbjörn Tännsjö. The moral import of modal realism. Theoria, 53 (2–3):87–96, August 1987. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bigelow:1987:BBS

[867] John Bigelow and Robert Pargetter. Beyond the blank stare. Theoria, 53 (2–3):97–114, August 1987. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Oderberg:1987:KQ

[868] David S. Oderberg. Kripke and ‘quus’. Theoria, 53(2–3):115–120, August 1987. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Williams:1987:PPD

[869] John N. Williams. The preface paradox dissolved. Theoria, 53(2–3):121– 140, August 1987. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Balaban:1987:ROP

[870] Oded Balaban. Relation and object in Plato’s approach to knowledge. Theoria, 53(2–3):141–159, August 1987. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hermeren:1987:REP

[871] Göran Hermerén. Review: Enzyklopädie Philosophic und Wissenschaftstheorie in 3 Bänden, Edited by Jürgen Mittelstrass. Theoria, 53(2–3): 160–164, August 1987. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Johanson:1988:ILB

[872] Arnold Johanson. Imperative logic as based on a Galois connection. Theoria, 54(1):1–24, April 1988. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Malachowski:1988:SFP

[873] Alan Malachowski. Searle on first person meaning and indeterminacy. Theoria, 54(1):25–30, April 1988. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Weaver:1988:RPU

[874] George Weaver. Reading proofs with understanding. Theoria, 54(1):31– 47, April 1988. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1988:Ra

[875] Anonymous. Review. Theoria, 54(1):48–78, April 1988. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Mckie:1988:CRS

[876] Johnr. Mckie. Conventionalism, realism, and spacetime structure. Theoria, 54(2):81–101, August 1988. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Nielsen:1988:MRM

[877] Kai Nielsen. Marxism and the rejection of morality. Theoria, 54(2):102– 128, August 1988. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1988:Rb

[878] Anonymous. Review. Theoria, 54(2):129–160, August 1988. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Lewis:1988:RI

[879] David Lewis. Relevant implication. Theoria, 54(3):161–174, December 1988. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Belnap:1988:SIC

[880] Nuel Belnap and Michael Perloff. Seeing to it that: a canonical form for agentives. Theoria, 54(3):175–199, December 1988. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sandoe:1988:SQS

[881] Peter Sandøe. Secondary qualities — subjective and intrinsic. Theoria, 54(3):200–219, December 1988. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1988:Rc

[882] Anonymous. Review. Theoria, 54(3):220–237, December 1988. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Rabinowicz:1989:AUP

[883] Wlodzimierz Rabinowicz. Act-utilitarian prisoner’s dilemmas. Theoria, 55(1):1–44, April 1989. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Stephenson:1989:BR

[884] Wendell Stephenson. Benevolence and resentment. Theoria, 55(1):45– 61, April 1989. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Persson:1989:RTD

[885] Ingmar Persson. Review: A theory of determinism. The mind, neuroscience, and life-hopes. Theoria, 55(1):62–76, April 1989. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Odegard:1989:CMI

[886] Douglas Odegard. Charity and moral imperatives. Theoria, 55(2):81–94, August 1989. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Corlett:1989:KPR

[887] J. Angelo Corlett. Is Kripke’s puzzle really a puzzle? Theoria, 55(2): 95–113, August 1989. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:1989:NOT

[888] Svenove Hansson. New operators for theory change. Theoria, 55(2):114– 132, August 1989. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1989:R

[889] Anonymous. Review. Theoria, 55(2):133–144, August 1989. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Rabinowicz:1989:HP

[890] Wlodzimierz Rabinowicz. Hare on prudence. Theoria, 55(3):145–151, December 1989. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hare:1989:PPP

[891] R. M. Hare. Prudence and past preferences: Reply to wlodzimierz rabinowicz. Theoria, 55(3):152–158, December 1989. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Persson:1989:USD

[892] Ingmar Persson. Universalizability and the summing of desires. Theoria, 55(3):159–170, December 1989. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hare:1989:USD

[893] R. M. Hare. Universalizability and the summing of desires: Reply to ingmar Persson. Theoria, 55(3):171–177, December 1989. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Wetterstrom:1989:I

[894] Thomas Wetterström. On intuitions. Theoria, 55(3):178–185, December 1989. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hare:1989:FRT

[895] R. M. Hare. Fanaticsm: Reply to Thomas Wetterström. Theoria, 55(3): 186–190, December 1989. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sandoe:1989:ALM

[896] Peter Sandøe. Amoralism — on the limits of moral thinking. Theoria, 55(3):191–204, December 1989. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hare:1989:ARP

[897] R. M. Hare. Amoralism: Reply to Peter Sandøe. Theoria, 55(3):205–210, December 1989. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sandoe:1989:RDS

[898] Peter Sandøe. Review: Douglas Seanor & N. Fotion (eds.): Hare and Critics. Theoria, 55(3):211–224, December 1989. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hallden:1990:POE

[899] Sören Halldén. Per Olof Ekelöf (1906–1990). Theoria, 56(1–2):1–2, April 1990. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Wright:1990:TMP

[900] Edmond Wright. Two more proofs of present qualia. Theoria, 56(1–2): 3–22, April 1990. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Weintraub:1990:OOP

[901] Ruth Weintraub. Objectivism without objective probabilities. Theoria, 56(1–2):23–41, April 1990. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Denyer:1990:EDM

[902] Nicholas Denyer. Ease and difficulty: a modal logic with deontic applications. Theoria, 56(1–2):42–61, April 1990. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Brown:1990:QQ

[903] Mark A. Brown. Questions and quantifiers. Theoria, 56(1–2):62–84, April 1990. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Loptson:1990:LIC

[904] Peter Loptson. Lockean ideas and 18th century British philosophy. Theoria, 56(1–2):85–106, April 1990. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1990:BR

[905] Anonymous. Book reviews. Theoria, 56(1–2):107–127, April 1990. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bigelow:1990:AQ

[906] John Bigelow and Robert Pargetter. Acquaintance with qualia. Theoria, 56(3):129–147, December 1990. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Percival:1990:ITI

[907] Philip Percival. Indices of truth and intensional operators. Theoria, 56(3):148–172, December 1990. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Klima:1990:NQG

[908] Gyula Klima and Gabriel Sandu. Numerical quantifiers in gametheoretical semantics. Theoria, 56(3):173–192, December 1990. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Nielsen:1990:TBP

[909] Kai Nielsen. On there being philosophical knowledge. Theoria, 56(3): 193–225, December 1990. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1990:R

[910] Anonymous. Review. Theoria, 56(3):226–235, December 1990. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Humberstone:1991:SSS

[911] I. L. Humberstone. A study of some ‘separated’ conditions on binary relations. Theoria, 57(1–2):1–16, April 1991. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kroon:1991:DDF

[912] Frederick W. Kroon. Denotation and description in free logic. Theoria, 57(1–2):17–41, April 1991. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Oddie:1991:AVE

[913] Graham Oddie and Peter Milne. Act and value: Expectation and the representability of moral theories. Theoria, 57(1–2):42–76, April 1991. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ranta:1991:CPW

[914] Aarne Ranta. Constructing possible worlds. Theoria, 57(1–2):77–99, April 1991. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Surendonk:1991:MMR

[915] Timothy J. Surendonk. Making maximal reliable action maximal. Theoria, 57(1–2):101–110, April 1991. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Singer:1991:AEH

[916] Peter Singer. Applied ethics in a hostile environment. Theoria, 57(1– 2):111–114, April 1991. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Crisp:1991:R

[917] Roger Crisp, Sören Halldén, and Torbjörn Tännsjö. Review. Theoria, 57 (1–2):115–126, April 1991. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Galavotti:1991:EEP

[918] Maria Carla Galavotti. Editorial: Editor’s preface. Theoria, 57(3):129– 131, December 1991. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sahlin:1991:ORP

[919] Nils-Eric Sahlin. Obtained by a reliable process and always leading to success. Theoria, 57(3):132–149, December 1991. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Simons:1991:RPU

[920] Peter Simons. Ramsey, particulars, and universals. Theoria, 57(3):150– 161, December 1991. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Majer:1991:RTT

[921] Ulrich Majer. Ramsey’s theory of truth and the truth of theories: A synthesis of pragmatism and intuitionism in Ramsey’s last philosophy. Theoria, 57(3):162–195, December 1991. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Egidi:1991:RWS

[922] Rosaria Egidi. Ramsey and Wittgenstein on scientific theories. Theoria, 57(3):196–210, December 1991. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Zabell:1991:RTP

[923] S. L. Zabell. Ramsey, truth, and probability. Theoria, 57(3):211–238, December 1991. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Galavotti:1991:NSP

[924] Maria Carla Galavotti. The notion of subjective probability in the work of Ramsey and de Finetti. Theoria, 57(3):239–259, December 1991. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Mares:1992:ADL

[925] Edwin D. Mares. Andersonian deontic logic. Theoria, 58(1):1–2, April 1992. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hermeren:1992:MM

[926] Goran Hermerén. Manfred moritz (1909–1990). Theoria, 58(1):3–20, April 1992. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Rott:1992:MBC

[927] Hans Rott. Modellings for belief change: Prioritization and entrenchment. Theoria, 58(1):21–57, April 1992. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hilpinen:1992:AWA

[928] Risto Hilpinen. On artifacts and works of art. Theoria, 58(1):58–82, April 1992. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1992:Ra

[929] Anonymous. Review. Theoria, 58(1):83–96, April 1992. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hallden:1992:KMW

[930] Sören Halldén. Konrad Marc-Wogau (1902–1991). Theoria, 58(2–3):97– 98, August 1992. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Divers:1992:MSM

[931] John Divers. Modal supereminence and modal realism. Theoria, 58(2– 3):99–115, August 1992. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Cresswell:1992:OSM

[932] M. J. Cresswell. The ontological status of matter in Aristotle. Theoria, 58 (2–3):116–130, August 1992. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Lindstrom:1992:RTR

[933] Sten Lindström and Włodzimierz Rabinowicz. The Ramsey test revisited. Theoria, 58(2–3):131–182, August 1992. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Magill:1992:EDS

[934] Kevin Magill. Epicurus, determinism and the security of knowledge. Theoria, 58(2–3):183–196, August 1992. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Horowitz:1992:FRI

[935] Amir Horowitz. Functional role and intentionality. Theoria, 58(2–3): 197–218, August 1992. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1992:Rb

[936] Anonymous. Review. Theoria, 58(2–3):219–239, August 1992. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1993:EP

[937] Anonymous. Editors’ preface. Theoria, 59(1–3):1–2, April 1993. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hallden:1993:ANA

[938] Sören Halldén. Autobiographical notes: Autobiographical notes. Theoria, 59(1–3):3–17, April 1993. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1993:AN

[939] Anonymous. Autobiographical notes. Theoria, 59(1–3):18–27, April 1993. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bjorklund:1993:PEK

[940] Stefan Björklund. The philosophy of everyday knowledge. Theoria, 59 (1–3):28–52, April 1993. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Furberg:1993:PPF

[941] Mats Furberg. Philosophical polyglottism and first person perspectives. Theoria, 59(1–3):53–79, April 1993. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hermeren:1993:EPR

[942] Goran Hermerén. Emotive properties:the role of abstraction, introspection and projection. Theoria, 59(1–3):80–112, April 1993. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Persson:1993:HUM

[943] Ingmar Persson. Hallden on the unity of the mind and the self. Theoria, 59(1–3):113–123, April 1993. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Levi:1993:EMD

[944] Isaac Levi. The evolutionary matrix of deliberation according to Halldén. Theoria, 59(1–3):124–143, April 1993. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Edman:1993:EEM

[945] Martin Edman. Evolutionary epistemology and the miracle of knowledge. Theoria, 59(1–3):144–157, April 1993. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hintikka:1993:PPI

[946] Jaakko Hintikka. On proper (popper?) and improper uses of information in epistemology. Theoria, 59(1–3):158–165, April 1993. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Levi:1993:EMR

[947] Isaac Levi. Evidentiary mechanisms and routine expansion. Theoria, 59 (1–3):166–177, April 1993. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sahlin:1993:WC

[948] Nils-Eric Sahlin. Worthy of choice. Theoria, 59(1–3):178–191, April 1993. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Schumm:1993:WDH

[949] George F. Schumm. Why does Halldén-completeness matter? Theoria, 59(1–3):192–206, April 1993. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Finn:1993:NLT

[950] Victor K. Finn and Revaz Grigolia. Nonsense logics and their algebraic properties. Theoria, 59(1–3):207–273, April 1993. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hallden:1993:RMC

[951] Sören Halldën. Reply to my commentators. Theoria, 59(1–3):274–275, April 1993. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hallden:1993:PEK

[952] Sören Halldën. The philosophy of everyday knowledge. Theoria, 59(1– 3):276–292, April 1993. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hallden:1993:PW

[953] Sören Halldën. Philosophical works. Theoria, 59(1–3):293–300, April 1993. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sobel:1994:PCQ

[954] Jordan Howard Sobel. Plato’s crito: a question of agreement. Theoria, 60(1):1–26, April 1994. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Swirydowicz:1994:NCR

[955] Kazimierz Swirydowicz. Normative consequence relation and consequence operations on the language of dyadic deontic logic. Theoria, 60(1):27–47, April 1994. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Haji:1994:COI

[956] Ishtiyaque Haji. Changing obligations and immutable blameworthiness. Theoria, 60(1):48–62, April 1994. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Gjelsvik:1994:R

[957] Olav Gjelsvik. Reviews. Theoria, 60(1):63–77, April 1994. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Odegard:1994:KTN

[958] Douglas Odegard. Knowing that there is no demon. Theoria, 60(2):81– 98, August 1994. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Nielsen:1994:HPP

[959] Kai Nielsen. How to proceed in philosophy: remarks after Habermas. Theoria, 60(2):99–128, August 1994. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

DiMaio:1994:ILA

[960] Maria Concetta Di Maio. Inductive logic: aims and procedures. Theoria, 60(2):129–153, August 1994. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1994:R

[961] Anonymous. Reviews. Theoria, 60(2):154–167, August 1994. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Pruwitz:1994:RSP

[962] Dug Pruwitz. The Rolf Schock Prize in Logic and Philosophy. Theoria, 60(3):169–170, December 1994. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Quine:1994:AO

[963] W. V. Quine. Assuming objects. Theoria, 60(3):171–183, December 1994. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Davidson:1994:QP

[964] Donald Davidson. On Quine’s philosophy. Theoria, 60(3):184–192, December 1994. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bergstrom:1994:IWV

[965] Lars Bergström and Dagfinn Føllesdal. Interview with Willard Van Orman Quine in November 1993. Theoria, 60(3):193–206, December 1994. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bergstrom:1994:IDD

[966] Lars Bergström and Dagfinn Føllesdal. Interview with Donald Davidson in November 1993. Theoria, 60(3):207–225, December 1994. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1994:EBD

[967] Anonymous. Exchange between Donald Davidson and W. V. Quine following Davidson’s lecture. Theoria, 60(3):226–231, December 1994. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Humberstone:1995:NI

[968] I. L. Humberstone. Negation by iteration. Theoria, 61(1):1–24, April 1995. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Jensen:1995:MSB

[969] Karsten Klint Jensen. Measuring the size of a benefit and its moral weight on the significance of John Broome’s: “Interpersonal Addition Theorem”. Theoria, 61(1):25–60, April 1995. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Roeper:1995:BBL

[970] Peter Roeper and Hugues Leblanc. Of a and b being logically independent of each other and of their having no common factual content. Theoria, 61(1):61–79, April 1995. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1995:Ra

[971] Anonymous. Reviews. Theoria, 61(1):80–102, April 1995. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sullivan:1995:WFM

[972] Peter M. Sullivan. Wittgenstein on “The Foundations of Mathematics”, June 1927. Theoria, 61(2):105–142, August 1995. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Persson:1995:CNM

[973] Ingmar Persson. Critical notice of Michael Smith: The moral problem. Theoria, 61(2):143–158, August 1995. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Smith:1995:RIP

[974] Michael Smith. Reply to ingmar Persson’s critical notice of: The moral problem. Theoria, 61(2):159–181, August 1995. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Olson:1995:WNH

[975] Eric T. Olson. Why i have no hands. Theoria, 61(2):182–197, August 1995. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Dummett:1995:BV

[976] Michael Dummett. Bivalence and vagueness. Theoria, 61(3):201–216, December 1995. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Jacquette:1995:BBT

[977] Dale Jacquette. The blue banana trick: Dennett on Jackson’s color scientist. Theoria, 61(3):217–230, December 1995. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Heathcote:1995:AII

[978] Adrian Heathcote. Abductive inference and invalidity. Theoria, 61(3): 231–260, December 1995. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Weiss:1995:DRM

[979] Bernhard Weiss. On the demise of Russell’s multiple relations theory of judgement. Theoria, 61(3):261–282, December 1995. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1995:Rb

[980] Anonymous. Reviews. Theoria, 61(3):283–296, December 1995. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1996:Pa

[981] Anonymous. Preface. Theoria, 62(1–2):1–2, April 1996. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hintikka:1996:WBT

[982] Jaakko Hintikka. Wittgenstein on being and time. Theoria, 62(1–2):3– 18, April 1996. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

LindstroM:1996:PLS

[983] Per LindströM. Provability logic — a short introduction. Theoria, 62 (1–2):19–61, April 1996. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Segerberg:1996:TRR

[984] Krister Segerberg. Three recipes for revision. Theoria, 62(1–2):62–73, April 1996. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bergstrom:1996:RC

[985] Lars Bergström. Reflections on consequentialism. Theoria, 62(1–2):74– 94, April 1996. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Rabinowicz:1996:WIW

[986] Wlodek Rabinowicz and Bertil Strömberg. What if I were in his shoes? On Hare’s argument for preference utilitarianism. Theoria, 62(1–2):95– 123, April 1996. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bykvist:1996:UDP

[987] Krister Bykvist. Utilitarian deontologies? On preference utilitarianism and agent-relative value. Theoria, 62(1–2):124–143, April 1996. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Carlson:1996:CPR

[988] Erik Carlson. Cyclical preferences and rational choice. Theoria, 62(1– 2):144–160, April 1996. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Danielsson:1996:RCE

[989] Sven Danielsson. The refutation of cyclic evaluations. Theoria, 62(1– 2):161–168, April 1996. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:1996:WPR

[990] Sven Ove Hansson. What is philosophy of risk? Theoria, 62(1–2):169– 186, April 1996. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Stenlund:1996:LM

[991] SÖRen Stenlund. Language and metaphysics. Theoria, 62(1–2):187– 211, April 1996. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1996:Pb

[992] Anonymous. Preface. Theoria, 62(3):213, December 1996. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Corblin:1996:MNP

[993] Francis Corblin. Multiple negation processing in natural language. Theoria, 62(3):214–259, December 1996. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Morrill:1996:GL

[994] Glyn Morrill. Grammar and logic. Theoria, 62(3):260–293, December 1996. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:1997:EIW

[995] Sven Ove Hansson. Editor’s introduction: What’s new isn’t always best. Theoria, 63(1–2):1–13, April 1997. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Makinson:1997:SR

[996] David Makinson. Screened revision. Theoria, 63(1–2):14–23, April 1997. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bellot:1997:SAC

[997] D. Bellot, C. Godefroid, P. Han, J. P. Prost, K. Schlechta, and E. Wurbel. A semantical approach to the concept of screened revision. Theoria, 63 (1–2):24–33, April 1997. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Schlechta:1997:NPB

[998] Karl Schlechta. Non-prioritized belief revision based on distances between models. Theoria, 63(1–2):34–53, April 1997. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Cantwell:1997:LSC

[999] John Cantwell. On the logic of small changes in hypertheories. Theoria, 63(1–2):54–89, April 1997. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Fuhrmann:1997:SB

[1000] André Fuhrmann. Solid belief. Theoria, 63(1–2):90–104, April 1997. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Olsson:1997:CIS

[1001] Erik J. Olsson. A coherence interpretation of semi-revision. Theoria, 63 (1–2):105–134, April 1997. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bennet:1997:EPE

[1002] Christian Bennet. Editor’s preface: Editor’s preface. Theoria, 63(3): 137–138, December 1997. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Beklemishev:1997:NLR

[1003] Lev Beklemishev. Notes on local reflection principles. Theoria, 63(3): 139–146, December 1997. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Buss:1997:BAC

[1004] Samuel R. Buss. Bounded arithmetic, cryptography and complexity. Theoria, 63(3):147–167, December 1997. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Feferman:1997:MRA

[1005] Solomon Feferman. My route to arithmetization. Theoria, 63(3):168–181, December 1997. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Lindstrom:1997:IRT

[1006] Per Lindström. Interpretability in reflexive theories — a survey. Theoria, 63(3):182–209, December 1997. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Shavrukov:1997:IDA

[1007] V. Yu. Shavrukov. Isomorphisms of diagonalizable algebras. Theoria, 63(3):210–221, December 1997. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Gillies:1998:DBT

[1008] Donald Gillies. Debates on Bayesianism and the theory of Bayesian networks. Theoria, 64(1):1–22, April 1998. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ryberg:1998:GRE

[1009] Jesper Ryberg. Generation-relative ethics — a critical note on Dasgupta. Theoria, 64(1):23–33, April 1998. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Olsson:1998:CAL

[1010] Erik J. Olsson. Competing for acceptance Lehrer’s rule and the paradoxes of justification. Theoria, 64(1):34–54, April 1998. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Graham:1998:IBG

[1011] Jody L. Graham. The Intellect’s burden: Geometrical inferences in Descartes’s theory of vision. Theoria, 64(1):55–83, April 1998. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tersman:1998:QE

[1012] Folke Tersman. Quine on ethics. Theoria, 64(1):84–98, April 1998. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1998:R

[1013] Anonymous. Reviews. Theoria, 64(1):99–118, April 1998. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Dummett:1998:TCS

[1014] Michael Dummett. Truth from the constructive standpoint. Theoria, 64 (2–3):122–138, August 1998. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bergstrom:1998:PVV

[1015] Lars Bergström. Prawitz’s version of verificationism. Theoria, 64(2– 3):139–156, August 1998. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Pagin:1998:BMT

[1016] Peter Pagin. Bivalence: meaning theory vs metaphysics. Theoria, 64(2– 3):157–186, August 1998. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sundholm:1998:PAP

[1017] Goran Sundholm. Proofs as acts and proofs as objects: Some questions for Dag Prawitz. Theoria, 64(2–3):187–216, August 1998. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Cellucci:1998:SLD

[1018] Carlo Cellucci. The scope of logic: deduction, abduction, analogy. Theoria, 64(2–3):217–242, August 1998. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Cozzo:1998:ETE

[1019] Cesare Cozzo. Epistemic truth and excluded middle. Theoria, 64(2– 3):243–282, August 1998. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Prawitz:1998:CP

[1020] Dag Prawitz. Comments on the papers. Theoria, 64(2–3):283–337, August 1998. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1998:PLP

[1021] Anonymous. Philosophical and logical publications of Dag Prawitz. Theoria, 64(2–3):338–343, August 1998. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:1999:ECP

[1022] Sven Ove Hansson. Editorial: Cooperation in philosophy. Theoria, 65(1): 1–2, April 1999. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hintikka:1999:QUP

[1023] Jaakko Hintikka. Quine’s ultimate presuppositions. Theoria, 65(1):3– 24, April 1999. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Arrhenius:1999:MAC

[1024] Gustaf Arrhenius. Mutual advantage contractarianism and future generations. Theoria, 65(1):25–35, April 1999. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bykvist:1999:ATP

[1025] Krister Bykvist. All time preferences? Theoria, 65(1):36–54, April 1999. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Munthe:1999:MI

[1026] Christian Munthe. The morality of interference. Theoria, 65(1):55–69, April 1999. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tersman:1999:CCC

[1027] Folke Tersman. Cannibals, Communists and cognitivists. Theoria, 65(1): 70–85, April 1999. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:1999:EPI

[1028] Anonymous. Editorial: Philosophy and interdisciplinary research. Theoria, 65(2–3):87–88, August 1999. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kappel:1999:GEP

[1029] Klemens Kappel. Guest Editor’s preface. Theoria, 65(2–3):89, August 1999. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Lippert-Rasmussen:1999:DSV

[1030] Kasper Lippert-Rasmussen. On denying a significant version of the constancy assumption. Theoria, 65(2–3):90–113, August 1999. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ronnow-Rasmussen:1999:PP

[1031] Toni Rønnow-Rasmussen. Particularism and principles. Theoria, 65(2– 3):114–126, August 1999. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Brannmark:1999:RE

[1032] Johan Brannmark. Rules and exceptions. Theoria, 65(2–3):127–143, August 1999. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Dancy:1999:LMA

[1033] Jonathan Dancy. On the logical and moral adequacy of particularism. Theoria, 65(2–3):144–155, August 1999. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Persson:1999:DDI

[1034] Ingmar Persson. Dancy on desire and internalism of reasons. Theoria, 65 (2–3):156–170, August 1999. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Strandberg:1999:DMP

[1035] Caj Strandberg. Dispositional moral properties and moral motivation. Theoria, 65(2–3):171–192, August 1999. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kappel:1999:DSP

[1036] Klemens Kappel. Has dancy shown a problem in consequentialism? Theoria, 65(2–3):193–211, August 1999. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Dancy:1999:MDA

[1037] Jonathan Dancy. Motivation, dispositions and aims. Theoria, 65(2–3): 212–224, August 1999. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2000:EPL

[1038] Sven Ove Hansson. Editorial: Is philosophy losing ground? Theoria, 66 (1):1–2, April 2000. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hochberg:2000:PTB

[1039] Herbert Hochberg. Propositions, truth and belief: The Wittgenstein-Russell dispute. Theoria, 66(1):3–40, April 2000. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Humberstone:2000:PP

[1040] Lloyd Humberstone. Parts and partitions. Theoria, 66(1):41–82, April 2000. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sternberg:2000:CP

[1041] Fredrik Sternberg. The contingent and the a priori. Theoria, 66(1):83– 85, April 2000. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Needham:2000:RTS

[1042] Paul Needham. Reviews: True Science: Apropos a Recent Collection of Duhem’s Essays. Theoria, 66(1):86–96, April 2000. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Danielsson:2000:RWS

[1043] Sven Danielsson. Reviews: What Shall We Do With Deontic Logic? Theoria, 66(1):97–114, April 2000. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tannsjo:2000:RMW

[1044] Torbjörn Tännsjö. Reviews: Must We, the Reader and I, Make Great Sacrifices In Order to Save Starving Children? Theoria, 66(1):115– 122, April 2000. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hunsson:2000:EPH

[1045] Sven Ove Hunsson. Editorial: Philosophy and its history. Theoria, 66 (2):123–125, August 2000. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Knuuttila:2000:GEP

[1046] Simo Knuuttila. Guest Editor’s preface. Theoria, 66(2):126–128, August 2000. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Boh:2000:FPM

[1047] Ivan Boh. Four phases of medieval epistemic logic. Theoria, 66(2):129– 144, August 2000. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ebbesen:2000:BDS

[1048] Sten Ebbesen. Boethius of Dacia: Science is a serious game. Theoria, 66 (2):145–158, August 2000. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

King:2000:PIM

[1049] Peter King. The problem of individuation in the middle ages. Theoria, 66 (2):159–184, August 2000. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Knuuttila:2000:MBM

[1050] Simo Knuuttila. The medieval background of modern modal conceptions. Theoria, 66(2):185–204, August 2000. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Yrjonsuufu:2000:DOL

[1051] Mikko Yrjönsuufu. Disputations, obligations and logical coherence. Theoria, 66(2):205–223, August 2000. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2000:EWS

[1052] Sven Ove Hansson. Editorial: Who should teach applied ethics? Theoria, 66(3):225–227, December 2000. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Pettit:2000:TDC

[1053] Philip Pettit. Theoria dialogue: Contractualism and consequentialism: A consequentialist perspective on contractualism. Theoria, 66(3):228–236, December 2000. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Scanlon:2000:CR

[1054] T. M. Scanlon. A contractualist reply. Theoria, 66(3):237–245, December 2000. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Carlson:2000:AHS

[1055] Erik Carlson. Aggregating harms — should we kill to avoid headaches? Theoria, 66(3):246–255, December 2000. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tannsjo:2000:ETT

[1056] Torbjörn Tännsjö. The expressivist theory of truth. Theoria, 66(3): 256–272, December 2000. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Stjernberg:2000:RCS

[1057] Fredrik Stjernberg. Reviews: Critical study: Recent work on Frege. Theoria, 66(3):273–288, December 2000. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2000:RRM

[1058] Anonymous. Reviews: R. M. Hare, Sorting out Ethics. Theoria, 66(3): 289–292, December 2000. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2000:RLN

[1059] Anonymous. Reviews: Lennart Nordenfelt, Talking About Health. A Philosophical Dialogue. Theoria, 66(3):293–298, December 2000. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Brinck:2000:RJL

[1060] Ingar Brinck. Reviews: José Luis Bermúdez, The Paradox of Self-Consciousness. Theoria, 66(3):299–306, December 2000. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2000:TBS

[1061] Sven Ove Hansson. A Theoria bibliography: Swedish theses in philosophy 1997–1999. Theoria, 66(3):307–311, December 2000. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2001:ELH

[1062] Sven Ove Hansson. Editorial: Learning from history. Theoria, 67(1): 1–3, April 2001. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Collin:2001:GEP

[1063] Finn Collin. Guest Editor’s preface. Theoria, 67(1):4–6, April 2001. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Son:2001:WAP

[1064] Jerrold Levin Son. Who’s afraid of a paraphrase? Theoria, 67(1):7– 23, April 2001. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

White:2001:LMF

[1065] Roger M. White. Literal meaning and “figurative meaning”. Theoria, 67(1):24–59, April 2001. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Zemach:2001:MTM

[1066] Eddy Zemach. A modal theory of metaphor. Theoria, 67(1):60–74, April 2001. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Collin:2001:MTC

[1067] Finn Collin and Anders Engstrøm. Metaphor and truth-conditional semantics: Meaning as process and product. Theoria, 67(1):75–92, April 2001. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2001:EPU

[1068] Sven Ove Hansson. Editorial: Philosophy as a unifying discipline. Theoria, 67(2):93–95, August 2001. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Maurin:2001:RMI

[1069] Anna-Sofia Maurin and Johannes Persson. Realistic metaphysics an interview with D. H. Mellor. Theoria, 67(2):96–113, August 2001. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Garson:2001:NSW

[1070] James W. Garson. Natural semantics: Why natural deduction is intuitionistic. Theoria, 67(2):114–139, August 2001. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2001:Ra

[1071] Anonymous. Reviews. Theoria, 67(2):140–186, August 2001. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2001:EAP

[1072] Sven Ove Hansson. Editorial: Against philosophical anorexia. Theoria, 67(3):187–188, December 2001. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Freund:2001:FMR

[1073] Michael Freund. Full meet revision on stratified bases. Theoria, 67(3): 189–213, December 2001. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Priest:2001:PBR

[1074] Graham Priest. Paraconsistent belief revision. Theoria, 67(3):214–228, December 2001. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Alter:2001:KHA

[1075] Torin Alter. Know-how, ability, and the ability hypothesis. Theoria, 67(3):229–239, December 2001. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2001:Rb

[1076] Anonymous. Reviews. Theoria, 67(3):240–280, December 2001. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2001:TBS

[1077] Sven Ove Hansson. A Theoria bibliography: Swedish theses in philosophy in the year 2000. Theoria, 67(3):281–285, December 2001. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2002:EHP

[1078] Anonymous. Editorial: How political is philosophy? Theoria, 68(1): 1–3, April 2002. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hugly:2002:TPS

[1079] Philip Hugly and Charles Sayward. There is a problem with substitutional quantification. Theoria, 68(1):4–12, April 2002. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Humberstone:2002:IA

[1080] Lloyd Humberstone. Invitation to autoepistemology. Theoria, 68(1):13– 51, April 2002. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Moreso:2002:LGC

[1081] Jose Juan Moreso, Pablo E. Navarro, and Cristina Redondo. Legal gaps and conclusive reasons. Theoria, 68(1):52–66, April 2002. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Brueckner:2002:BMA

[1082] Anthony Brueckner. Blackburn’s modal argument against moral realism. Theoria, 68(1):67–70, April 2002. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2002:Ra

[1083] Anonymous. Reviews. Theoria, 68(1):71–86, April 2002. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2002:EEM

[1084] Anonymous. Editorial: Educate more philosophers! Theoria, 68(2):89– 90, August 2002. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Gullan-Whur:2002:SEW

[1085] Margaret Gullan-Whur. Spinoza and the equality of women. Theoria, 68(2):91–111, August 2002. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Peterson:2002:APM

[1086] Martin Peterson. An argument for the principle of maximizing expected utility. Theoria, 68(2):112–128, August 2002. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Pritchard:2002:RSE

[1087] Duncan Pritchard. Radical scepticism, epistemological externalism, and closure. Theoria, 68(2):129–161, August 2002. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Raikka:2002:RCW

[1088] Juha Räikkä. The repugnant conclusion and the welfare of actual people. Theoria, 68(2):162–169, August 2002. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Xu:2002:NEG

[1089] Ming Xu. Normal extensions of g.3. Theoria, 68(2):170–176, August 2002. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Beckman:2002:RKF

[1090] Ludvig Beckman. Reviews: Katrin Flikschuh, 2000. Kant and modern political philosophy. Cambridge: Cambridge University. Theoria, 68(2): 177–181, August 2002. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2002:EPR

[1091] Anonymous. Editorial: Peer review in philosophy. Theoria, 68(3):183– 184, December 2002. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Moore:2002:AI

[1092] Matthew E. Moore. Archimedean intuitions. Theoria, 68(3):185–204, December 2002. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Nagasawa:2002:KAA

[1093] Yujin Nagasawa. The knowledge argument against dualism. Theoria, 68(3):205–223, December 2002. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Toribio:2002:MBA

[1094] Josefa Toribio. Mindful belief: Accountability, expertise and cognitive kinds. Theoria, 68(3):224–249, December 2002. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2002:Rb

[1095] Anonymous. Reviews. Theoria, 68(3):250–270, December 2002. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hunsson:2002:TBS

[1096] Sven Ove Hunsson. A Theoria bibliography:swedish theses in philosophy 2001. Theoria, 68(3):271–275, December 2002. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2003:EAP

[1097] Sven Ove Hansson. Editorial: Applying philosophy. Theoria, 69(1–2):1– 3, August 2003. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Holtug:2003:GW

[1098] Nils Holtug. Good for whom? Theoria, 69(1–2):4–20, August 2003. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Temkin:2003:PVI

[1099] Larry S. Temkin. Personal versus impersonal principles: Reconsidering the slogan. Theoria, 69(1–2):21–31, August 2003. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Rasmussen:2003:MDI

[1100] Kasper Lippert Rasmussen. Measuring the disvalue of inequality over time. Theoria, 69(1–2):32–45, August 2003. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Temkin:2003:DSE

[1101] Larry S. Temkin. Determining the scope of egalitarian concern: A partial defense of complete lives egalitarianism. Theoria, 69(1–2):46–59, August 2003. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Rabinowicz:2003:SIB

[1102] Wlodek Rabinowicz. The size of inequality and its badness some reflections around Temkin’s inequality. Theoria, 69(1–2):60–84, August 2003. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Temkin:2003:MIB

[1103] Larry S. Temkin. Measuring inequality’s badness: Does size matter? If so, how, if not, what does? Theoria, 69(1–2):85–108, August 2003. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Persson:2003:BUI

[1104] Ingmar Persson. The badness of unjust inequality. Theoria, 69(1–2): 109–124, August 2003. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Temkin:2003:ERE

[1105] Larry S. Temkin. Exploring the roots of egalitarian concerns. Theoria, 69 (1–2):125–151, August 2003. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2003:EPS

[1106] Sven Ove Hansson. Editorial: Is philosophy science? Theoria, 69(3): 153–156, December 2003. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2003:DLR

[1107] Sven Ove Hansson. David Lewis on rights introduction. Theoria, 69(3): 157–159, December 2003. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Lewis:2003:RR

[1108] David Lewis. Rights to rights. Theoria, 69(3):160–165, December 2003. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Cresswell:2003:NCS

[1109] M. J. Cresswell. Non-contradiction and substantial predication. Theoria, 69(3):166–183, December 2003. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Morrisson:2003:NPC

[1110] Iain Morrisson. Nietzsche’s political confusion. Theoria, 69(3):184–210, December 2003. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Peressini:2003:PRM

[1111] Anthony Peressini. Proof, reliability, and mathematical knowledge. Theoria, 69(3):211–232, December 2003. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2003:R

[1112] Anonymous. Review. Theoria, 69(3):233–253, December 2003. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2003:TBS

[1113] Sven Ove Hansson. A Theoria bibliography: Swedish theses in philosophy 2002. Theoria, 69(3):254–256, December 2003. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2004:Ea

[1114] Anonymous. Editorial. Theoria, 70(1):1–2, April 2004. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Angner:2004:RRT

[1115] Erik Angner. Revisiting Rawls: a theory of justice in the light of Levi’s theory of decision. Theoria, 70(1):3–21, April 2004. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Campbell:2004:CYS

[1116] Scott Campbell. Can you survive a brain-zap? Theoria, 70(1):22–27, April 2004. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Stevens:2004:RPT

[1117] Graham Stevens. From Russell’s paradox to the theory of judgement: Wittgenstein and Russell on the unity of the proposition. Theoria, 70 (1):28–61, April 2004. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Walker:2004:SNC

[1118] Mark Alan Walker. Skepticism and naturalism: Can philosophical skepticism be scientifically tested? Theoria, 70(1):62–97, April 2004. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Johansson:2004:R

[1119] Lars-Göran Johansson. Reviews. Theoria, 70(1):98–105, April 2004. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Olsson:2004:R

[1120] Erik J. Olsson. Reviews. Theoria, 70(1):106–111, April 2004. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Persson:2004:R

[1121] Anders J. Persson. Reviews. Theoria, 70(1):112–116, April 2004. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2004:Eb

[1122] Anonymous. Editorial. Theoria, 70(2–3):117–118, December 2004. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2004:GEP

[1123] Anonymous. Guest Editor’s preface. Theoria, 70(2–3):119–120, December 2004. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

deGaynesford:2004:RO

[1124] Maximilian de Gaynesford. On referring to oneself. Theoria, 70(2–3): 121–161, December 2004. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic). See reply [1125].

Mcdowell:2004:RMG

[1125] John Mcdowell. Reply to Maximilian de Gaynesford. Theoria, 70(2–3): 162–166, December 2004. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic). See [1124].

Gjelsvik:2004:E

[1126] Olav Gjelsvik. Experience. Theoria, 70(2–3):167–191, December 2004. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Mcdowell:2004:ROG

[1127] John Mcdowell. Reply to olav gjelsvik. Theoria, 70(2–3):192–196, December 2004. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Gluer:2004:P

[1128] Kathrin Glüer. On perceiving that. Theoria, 70(2–3):197–212, December 2004. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Mcdowell:2004:RKG

[1129] John Mcdowell. Reply to Kathrin Glüer. Theoria, 70(2–3):213–215, December 2004. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Macbeth:2004:OKM

[1130] Danielle Macbeth. Overcoming Kant: McDowell, and Sellars, on judgment. Theoria, 70(2–3):216–242, December 2004. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic). See reply [1131].

Mcdowell:2004:RDM

[1131] John Mcdowell. Reply to Danielle Macbeth. Theoria, 70(2–3):243–249, December 2004. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic). See [1130].

Macfarlane:2004:MK

[1132] John Macfarlane. McDowell’s Kantianism. Theoria, 70(2–3):250–265, December 2004. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Mcdowell:2004:RJM

[1133] John Mcdowell. Reply to John MacFarlane. Theoria, 70(2–3):266–270, December 2004. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Wikforss:2004:DKO

[1134] Åsa Wikforss. Direct knowledge and other minds. Theoria, 70(2–3): 271–293, December 2004. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Mcdowell:2004:RAW

[1135] John Mcdowell. Reply to Åsa Wikforss. Theoria, 70(2–3):294–297, December 2004. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Brandom:2004:CNB

[1136] Robert Brandom. Critical notice of blind and worried or the sorter resorted. Theoria, 70(2–3):298–302, December 2004. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2004:TBS

[1137] Anonymous. A Theoria bibliography:swedish theses in philosophy 2003. Theoria, 70(2–3):303–305, December 2004. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2005:Ea

[1138] Anonymous. Editorial. Theoria, 71(1):1–2, March 2005. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Maurin:2005:RMI

[1139] Anna-Sofia Maurin and Ingar Brinck. Revisionary metaphysics: An interview with D. M. Armstrong. Theoria, 71(1):3–19, March 2005. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Danielsson:2005:TRP

[1140] Sven Danielsson. Taking Ross’s paradox seriously a note on the original problems of deontic logic. Theoria, 71(1):20–28, March 2005. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hershenov:2005:PPP

[1141] David B. Hershenov. Persons as proper parts of organisms. Theoria, 71(1):29–37, March 2005. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Suikkanen:2005:CRR

[1142] Jussi Suikkanen. Contractualist replies to the redundancy objections. Theoria, 71(1):38–58, March 2005. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Watkins:2005:FPE

[1143] Jeremy Watkins. Forgiveness and its place in ethics. Theoria, 71(1):59– 77, March 2005. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Halldenius:2005:R

[1144] Lena Halldenius. Reviews. Theoria, 71(1):78–84, March 2005. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Olsson:2005:R

[1145] Erik J. Olsson. Reviews. Theoria, 71(1):85–87, March 2005. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2005:Eb

[1146] Anonymous. Editorial. Theoria, 71(2):89–91, June 2005. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Green:2005:CPD

[1147] Karen Green. The context principle and Dummett’s argument for antirealism. Theoria, 71(2):92–117, June 2005. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Jensen:2005:UP

[1148] Karsten Klint Jensen and Kasper Lippert-Rasmussen. Understanding particularism. Theoria, 71(2):118–137, June 2005. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Humberstone:2005:BTP

[1149] Lloyd Humberstone. Béziau’s translation paradox. Theoria, 71(2):138– 181, June 2005. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Nilsson:2005:BTR

[1150] Jonas Nilsson. A bootstrap theory of rationality. Theoria, 71(2):182– 199, June 2005. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2005:EPI

[1151] Sven Ove Hansson. Editorial: Philosophers and intellectuals. Theoria, 71(3):201–202, September 2005. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Olson:2005:RR

[1152] Jonas Olson and Frans Svensson. Regimenting reasons. Theoria, 71(3): 203–214, September 2005. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Loomis:2005:LFP

[1153] Eric J. Loomis. Logical form and propositional function in the Tractatus. Theoria, 71(3):215–240, September 2005. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Humberstone:2005:CSA

[1154] Lloyd Humberstone. Contrariety and subcontrariety: The anatomy of negation (with special reference to an example of J.-Y. Béziau). Theoria, 71(3):241–262, September 2005. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2005:Ra

[1155] Anonymous. Reviews. Theoria, 71(3):263–286, September 2005. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2005:TBT

[1156] Sven Ove Hansson. A Theoria bibliography: A Theoria bibliography. Theoria, 71(3):287–289, September 2005. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2005:EPT

[1157] Sven Ove Hansson. Editorial: Philosophical terminology. Theoria, 71(4): 291–293, December 2005. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Jonsson:2005:TIC

[1158] Martin L. Jönsson and Ingar Brinck. A Theoria interview: Compositionality and other issues in the philosophy of mind and language an interview with Jerry Fodor. Theoria, 71(4):294–308, December 2005. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Miller:2005:NDE

[1159] Kristie Miller. A new definition of endurance. Theoria, 71(4):309–332, December 2005. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Stapleford:2005:TAS

[1160] Scott Stapleford. Transcendental arguments: Superfluity and scepticism. Theoria, 71(4):333–367, December 2005. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2005:Rb

[1161] Anonymous. Reviews. Theoria, 71(4):368–384, December 2005. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2006:EPS

[1162] Sven Ove Hansson. Editorial: Philosophical schools. Theoria, 72(1):1– 4, March 2006. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Almeida:2006:SEA

[1163] Michael J. Almeida. On Stone’s evidential atheism. Theoria, 72(1):5– 22, March 2006. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2006:WEF

[1164] Bengt Hansson. Why explanations? Fundamental, and less fundamental ways of understanding the world. Theoria, 72(1):23–59, March 2006. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Wrenn:2006:EE

[1165] Chase B. Wrenn. Epistemology as engineering? Theoria, 72(1):60–79, March 2006. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sundstrom:2006:RDP

[1166] Pär Sundström. Reviews: David Papineau, Thinking about Consciousness: Oxford: Clarendon Press, 2002, 266 pages, ISBN 0-19-924382-4. Theoria, 72(1):80–86, March 2006. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sunnier:2006:RTT

[1167] L. Wayne Sunnier. Reviews: Torbjörn Tännsjö, Understanding Ethics: An Introduction to Moral Theory. Edinburgh University Press, 2002. vii + 143 pp., ISBN 0-7486-1638-1. Theoria, 72(1):87–90, March 2006. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Rozemond:2006:RLA

[1168] Marleen Rozemond. Reviews: Lilli Alanen, Descartes’s Concept of the Mind, Harvard University Press, 2003, 455 pages, ISBN 0-674-01043-4. Theoria, 72(1):91–95, March 2006. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2006:ECE

[1169] Sven Ove Hansson. Editorial: Condensed examples in philosophy. Theoria, 72(2):97–99, June 2006. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Koslow:2006:RIR

[1170] Arnold Koslow. The representational inadequacy of Ramsey sentences. Theoria, 72(2):100–125, June 2006. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Petersen:2006:RRC

[1171] Thomas Søbirk Petersen. On the repugnance of the repugnant conclusion. Theoria, 72(2):126–137, June 2006. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Spaak:2006:RPS

[1172] Torben Spaak. Reviews: Philip Soper, The Ethics of Deference. Cambridge Studies in Philosophy and Law, Cambridge University Press, 2002, 189 pages, ISBN 0-521-81047-7. Theoria, 72(2):138–147, June 2006. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Miller:2006:RIH

[1173] Jon Miller. Reviews: Ian Hacking, Historical Ontology. Cambridge, MA: Harvard University Press, 2002, vii + 279 pp. ISBN 0-674-00616-X (hardcover). Theoria, 72(2):148–153, June 2006. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hellsten:2006:RPH

[1174] Sirkku K. Hellsten. Reviews: P. H. Coetzee and A. P. J. Roux (eds.), The African Philosophy Reader, 2nd ed. London: Routledge 2003. Theoria, 72(2):154–161, June 2006. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Peterson:2006:RJD

[1175] Martin Peterson. Reviews: Jonathan Dancy, Ethics Without Principles. Oxford University Press 2004, 229 pp. ISBN 0-19-927002-3. Theoria, 72 (2):162–165, June 2006. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Shogenji:2006:RLB

[1176] Toinoji Shogenji. Reviews: Luc Bovens and Stephan Hartmann, Bayesian Epistemology. Oxford University Press, 2004, pp. ix + 159. ISBN 0-19-926975-0 (hardback), ISBN 0-19-927040-6 (paperback). Theoria, 72(2):166–171, June 2006. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2006:EMS

[1177] Sven Ove Hansson. Editorial: The modal status of philosophy. Theoria, 72(3):173–176, September 2006. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bellotti:2006:SSP

[1178] Luca Bellotti. Skolem, the Skolem ‘paradox’ and informal mathematics. Theoria, 72(3):177–212, September 2006. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tannsjo:2006:UTE

[1179] Torbjörn Tännsjö. Understanding through explanation in ethics. Theoria, 72(3):213–220, September 2006. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansen:2006:PDL

[1180] Jörg Hansen. The paradoxes of Deontic logic: Alive and kicking. Theoria, 72(3):221–232, September 2006. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic). See reply [1215].

Tersman:2006:RDD

[1181] Folke Tersman. Review: Davidson, Donald: Problems of Rationality. Oxford: Clarendon, 2004, pp. xx + 280. ISBN 0-19-823754-5. Theoria, 72(3):233–239, September 2006. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Weir:2006:RRF

[1182] Alan Weir. Review: Roger F. Gibson, Jr. (ed.), The Cambridge Companion to Quine, Cambridge: Cambridge University Press, 2004, xx + 323, ISBN 0-521-63056-8. Theoria, 72(3):240–247, September 2006. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ledwig:2006:RAB

[1183] Marion Ledwig. Review: Alexander Broadie (ed.), Thomas Reid on Logic, Rhetoric and the Fine Arts. Edinburgh University Press, Edinburgh (2005), 350 pp. ISBN 0-7486-1684-5 (hardback). Theoria, 72(3): 248–252, September 2006. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Grune-Yanoff:2006:RJB

[1184] Till Grüne-Yanoff. Review: John B. Davis, Alain Marciano and Jochen Runde (eds.), The Elgar Companion to Economics and Philosophy, Edward Elgar (2004), 509 + xxii pp., ISBN 1-84064-964-x. Theoria, 72(3): 253–258, September 2006. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2006:TBT

[1185] Sven Ove Hansson. A Theoria bibliography: A Theoria bibliography: A Theoria bibliography: Swedish theses in philosophy 2005. Theoria, 72(3): 259–262, September 2006. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2006:EDL

[1186] Sven Ove Hansson. Editorial: A dialogue on logic. Theoria, 72(4): 263–268, December 2006. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Borge:2006:CTA

[1187] Steffen Borge. Counterpart theory and the argument from modal concerns. Theoria, 72(4):269–285, December 2006. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bergstrom:2006:QR

[1188] Lars Bergström. Quine’s relativism. Theoria, 72(4):286–298, December 2006. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Dilworth:2006:PIF

[1189] John Dilworth. Perception, introspection and functional consonance. Theoria, 72(4):299–318, December 2006. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Weintraub:2006:WDD

[1190] Ruth Weintraub. What Descartes’ demon can do and his dream cannot. Theoria, 72(4):319–335, December 2006. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Wansing:2006:RDM

[1191] Heinrich Wansing. Review: David Makinson, Bridges from Classical to Nonmonotonic Logic, Texts in Computing, Vol. X.5, King’s College Publications, London, 2005. xvi + 216 pp. ISBN 1-904987-00-1. Theoria, 72(4):336–340, December 2006. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Puustinen:2006:RPJ

[1192] Raimo Puustinen. Review: Philip J. van der Eijk, Medicine and Philosophy in Classical Antiquiw. Doctors and Philosophers on Nature, Soul, Health and Disease. Cambridge University Press, Cambridge 2005, xiv + 404 pp. ISBN 0-521-81800-1. Theoria, 72(4):341–345, December 2006. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Thorseth:2006:RSO

[1193] May Thorseth. Review: Sven Ove Hansson & Elin Palm (eds.): The Ethics of Workplace Privacy, RLE.- Peter Lang, Brussels, 2005. 186 pp. Theoria, 72(4):346–351, December 2006. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hunsson:2007:ERE

[1194] Sven Ove Hunsson. Editorial: Rationality and emotions. Theoria, 73(1): 1–2, March 2007. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Carlson:2007:HVN

[1195] Erik Carlson. Higher values and non-Archimedean additivity. Theoria, 73(1):3–27, March 2007. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Zouhar:2007:DDR

[1196] Marián Zouhar. Definite descriptions, reference, and inference. Theoria, 73(1):28–45, March 2007. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Vessel:2007:PNO

[1197] Jean-Paul Vessel. The probabilistic nature of objective consequentialism. Theoria, 73(1):46–67, March 2007. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2007:RME

[1198] Anonymous. Review: Marc Elliott Bobro, Self and Substance in Leibniz. Theoria, 73(1):68–72, March 2007. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2007:RYB

[1199] Anonymous. Review: Yemima Ben-Menahem (ed). Hilary Putnam. Contemporary Philosophy in Focus. Theoria, 73(1):73–81, March 2007. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2007:RCM

[1200] Anonymous. Review: Christia Mercer and Eileen O’Neill (eds.), Early Modern Philosophy: Mind, Matter, and Metaphysics. Theoria, 73(1):82– 86, March 2007. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2007:REO

[1201] Anonymous. Review: Erik Olsson, Against Coherence: Truth, Probability and Justification. Theoria, 73(1):87–93, March 2007. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2007:EAP

[1202] Sven Ove Hansson. Editorial: Against programmatic ignorance. Theoria, 73(2):95–97, June 2007. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bykvist:2007:VNI

[1203] Krister Bykvist. Violations of normative invariance: Some thoughts on shifty oughts. Theoria, 73(2):98–120, June 2007. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Alm:2007:NCV

[1204] David Alm. Non-cognitivism and validity. Theoria, 73(2):121–147, June 2007. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Clarke:2007:IDN

[1205] David Clarke. Intelligent design: neither scientific nor religious. Theoria, 73(2):148–171, June 2007. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Pritchard:2007:RDP

[1206] Duncan Pritchard. Reviews: Duncan Pritchard, Epistemic Luck. Theoria, 73(2):173–178, June 2007. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2007:RMC

[1207] Anonymous. Reviews: Martha C. Nussbaum, Frontiers of Justice: Disabiliy, Nationality. Theoria, 73(2):179–186, June 2007. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2007:EDD

[1208] Anonymous. Editorial: A dialogue on definitions. Theoria, 73(3):187– 189, September 2007. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bjornsson:2007:MNC

[1209] Gunnar Björnsson and Arvid Båve. Meaning as a normative concept: An interview with Allan Gibbard. Theoria, 73(3):190–206, September 2007. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hilpinen:2007:PA

[1210] Risto Hilpinen. On practical abduction. Theoria, 73(3):207–220, September 2007. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Wreen:2007:SFA

[1211] Michael J. Wreen. A second form of argument from analogy. Theoria, 73(3):221–239, September 2007. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2007:RAG

[1212] Anonymous. Reviews: Anil Gupta, Empiricism and Experience. Theoria, 73(3):240–243, September 2007. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2007:RRS

[1213] Anonymous. Reviews: Richard Sorabji and David Rodin (eds.). The Ethics of War: Shared Problems in Different Traditions. Theoria, 73(3): 244–247, September 2007. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2007:RDH

[1214] Anonymous. Reviews: Douglas Husak and Peter de Marneffe: The Legalization of Drugs. Theoria, 73(3):248–255, September 2007. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Danielsson:2007:RPA

[1215] Sven Danielsson. Ross’s Paradox again: Reply to Hansen. Theoria, 73(3): 256–258, September 2007. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic). See [1180].

Anonymous:2007:TBT

[1216] Anonymous. A Theoria bibliography: A Theoria bibliography: Swedish theses in philosophy 2006. Theoria, 73(3):259–262, September 2007. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2007:TBI

[1217] Anonymous. A Theoria bibliography: Index of Swedish theses in philosophy 1997–2006. Theoria, 73(3):265–279, September 2007. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2007:AA

[1218] Anonymous. Addendum: Addendum. Theoria, 73(3):280, September 2007. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2007:EWP

[1219] Anonymous. Editorial: Why philosophize in English? Theoria, 73(4): 281–283, December 2007. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2007:CT

[1220] Anders Hansson. The concept of tolerance. Theoria, 73(4):284–303, December 2007. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Reisner:2007:ENG

[1221] Andrew E. Reisner. Evidentialism and the numbers game. Theoria, 73(4): 304–316, December 2007. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Johansson:2007:CCE

[1222] Lars-Göran Johansson. Causation in collisions — an empiricist but non-Humean account. Theoria, 73(4):317–333, December 2007. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2007:RDR

[1223] Anonymous. Reviews: D. Rutherford (ed.) The Cambridge Companion to Early Modern Philosophy. Theoria, 73(4):334–339, December 2007. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2007:RFT

[1224] Anonymous. Reviews: Folke Tersman, Moral Disagreement. Theoria, 73(4):340–344, December 2007. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2007:RRC

[1225] Anonymous. Reviews: Robert Callergård, An Essay on Thomas Reid’s Philosophy of Science. Theoria, 73(4):345–353, December 2007. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2007:RSH

[1226] Anonymous. Reviews: Sally Haslanger and Charlotte Witt (eds.): Adoption Matters: Philosophical and Feminist Essays. Theoria, 73(4):354– 358, December 2007. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2008:ENE

[1227] Sven Ove Hansson. Editorial: A new era for Theoria. Theoria, 74(1): 1–2, March 2008. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Lind:2008:TIP

[1228] Andreas Lind and Johan Brännmark. A Theoria interview: Particularism in question: an interview with Jonathan Dancy. Theoria, 74(1):3–17, March 2008. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Rabinowicz:2008:VR

[1229] Wlodek Rabinowicz. Value relations. Theoria, 74(1):18–49, March 2008. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Zangwill:2008:BMD

[1230] Nick Zangwill. Besires and the motivation debate. Theoria, 74(1):50– 59, March 2008. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Grune-Yanoff:2008:AEI

[1231] Till Grüne-Yanoff. Action explanations are not inherently normative. Theoria, 74(1):60–78, March 2008. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Serck-Hanssen:2008:RKH

[1232] Camilla Serck-Hanssen. Reviews: Kant on the Human Standpoint — By Beatrice Longuenesse. Theoria, 74(1):79–85, March 2008. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Girle:2008:RML

[1233] Roderic A. Girle. Reviews: Modal Logic for Philosophers — By James W. Garson. Theoria, 74(1):86–90, March 2008. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sirkel:2008:RAO

[1234] Riin Sirkel. Reviews: Aristotle and Other Platonists — By Lloyd P. Gerson. Theoria, 74(1):91–95, March 2008. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2008:EPP

[1235] Sven Ove Hansson. Editorial: Philosophical plagiarism. Theoria, 74(2): 97–101, May 2008. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Busch:2008:NNM

[1236] Jacob Busch. No new miracles, same old tricks. Theoria, 74(2):102– 114, May 2008. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sobel:2008:HHO

[1237] Jordan Howard Sobel. ‘hoist with his owne petar’: On the undoing of a liar paradox. Theoria, 74(2):115–145, May 2008. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Vahid:2008:RIM

[1238] Hamid Vahid. Radical interpretation and Moore’s paradox. Theoria, 74(2):146–163, May 2008. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Johansson:2008:RFP

[1239] Lars-Göran Johansson. Reviews: A Future for Presentism — By Craig Bourne. Theoria, 74(2):164–168, May 2008. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Brannmark:2008:RMR

[1240] Johan Brännmark. Reviews: Moral Repair: Reconstructing Moral Relations after Wrongdoing — By Margaret Urban Walker. Theoria, 74 (2):169–172, May 2008. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Fleurbaey:2008:REN

[1241] Marc Fleurbaey. Reviews: Egalitarianism. New Essays on the Nature and Value of Equality — Edited by Nils Holtug and Kasper Lippert-Rasmussen. Theoria, 74(2):173–177, May 2008. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2008:EOP

[1242] Sven Ove Hansson. Editorial: The obligations of philosophers. Theoria, 74(3):179–180, September 2008. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kripke:2008:FTS

[1243] Saul A. Kripke. Frege’s theory of sense and reference: Some exegetical notes. Theoria, 74(3):181–218, September 2008. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Feldman:2008:WLS

[1244] Fred Feldman. Whole life satisfaction concepts of happiness. Theoria, 74(3):219–238, September 2008. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Douven:2008:UAS

[1245] Igor Douven and Diederik Olders. Unger’s argument for skepticism revisited. Theoria, 74(3):239–250, September 2008. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2008:TBS

[1246] Sven Ove Hansson. A Theoria bibliography: Swedish theses in philosophy 2007. Theoria, 74(3):251–254, September 2008. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Stjernberg:2008:RKP

[1247] Fredrik Stjernberg. Reviews: The Knowability Paradox — By Jonathan Kvanvig. Theoria, 74(3):255–262, September 2008. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hill:2008:RCC

[1248] Benjamin Hill. Reviews: The Cambridge Companion to Locke’s “Essay Concerning Human Understanding” — Edited by Lex Newman. Theoria, 74(3):263–265, September 2008. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2008:EPM

[1249] Sven Ove Hansson. Editorial: Philosophy as mere rhetoric? Theoria, 74(4):267–270, December 2008. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Aaqvist:2008:SLS

[1250] Lennart Åqvist. Some logico-semantical themes in Karl Olivecrona’s philosophy of law: A non-exegetical approach. Theoria, 74(4):271–294, December 2008. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Cozzo:2008:CTS

[1251] Cesare Cozzo. On the Copernican turn in semantics. Theoria, 74(4): 295–317, December 2008. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Benovsky:2008:TCP

[1252] Jiri Benovsky. Two concepts of possible worlds — or only one? Theoria, 74(4):318–330, December 2008. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Cantwell:2008:CMC

[1253] John Cantwell. Changing the modal context. Theoria, 74(4):331–351, December 2008. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Spaak:2008:RNJ

[1254] Torben Spaak. Reviews: Naturalizing Jurisprudence — By Brian Leiter. Theoria, 74(4):352–362, December 2008. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Brannmark:2008:RNE

[1255] Johan Brännmark. Reviews: Normative Ethics: 5 Questions — Edited by Thomas S. Petersen and Jesper Ryberg. Theoria, 74(4):363–366, December 2008. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Callergaard:2008:RTH

[1256] Robert Callergård. Reviews: A Treatise of Human Nature — a critical edition — By David Hume. Theoria, 74(4):367–368, December 2008. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2009:ESF

[1257] Sven Ove Hansson. Editorial: Seventy-five years of Theoria. Theoria, 75 (1):1, March 2009. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2009:HT

[1258] Sven Ove Hansson. A history of Theoria. Theoria, 75(1):2–27, March 2009. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Brueckner:2009:EPA

[1259] Anthony Brueckner. Endurantism and the psychological approach to personal identity. Theoria, 75(1):28–33, March 2009. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Campbell:2009:WWS

[1260] Neil Campbell. Why we should lower our expectations about the explanatory gap. Theoria, 75(1):34–51, March 2009. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sandu:2009:RDL

[1261] Gabriel Sandu. Reviews: Dependence Logic: A New Approach to Independence Friendly Logic — By Juokko Väänänen. Theoria, 75(1):52–64, March 2009. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sandin:2009:RMP

[1262] Per Sandin. Reviews: 10 Moral Paradoxes — By Saul Smilansky. Theoria, 75(1):65–66, March 2009. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2009:E

[1263] Anonymous. Erratum. Theoria, 75(1):67, March 2009. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2009:EAP

[1264] Sven Ove Hansson. Editorial: Analytic philosophy. Theoria, 75(2):69– 74, May 2009. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2009:TNR

[1265] Sven Ove Hansson. Thomas Nagel — Rolf Schock Prize Laureate. Theoria, 75(2):75, May 2009. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bergstrom:2009:TNR

[1266] Lars Bergström. Thomas Nagel — recipient of the Rolf Schock Prize in Logic and Philosophy, 2008. Theoria, 75(2):76–78, May 2009. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Broome:2009:TNR

[1267] John Broome. Thomas Nagel — Rolf Schock Prize Laureate: Motivation. Theoria, 75(2):79–99, May 2009. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Lotfi:2009:WRF

[1268] Shidan Lotfi. Wittgenstein’s rule-following considerations and moral particularism. Theoria, 75(2):100–116, May 2009. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Mcbride:2009:KCU

[1269] Mark Mcbride. Is knowledge closed under known entailment? The strange case of Hawthorne’s “heavyweight conjunct” (and other strange cases). Theoria, 75(2):117–128, May 2009. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ruloff:2009:ESI

[1270] Colin Ruloff. Epistemic supervenience and internalism: A trilemma. Theoria, 75(2):129–151, May 2009. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2009:EBS

[1271] Sven Ove Hansson. Editorial: Bentham and the shoemaker. Theoria, 75(3):153–155, September 2009. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2009:TBS

[1272] Sven Ove Hansson. A Theoria bibliography: Swedish theses in philosophy 2008. Theoria, 75(3):156–160, September 2009. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Klausen:2009:TNE

[1273] Søren Harnow Klausen. Two notions of epistemic normativity. Theoria, 75(3):161–178, September 2009. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Zelenak:2009:ERS

[1274] Eugen Zeleňák. On explanatory relata in singular causal explanation. Theoria, 75(3):179–195, September 2009. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Johansson:2009:S

[1275] Jens Johansson. Am i a series? Theoria, 75(3):196–205, September 2009. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Stapleford:2009:LSK

[1276] Scott Stapleford. Locke on sensitive knowledge as knowledge. Theoria, 75(3):206–231, September 2009. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kim:2009:MRD

[1277] Sungsu Kim. Multiple realizations, diverse implementations and antireductionism. Theoria, 75(3):232–244, September 2009. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Brueckner:2009:MPP

[1278] Anthony Brueckner. Moore-paradoxicality and the principle of charity. Theoria, 75(3):245–247, September 2009. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2009:EPT

[1279] Sven Ove Hansson. Editorial: Philosophy and the two cultures. Theoria, 75(4):249–251, December 2009. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Rosenkrantz:2009:TTD

[1280] Max Rosenkrantz. The Tractatus theory of descriptions. Theoria, 75(4): 252–271, December 2009. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Seidman:2009:VC

[1281] Jeffrey Seidman. Valuing and caring. Theoria, 75(4):272–303, December 2009. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hauska:2009:DU

[1282] Jan Hauska. Dispositions unmasked. Theoria, 75(4):304–335, December 2009. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Northcott:2009:LSN

[1283] Robert Northcott. On Lewis, Schaffer and the non-reductive evaluation of counterfactuals. Theoria, 75(4):336–343, December 2009. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tahko:2009:MCP

[1284] Tuomas E. Tahko. On the modal content of a posteriori necessities. Theoria, 75(4):344–357, December 2009. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Brannmark:2009:RCA

[1285] Johan Brännmark. Reviews: The Constitution of Agency: Essays on Practical Reason and Moral Psychology — By Christine M. Korsgaard.

Theoria, 75(4):358–361, December 2009. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hill:2009:RTS

[1286] David W. Hill. Reviews: Trouble with Strangers: A Study of Ethics — By Terry Eagleton. Theoria, 75(4):362–365, December 2009. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Penco:2009:RBS

[1287] Carlo Penco. Reviews: Between Saying and Doing: Towards an Analytic Pragmatism — By Robert B. Brandom. Theoria, 75(4):366–369, December 2009. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2010:EWM

[1288] Sven Ove Hansson. Editorial: Women and minorities in philosophy. Theoria, 76(1):1–3, March 2010. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Dzhafarov:2010:SVC

[1289] Ehtibar N. Dzhafarov and Damir D. Dzhafarov. Sorites without vagueness I: Classificatory Sorites. Theoria, 76(1):4–24, March 2010. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Dzhafarov:2010:SVI

[1290] Ehtibar N. Dzhafarov and Damir D. Dzhafarov. Sorites without vagueness II: Comparative sorites. Theoria, 76(1):25–53, March 2010. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

OSullivan:2010:TRS

[1291] Brendan O’Sullivan. Taking referentialism seriously: A response to the modal argument. Theoria, 76(1):54–67, March 2010. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Mikkonen:2010:LFU

[1292] Jukka Mikkonen. Literary fictions as utterances and artworks. Theoria, 76(1):68–90, March 2010. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ronzitti:2010:NUM

[1293] Giuseppina Ronzitti. A note on an unpublished manuscript by Erik Stenius. Theoria, 76(1):91–96, March 2010. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2010:EWP

[1294] Sven Ove Hansson. Editorial: Wedberg on philosophical analysis. Theoria, 76(2):97–99, May 2010. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Vaananen:2010:OMP

[1295] Jouko Väänänen and Dag Westerståhl. Obituary: In memoriam: Per Lindström. Theoria, 76(2):100–107, May 2010. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Manninen:2010:LEE

[1296] Juha Manninen and Johan Strang. Letter to the Editor: Eino Kaila’s political views: A comment on Sven Ove Hansson’s “A History of Theoria”. Theoria, 76(2):108–111, May 2010. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Podlaskowski:2010:UTD

[1297] Adam C. Podlaskowski. Unbelievable thoughts and doxastic oughts. Theoria, 76(2):112–118, May 2010. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Carlson:2010:PD

[1298] Erik Carlson. Parity demystified. Theoria, 76(2):119–128, May 2010. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hendley:2010:AWE

[1299] Steven Hendley. Answerable to the world: Experience and practical intentionality in Brandom’s and McDowell’s “intramural” debate. Theoria, 76(2):129–151, May 2010. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Famakinwa:2010:MCI

[1300] J. O. Famakinwa. The moderate communitarian individual and the primacy of duties. Theoria, 76(2):152–166, May 2010. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Wolff:2010:TBPa

[1301] Ernst Wolff. Technicity of the body as part of the socio-technical system: the contributions of Mauss and Bourdieu. Theoria, 76(2):167–187, May 2010. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2010:EMP

[1302] Sven Ove Hansson. Editorial: Methodological pluralism in philosophy. Theoria, 76(3):189–191, September 2010. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Rabinowicz:2010:OMJ

[1303] Wlodek Rabinowicz. Obituary: In memoriam: Jordan Howard Sobel (1929–2010). Theoria, 76(3):192–196, September 2010. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Turri:2010:DPE

[1304] John Turri. Does perceiving entail knowing? Theoria, 76(3):197–206, September 2010. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Campbell:2010:EEI

[1305] Neil Campbell. Explanatory exclusion and the intensionality of explanation. Theoria, 76(3):207–220, September 2010. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Williams:2010:MPD

[1306] John N. Williams. Moore’s paradox, defective interpretation, justified belief and conscious belief. Theoria, 76(3):221–248, September 2010. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Cresswell:2010:AEC

[1307] M. J. Cresswell. Abstract entities in the causal order. Theoria, 76(3): 249–265, September 2010. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2010:TBS

[1308] Sven Ove Hansson. A Theoria bibliography: Swedish theses in philosophy 2009. Theoria, 76(3):266–269, September 2010. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

VanDitmarsch:2010:RNE

[1309] Hans Van Ditmarsch. Reviews: New Essays on the Knowability Paradox — Edited by Joe Salerno. Theoria, 76(3):270–273, September 2010. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Johansson:2010:RPA

[1310] Linda Johansson. Reviews: The Philosophy of Animal Minds — Edited by Robert W. Lurz. Theoria, 76(3):274–279, September 2010. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Mansnerus:2010:RET

[1311] Erika Mansnerus. Reviews: The Ethics of Technological Risk — Edited by Lotte Asveld and Sabine Roeser. Theoria, 76(3):280–283, September 2010. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2010:ERT

[1312] Sven Ove Hansson. Editorial: The right to be technical. Theoria, 76(4): 285–286, December 2010. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Botterill:2010:TKC

[1313] George Botterill. Two kinds of causal explanation. Theoria, 76(4):287– 313, December 2010. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Horta:2010:DTM

[1314] Oscar Horta. Discrimination in terms of moral exclusion. Theoria, 76(4): 314–332, December 2010. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Wolff:2010:TBPb

[1315] Ernst Wolff. Technicity of the body as part of the socio-technical system: the contributions of Mauss and Bourdieu. Theoria, 76(4):333–354, December 2010. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Martens:2010:WHC

[1316] David B. Martens. William heytesbury and the conditions for knowledge. Theoria, 76(4):355–374, December 2010. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Wolenski:2010:RTO

[1317] Jan Woleński. Reviews: The Theory of Ontic Modalities — By Uwe Meixner. Theoria, 76(4):375–379, December 2010. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Cook:2010:RVD

[1318] Roy T. Cook. Reviews: Vagueness and Degrees of Truth — By Nicholas J. J. Smith. Theoria, 76(4):380–384, December 2010. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2011:EPD

[1319] Sven Ove Hansson. Editorial: Philosophy in the defence of science. Theoria, 77(1):1–3, March 2011. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Chen:2011:ITW

[1320] Bo Chen. An interview with Timothy Williamson. Theoria, 77(1):4– 31, March 2011. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

LeMorvan:2011:KIT

[1321] Pierre Le Morvan. Knowledge, ignorance and true belief. Theoria, 77 (1):32–41, March 2011. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Huckfeldt:2011:PCI

[1322] Vaughn Huckfeldt. Prudence, commitments and intertemporal conflicts. Theoria, 77(1):42–54, March 2011. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Petersson:2011:AHH

[1323] Bo Petersson. Axel Hägerström and his early version of error theory. Theoria, 77(1):55–70, March 2011. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tsai:2011:LKH

[1324] Cheng-Hung Tsai. Linguistic know-how: The limits of intellectualism. Theoria, 77(1):71–86, March 2011. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Wolenski:2011:RBL

[1325] Jan Woleński. Reviews: Der Begriff der logischen Form in der analytischen Philosophie. Russell in Auseinandersetzung mit Frege, Meinong und Wittgenstein — By Elena Tatievskaya. Theoria, 77(1):87–89, March 2011. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Petersson:2011:RRM

[1326] Bo Petersson. Reviews: A Real Mind. The Life and Work of Axel Hägerström — By Patricia Mindus. Theoria, 77(1):90–99, March 2011. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2011:EWO

[1327] Sven Ove Hansson. Editorial: Writing our own history. Theoria, 77 (2):101–103, May 2011. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2011:TRT

[1328] Bengt Hansson, Hans Van Ditmarsch, Pascal Engel, Sven Ove Hansson, Vincent Hendricks, Søren Holm, Pauline Jacobson, Anthonie Meijers, Henry S. Richardson, and Hans Rott. A Theoria round table on philosophy publishing. Theoria, 77(2):104–116, May 2011. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Jespersen:2011:TCT

[1329] Bjørn Jespersen and Massimiliano Carrara. Two conceptions of technical malfunction. Theoria, 77(2):117–138, May 2011. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Busch:2011:IAD

[1330] Jacob Busch. Is the indispensability argument dispensable? Theoria, 77(2):139–158, May 2011. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Boudry:2011:EPP

[1331] Maarten Boudry and Filip Buekens. The epistemic predicament of a pseudoscience: Social constructivism confronts Freudian psychoanalysis. Theoria, 77(2):159–179, May 2011. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Carrara:2011:WKS

[1332] Massimiliano Carrara and Davide Fassio. Why knowledge should not be typed: An argument against the type solution to the knowability paradox. Theoria, 77(2):180–193, May 2011. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2011:ETA

[1333] Sven Ove Hansson. Editorial: The topics of applied ethics. Theoria, 77(3):195–197, September 2011. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Arrhenius:2011:ET

[1334] Gustaf Arrhenius, Ingar Brinck, Kathrin Glüer-Pagin, Lena Halldenius, Anna-Sofia Maurin, Folke Tersman, and Åsa Wikforss. To the Editor of Theoria. Theoria, 77(3):198, September 2011. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Khetrapal:2011:ET

[1335] Neha Khetrapal. To the Editor of Theoria. Theoria, 77(3):198–200, September 2011. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2011:LEE

[1336] Sven Ove Hansson. Letters to the Editor: The Editor’s reply. Theoria, 77(3):200, September 2011. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Jager:2011:PRV

[1337] Christoph Jäger. Process reliabilism and the value problem. Theoria, 77(3):201–213, September 2011. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic). See reply [1338] and response [1339].

Olsson:2011:KLL

[1338] Erik J. Olsson and Martin L. Jönsson. Kinds of learning and the likelihood of future true beliefs: Reply to Jäger on reliabilism and the value problem. Theoria, 77(3):214–222, September 2011. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic). See [1337] and reply [1339].

Jager:2011:RFT

[1339] Christoph Jäger. Reliability and future true belief: Reply to Olsson and Jönsson. Theoria, 77(3):223–237, September 2011. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic). See [1339].

Svensson:2011:HWB

[1340] Frans Svensson. Happiness, well-being, and their relation to virtue in Descartes’ ethics. Theoria, 77(3):238–260, September 2011. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Callanan:2011:MSD

[1341] John. J. Callanan. Making sense of doubt: Strawson’s anti-scepticism. Theoria, 77(3):261–278, September 2011. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2011:TBS

[1342] Sven Ove Hansson. A Theoria bibliography: Swedish theses in philosophy 2010. Theoria, 77(3):279–281, September 2011. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Grattan-Guinness:2011:REL

[1343] I. Grattan-Guinness. Reviews: The Evolution of Logic — By W. D. Hart. Theoria, 77(3):282–283, September 2011. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Johansson:2011:RPI

[1344] Jens Johansson. Reviews: Persons, Interests, and Justice — By Nils Holtug. Theoria, 77(3):284–287, September 2011. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2011:EE

[1345] Anonymous. ERRATUM: Erratum. Theoria, 77(3):i, September 2011. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2011:EIT

[1346] Sven Ove Hansson. Editorial: Information technology and the organization of philosophical research. Theoria, 77(4):289–291, December 2011. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Trueman:2011:PFE

[1347] Robert Trueman. Propositional functions in extension. Theoria, 77(4): 292–311, December 2011. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Buckareff:2011:AID

[1348] Andrei A. Buckareff. Action-individuation and doxastic agency. Theoria, 77(4):312–332, December 2011. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Roberts:2011:AS

[1349] Melinda A. Roberts. The asymmetry: A solution. Theoria, 77(4):333– 367, December 2011. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Brueckner:2011:IS

[1350] Anthony Brueckner. Idealism and scepticism. Theoria, 77(4):368–371, December 2011. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2012:ELL

[1351] Sven Ove Hansson. Editorial: From Latin to linguistic confusion to English: Language shifts in philosophy. Theoria, 78(1):1–5, February 2012. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Rickless:2012:WHF

[1352] Samuel C. Rickless. Why and how to fill an unfilled proposition. Theoria, 78(1):6–25, February 2012. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tropman:2012:CCM

[1353] Elizabeth Tropman. Can Cornell moral realism adequately account for moral knowledge? Theoria, 78(1):26–46, February 2012. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic). See reply [1419].

Baumberger:2012:IA

[1354] Christoph Baumberger and Georg Brun. Identities of artefacts. Theoria, 78(1):47–74, February 2012. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Wolfendale:2012:RMD

[1355] Jessica Wolfendale. Reviews: Moral Dilemmas of Modern War: Torture, Assassination, and Blackmail in an Age of Asymmetric Conflict — By Michael L. Gross. Theoria, 78(1):75–79, February 2012. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Angere:2012:RPI

[1356] Staffan Angere. Reviews: The Philosophy of Information — By Luciano Floridi. Theoria, 78(1):80–83, February 2012. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Moberger:2012:RRN

[1357] Victor Moberger. Reviews: From Reasons to Norms: On the Basic Question in Ethics — By Torbjörn Tännsjö. Theoria, 78(1):84–88, February 2012. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2012:EPPa

[1358] Sven Ove Hansson. Editorial: Philosophy and public policy. Theoria, 78(2):89–92, May 2012. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

VanDitmarsch:2012:EKH

[1359] Hans Van Ditmarsch, Wiebe Van Der Hoek, and Petar Iliev. Everything is knowable — how to get to know whether a proposition is true. Theoria, 78(2):93–114, May 2012. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

French:2012:DPS

[1360] Craig French. Does propositional seeing entail propositional knowledge? Theoria, 78(2):115–127, May 2012. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Keskinen:2012:QOR

[1361] Antti Keskinen. Quine on objects: Realism or anti-realism? Theoria, 78(2):128–145, May 2012. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Klement:2012:FCC

[1362] Kevin C. Klement. Frege’s changing conception of number. Theoria, 78(2):146–167, May 2012. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Jerkert:2012:RSA

[1363] Jesper Jerkert. Review: Science in the Age of Computer Simulation — By Eric Winsberg. Theoria, 78(2):168–175, May 2012. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2012:TBS

[1364] Sven Ove Hansson. A Theoria bibliography: Swedish theses in philosophy 2011. Theoria, 78(3):177–180, September 2012. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2012:EPPb

[1365] Sven Ove Hansson. Editorial: Progress in philosophy? A dialogue. Theoria, 78(3):181–185, September 2012. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Berkovski:2012:PMH

[1366] Sandy Berkovski. The possibility of modified hedonism. Theoria, 78(3): 186–212, September 2012. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kasaa:2012:ELM

[1367] Martin Kaså. Experimental logics, mechanism and knowable consistency. Theoria, 78(3):213–224, September 2012. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bo:2012:DRK

[1368] Chen Bo. A descriptivist refutation of Kripke’s modal argument and of Soames’s defence. Theoria, 78(3):225–260, September 2012. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bjornberg:2012:REP

[1369] Karin Edvardsson Björnberg. Reviews: Ethics and Public Policy: A Philosophical Inquiry — By Jonathan Wolff. Theoria, 78(3):261–267, September 2012. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ekendahl:2012:RPV

[1370] Karl Ekendahl. Reviews: Personal Value — By Toni Rønnow-Rasmussen. Theoria, 78(3):268–272, September 2012. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2012:EPPc

[1371] Sven Ove Hansson. Editorial: Precision in philosophy. Theoria, 78(4): 273–275, December 2012. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Duke:2012:AST

[1372] George Duke. Abstract singular terms and thin reference. Theoria, 78(4): 276–292, December 2012. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Olinder:2012:RDN

[1373] Ragnar Francén Olinder. Rescuing doxastic normativism. Theoria, 78(4): 293–308, December 2012. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Zhong:2012:ECM

[1374] Lei Zhong. An explanatory challenge to moral reductionism. Theoria, 78(4):309–325, December 2012. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Wolfson:2012:AKA

[1375] Ben Wolfson. Agential knowledge, action and process. Theoria, 78(4): 326–357, December 2012. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Neuber:2012:ISM

[1376] Matthias Neuber. Invariance, structure, measurement — Eino Kaila and the history of logical empiricism. Theoria, 78(4):358–383, December 2012. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2013:EWP

[1377] Sven Ove Hansson. Editorial: Welcome to philosophyland. Theoria, 79 (1):1–7, February 2013. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Baltimore:2013:CPM

[1378] Joseph A. Baltimore. Careful, physicalists: Mind–body supervenience can be too superduper. Theoria, 79(1):8–21, February 2013. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Echeverri:2013:PSR

[1379] Santiago Echeverri. Is perception a source of reasons? Theoria, 79 (1):22–56, February 2013. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Raven:2013:LMT

[1380] Michael J. Raven. Is Lewis’s mixed theory mixed up? Theoria, 79 (1):57–75, February 2013. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sonderholm:2013:UIN

[1381] Jorn Sonderholm. Unreliable intuitions: A new reply to the moral twin-Earth argument. Theoria, 79(1):76–88, February 2013. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Moller:2013:RPP

[1382] Niklas Möller. Review: The Price of Precaution and the Ethics of Risk — By Christian Munthe. Theoria, 79(1):89–92, February 2013. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2013:EPI

[1383] Sven Ove Hansson. Editorial: Philosophy and intercultural dialogue. Theoria, 79(2):93–95, May 2013. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Konigsberg:2013:PUS

[1384] Amir Konigsberg. The problem with uniform solutions to peer disagreement. Theoria, 79(2):96–126, May 2013. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Miller:2013:IOD

[1385] Christian Miller. Identifying with our desires. Theoria, 79(2):127–154, May 2013. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Buford:2013:DIM

[1386] Christopher T. Buford. Does indeterminacy matter? Theoria, 79(2):155– 166, May 2013. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Turri:2013:TO

[1387] John Turri. That’s outrageous. Theoria, 79(2):167–171, May 2013. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bergstrom:2013:RCC

[1388] Lars Bergström. Reviews: Confessions of a Confirmed Extensionalist and Other Essays — By W. V. Quine. Theoria, 79(2):172–179, May 2013. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Jerkert:2013:RPE

[1389] Jesper Jerkert. Reviews: The Philosophy of Evidence-Based Medicine — By Jeremy Howick. Theoria, 79(2):180–186, May 2013. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2013:ETC

[1390] Sven Ove Hansson. Editorial: Ten commandments for journal referees. Theoria, 79(3):187–188, September 2013. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2013:HPR

[1391] Sven Ove Hansson. Hilary Putnam — Rolf Schock Prize Laureate. Theoria, 79(3):189, September 2013. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Pagin:2013:HPR

[1392] Peter Pagin. Hilary Putnam — Rolf Schock Laureate: Introduction. Theoria, 79(3):190–191, September 2013. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Putnam:2013:HPR

[1393] Hilary Putnam. Hilary Putnam — Rolf Schock Laureate: The development of externalist semantics. Theoria, 79(3):192–203, September 2013. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Boyd:2013:HPR

[1394] Richard Boyd. Hilary Putnam — Rolf Schock Laureate: Semantic externalism and knowing our own minds: Ignoring twin-Earth and doing naturalistic philosophy. Theoria, 79(3):204–228, September 2013. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Burge:2013:HPR

[1395] Tyler Burge. Hilary Putnam — Rolf Schock Laureate: Some remarks on Putnam’s contributions to semantics. Theoria, 79(3):229–241, September 2013. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Wikforss:2013:HPR

[1396] Åsa Wikforss. Hilary Putnam — Rolf Schock Laureate: Bachelors, energy, cats and water: Putnam on kinds and kind terms. Theoria, 79(3): 242–261, September 2013. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Reynolds:2013:ESH

[1397] Steven L. Reynolds. Effective sceptical hypotheses. Theoria, 79(3):262– 278, September 2013. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Steglich-Petersen:2013:NGA

[1398] Asbjørn Steglich-Petersen. The no guidance argument. Theoria, 79(3): 279–283, September 2013. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2013:TBS

[1399] Sven Ove Hansson. A Theoria bibliography: Swedish theses in philosophy 2012: A Theoria bibliography. Theoria, 79(3):284–286, September 2013. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2013:EFP

[1400] Sven Ove Hansson. Editorial: Formal philosophy — a guarded defence. Theoria, 79(4):287–289, December 2013. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Puryear:2013:ISR

[1401] Stephen Puryear. Idealism and scepticism: A reply to Brueckner. Theoria, 79(4):290–293, December 2013. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Atkinson:2013:TPS

[1402] David Atkinson and Jeanne Peijnenburg. Transitivity and partial screening off. Theoria, 79(4):294–308, December 2013. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Deery:2013:ALP

[1403] Oisín Deery. Absences and late preemption. Theoria, 79(4):309–325, December 2013. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Batchelor:2013:CTC

[1404] Roderick Batchelor. Complexes and their constituents. Theoria, 79(4): 326–352, December 2013. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tripodi:2013:MKM

[1405] Paolo Tripodi. A myth to kill a myth? On McDowell’s interpretation of Sellars’ empiricism and the philosophy of mind. Theoria, 79(4):353–377, December 2013. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Janvid:2013:RES

[1406] Mikael Janvid. Review: Essays on Skepticism — By Anthony Brueckner. Theoria, 79(4):378–382, December 2013. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2014:EBE

[1407] Sven Ove Hansson. Editorial: Beyond “experimental philosophy”. Theoria, 80(1):1–3, February 2014. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Moore:2014:CPR

[1408] Dwayne Moore. On the constitutive property reply: Commentary on Campbell. Theoria, 80(1):4–25, February 2014. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Rubin:2014:TRM

[1409] Michael Rubin. On two responses to moral twin Earth. Theoria, 80 (1):26–43, February 2014. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Freitag:2014:SSD

[1410] Wolfgang Freitag. Safety, sensitivity and “distant” epistemic luck. Theoria, 80(1):44–61, February 2014. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic). See [1468].

Breakey:2014:PML

[1411] Hugh Breakey. Parsing macpherson: The last rites of Locke the possessive individualist. Theoria, 80(1):62–83, February 2014. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Salay:2014:DRP

[1412] Nancy Salay. Dress rehearsals, previews, and encores: A new account of mental representation. Theoria, 80(1):84–97, February 2014. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Brueckner:2014:MIS

[1413] Anthony Brueckner. More on idealism and skepticism. Theoria, 80(1): 98–99, February 2014. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sahlin:2014:US

[1414] Nils-Eric Sahlin and Paul Weirich. Unsharp sharpness. Theoria, 80(1): 100–103, February 2014. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Brennan:2014:RIR

[1415] Geoffrey Brennan. Review: Ideology, Rationality, and Revolution: An Essay on the Persistence of Oppression — By Nicolas Olsson-Yaouzis. Theoria, 80(1):104–112, February 2014. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2014:ENS

[1416] Sven Ove Hansson. Editorial: The normative stance. Theoria, 80(2):113– 115, May 2014. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

vanBenthem:2014:PSD

[1417] Johan van Benthem, Davide Grossi, and Fenrong Liu. Priority structures in Deontic logic. Theoria, 80(2):116–152, May 2014. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Baron:2014:CID

[1418] Sam Baron. Can indispensability-driven Platonists be (serious) presentists? Theoria, 80(2):153–173, May 2014. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Long:2014:DCR

[1419] Joseph Long. In defence of Cornell realism: a reply to Elizabeth Tropman. Theoria, 80(2):174–183, May 2014. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic). See [1353].

Tropman:2014:WCM

[1420] Elizabeth Tropman. Why Cornell moral realism cannot provide an adequate account of moral knowledge. Theoria, 80(2):184–190, May 2014. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Torpman:2014:RCM

[1421] Olle Torpman. Reviews: Climate Matters — Ethics in a Warming World — By John Broome. Theoria, 80(2):191–195, May 2014. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Zackariasson:2014:RPP

[1422] Ulf Zackariasson. Reviews: Pragmatic Pluralism and the Problem of God — By Sami Pihlström. Theoria, 80(2):196–200, May 2014. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2014:EPL

[1423] Sven Ove Hansson. Editorial: Philosophy and linguistic relativity. Theoria, 80(3):201–204, September 2014. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Uberti:2014:GBI

[1424] Luca Jacopo Uberti. Good and bad idealizations in political theory. Theoria, 80(3):205–231, September 2014. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Connelly:2014:RWL

[1425] James Connelly. Russell and Wittgenstein on logical form and judgement: What did Wittgenstein try that wouldn’t work? Theoria, 80(3):232–254, September 2014. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Friedrich:2014:BM

[1426] Daniel Friedrich. Belief and motivation. Theoria, 80(3):255–268, September 2014. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2014:TBS

[1427] Sven Ove Hansson. A Theoria bibliography: Swedish theses in philosophy 2013. Theoria, 80(3):269–271, September 2014. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Brulde:2014:RHG

[1428] Bengt Brülde. Reviews: Happiness and the Good Life — By Mike W. Martin. Theoria, 80(3):272–276, September 2014. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kauss:2014:RNB

[1429] Dominik Kauss. Reviews: Necessary Beings — An Essay on Ontology, Modality, and the Relations Between Them — By Bob Hale. Theoria, 80(3):277–282, September 2014. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bergstrom:2014:RQV

[1430] Lars Bergström. Reviews: Quine versus Davidson: Truth, Reference, and Meaning — By Gary Kemp. Theoria, 80(3):283–287, September 2014. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2014:EPO

[1431] Sven Ove Hansson. Editorial: Predatory open access. Theoria, 80(4): 289–291, December 2014. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Dieveney:2014:QMD

[1432] Patrick Dieveney. Quantification and metaphysical discourse. Theoria, 80(4):292–318, December 2014. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Aaqvist:2014:DTL

[1433] Lennart Åqvist. Deontic tense logic with historical necessity, frame constants, and a solution to the epistemic obligation paradox (the “knower”). Theoria, 80(4):319–349, December 2014. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Barz:2014:IGM

[1434] Wolfgang Barz. Introspection as a game of make-believe. Theoria, 80(4): 350–367, December 2014. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Johansson:2014:RPS

[1435] Lars-Göran Johansson. Reviews: Philosophy of Science. A Unified Approach — By Gerhard Schurz. Theoria, 80(4):368–376, December 2014. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Svensson:2014:RML

[1436] Frans Svensson. Reviews: Meaning in Life: An Analytic Study — By Thaddeus Metz. Theoria, 80(4):377–384, December 2014. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2015:EPJ

[1437] Sven Ove Hansson. Editorial: Is philosophy just a set of empty ideas? Theoria, 81(1):1–3, February 2015. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Eilan:2015:RRN

[1438] Naomi Eilan. A relational response to Newman’s objection to Russell’s causal theory of perception. Theoria, 81(1):4–26, February 2015. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Williams:2015:EDM

[1439] John N. Williams. Eliminativism, dialetheism and Moore’s paradox. Theoria, 81(1):27–47, February 2015. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Skiba:2015:ISR

[1440] Lukas Skiba. On indirect sense and reference. Theoria, 81(1):48–81, February 2015. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Evers:2015:RMN

[1441] Daan Evers. Reviews: Meaning and Normativity — By Allan Gibbard. Theoria, 81(1):82–86, February 2015. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Gustafsson:2015:RMF

[1442] Johan E. Gustafsson. Reviews: Measures of Freedom of Choice — By Karin Enflo. Theoria, 81(1):87–92, February 2015. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2015:EED

[1443] Sven Ove Hansson. Editorial: The ethics of doing philosophy. Theoria, 81(2):93–96, May 2015. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Steglich-Petersen:2015:KAL

[1444] Asbjørn Steglich-Petersen. Knowing the answer to a loaded question. Theoria, 81(2):97–125, May 2015. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Buekens:2015:DSL

[1445] Filip Buekens and Maarten Boudry. The dark side of the loon. explaining the temptations of obscurantism. Theoria, 81(2):126–142, May 2015. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Naibo:2015:UDI

[1446] Alberto Naibo and Mattia Petrolo. Are uniqueness and deducibility of identicals the same? Theoria, 81(2):143–181, May 2015. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Lundgren:2015:RRE

[1447] Björn Lundgren. Review: Reference and Existence: The John Locke Lectures — By Saul A. Kripke. Theoria, 81(2):182–188, May 2015. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2015:EPD

[1448] Sven Ove Hansson. Editorial: Philosophy — a domain for the white male genius? Theoria, 81(3):189–191, September 2015. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Pailos:2015:SMP

[1449] Federico Pailos and Lucas Rosenblatt. Solving multimodal paradoxes. Theoria, 81(3):192–210, September 2015. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tallant:2015:PTN

[1450] Jonathan Tallant. Presentism, truthmaking and necessary connections. Theoria, 81(3):211–221, September 2015. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Garthoff:2015:PPR

[1451] Jon Garthoff. The priority and posteriority of right. Theoria, 81(3): 222–248, September 2015. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Gert:2015:PPP

[1452] Joshua Gert. Parity, preference and puzzlement. Theoria, 81(3):249–271, September 2015. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Gluer:2015:SNG

[1453] Kathrin Glüer and Åsa Wikforss. Still no guidance: Reply to steglich-Petersen. Theoria, 81(3):272–279, September 2015. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2015:TBS

[1454] Sven Ove Hansson. A Theoria bibliography: Swedish theses in philosophy 2014. Theoria, 81(3):280–282, September 2015. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2015:EWC

[1455] Sven Ove Hansson. Editorial: “Who can write my dissertation for me?”. Theoria, 81(4):283–288, December 2015. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sandu:2015:OJH

[1456] Gabriel Sandu. Obituary: Jaakko Hintikka in memoriam. Theoria, 81(4): 289–292, December 2015. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Green:2015:FEN

[1457] Karen Green. Frege on existence and non-existence. Theoria, 81(4): 293–310, December 2015. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kraft:2015:EDG

[1458] Tim Kraft. Epistemological Disjunctivism’s genuine access problem. Theoria, 81(4):311–332, December 2015. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ashton:2015:UUB

[1459] Natalie Alana Ashton. Undercutting underdetermination-based scepticism. Theoria, 81(4):333–354, December 2015. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Horgan:2015:EIM

[1460] Terry Horgan and Mark Timmons. Exploring intuitions on moral twin Earth: A reply to sonderholm. Theoria, 81(4):355–375, December 2015. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Palm:2015:RRA

[1461] Elin Palm. Reviews: Rightlessness in an Age of Rights: Hannah Arendt and the Contemporary Struggles of Migrants — By Ayten Gündoğdu. Theoria, 81(4):376–379, December 2015. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ahlin:2015:REI

[1462] Jesper Ahlin. Reviews: The Ethics of Immigration — By Joseph H. Carens. Theoria, 81(4):380–384, December 2015. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2016:EPF

[1463] Sven Ove Hansson. Editorial: Philosophy and freedom of speech. Theoria, 82(1):1–2, February 2016. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Allo:2016:LRR

[1464] Patrick Allo. Logic, reasoning and revision. Theoria, 82(1):3–31, February 2016. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Inkpin:2016:PRP

[1465] Andrew Inkpin. Projection, recognition, and pictorial diversity. Theoria, 82(1):32–55, February 2016. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Rosenkrantz:2016:PFB

[1466] Max Rosenkrantz. The problem of false belief and the failure of the theory of descriptions. Theoria, 82(1):56–80, February 2016. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Avnur:2016:IR

[1467] Yuval Avnur. On an irrelevant regress. Theoria, 82(1):81–88, February 2016. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Broncano-Berrocal:2016:NLD

[1468] Fernando Broncano-Berrocal. No luck in the distance: A reply to Freitag. Theoria, 82(1):89–100, February 2016. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic). See [1410] and reply [1488].

Hansson:2016:EPP

[1469] Sven Ove Hansson. Editorial: Progress in philosophy — a centennial perspective. Theoria, 82(2):101–103, May 2016. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Rabinowicz:2016:DPR

[1470] Wlodek Rabinowicz. Derek Parfit — Rolf Schock Laureate: Derek Parfit’s contributions to philosophy. Theoria, 82(2):104–109, May 2016. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Parfit:2016:DPR

[1471] Derek Parfit. Derek Parfit — Rolf Schock Laureate: Can we avoid the repugnant conclusion? Theoria, 82(2):110–127, May 2016. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Temkin:2016:DPR

[1472] Larry S. Temkin. Derek Parfit — Rolf Schock Laureate: On the nature and structure of self-interest, morality and practical reasoning. Theoria, 82(2):128–147, May 2016. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Persson:2016:DPR

[1473] Ingmar Persson. Derek Parfit — Rolf Schock Laureate: Parfit on personal identity: Its analysis and (un)importance. Theoria, 82(2):148–165, May 2016. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Arrhenius:2016:DPR

[1474] Gustaf Arrhenius. Derek Parfit — Rolf Schock Laureate: Population ethics and different-number-based imprecision. Theoria, 82(2):166–181, May 2016. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Chang:2016:DPR

[1475] Ruth Chang. Derek Parfit — Rolf Schock Laureate: Parity, imprecise comparability and the repugnant conclusion. Theoria, 82(2):182–214, May 2016. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2016:EZA

[1476] Sven Ove Hansson. Editorial: Zombie arguments and the progress of philosophy. Theoria, 82(3):215–216, September 2016. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Dellsen:2016:ERU

[1477] Finnur Dellsén. Explanatory rivals and the ultimate argument. Theoria, 82(3):217–237, September 2016. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Archer:2016:ASA

[1478] Alfred Archer. Are acts of supererogation always praiseworthy? Theoria, 82(3):238–255, September 2016. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Monsere:2016:CPA

[1479] Annelies Monseré. The charge from psychology and Art’s definition. Theoria, 82(3):256–273, September 2016. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ho:2016:ENP

[1480] Tsung-Hsing Ho. Epistemic normativity as performance normativity. Theoria, 82(3):274–284, September 2016. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2016:TBS

[1481] Sven Ove Hansson. A Theoria bibliography: Swedish theses in philosophy 2015. Theoria, 82(3):285–287, September 2016. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ahteensuu:2016:RAA

[1482] Marko Ahteensuu. Reviews: Animals and African Ethics — By Kai Horsthemke. Theoria, 82(3):288–294, September 2016. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Cabezas:2016:REC

[1483] Mar Cabezas. Reviews: Emotion — By Carolyn Price. Theoria, 82(3): 295–298, September 2016. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2016:EEE

[1484] Sven Ove Hansson. Editorial: Ethical expertise. Theoria, 82(4):299–301, December 2016. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Lockhart:2016:FAP

[1485] Thomas Lockhart. Frege on anti-psychologism and the role of logic in thinking. Theoria, 82(4):302–328, December 2016. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

McGrew:2016:BFA

[1486] Lydia McGrew. Bayes factors all the way: Toward a new view of coherence and truth. Theoria, 82(4):329–350, December 2016. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Klev:2016:PTA

[1487] Ansten Klev. A proof-theoretic account of the miners paradox. Theoria, 82(4):351–369, December 2016. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Freitag:2016:DSR

[1488] Wolfgang Freitag. Distance from safety! A reply to Broncano-Berrocal’s “No Luck in the Distance”. Theoria, 82(4):370–373, December 2016. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic). See [1468].

Lagerlund:2016:RAH

[1489] Henrik Lagerlund. Reviews: Axel Hägerström and Modern Social Thought — Edited by Sven Eliaeson, Patricia Mindus, Stephen P. Turner. Theoria, 82(4):374–378, December 2016. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sandqvist:2016:RCC

[1490] Tor Sandqvist. Reviews: Constructivity and Computability in Historical and Philosophical Perspective — Edited by Jaques Dubucs and Michel Bourdeau. Theoria, 82(4):379–382, December 2016. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2017:EPU

[1491] Sven Ove Hansson. Editorial: The paradox of useful research. Theoria, 83(1):1–3, February 2017. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Meylan:2017:CCD

[1492] Anne Meylan. The consequential conception of doxastic responsibility. Theoria, 83(1):4–28, February 2017. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Novaes:2017:RBN

[1493] Catarina Dutilh Novaes and Herman Veluwenkamp. Reasoning biases, non-monotonic logics and belief revision. Theoria, 83(1):29–52, February 2017. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Paoletti:2017:FNP

[1494] Michele Paolini Paoletti. From falsemakers to negative properties. Theoria, 83(1):53–77, February 2017. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Blancke:2017:WDI

[1495] Stefaan Blancke, Maarten Boudry, and Massimo Pigliucci. Why do irrational beliefs mimic science? The cultural evolution of pseudoscience. Theoria, 83(1):78–97, February 2017. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2017:EWS

[1496] Sven Ove Hansson. Editorial: Who should be author? Theoria, 83(2): 99–102, May 2017. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Thorn:2017:ADC

[1497] Paul D. Thorn. Against deductive closure. Theoria, 83(2):103–119, May 2017. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Logins:2017:CSE

[1498] Artūrs Logins. Common sense and evidence: Some neglected arguments in favour of E = K. Theoria, 83(2):120–137, May 2017. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Besch:2017:SMR

[1499] Thomas M. Besch. Schmidtz on moral recognition rules: A critique. Theoria, 83(2):138–153, May 2017. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Cantwell:2017:MCR

[1500] John Cantwell, Sten Lindström, and Wlodek Rabinowicz. McGee’s counterexample to the Ramsey test. Theoria, 83(2):154–168, May 2017. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2017:EUM

[1501] Sven Ove Hansson. Editorial: The uses and misuses of philosophical scepticism. Theoria, 83(3):169–174, September 2017. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bonardi:2017:NTB

[1502] Paolo Bonardi. Names of truth bearers, and “that”-clauses: A dilemma for Millians. Theoria, 83(3):175–184, September 2017. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Morganti:2017:CRP

[1503] Matteo Morganti and Attila Tanyi. Can reasons be propositions? Against Dancy’s attack on propositionalism. Theoria, 83(3):185–205, September 2017. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Vahid:2017:DCP

[1504] Hamid Vahid. Deontological conservatism and perceptual justification. Theoria, 83(3):206–224, September 2017. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Gruber:2017:RPA

[1505] Monika Gruber. Ramsey’s pragmatic approach to truth and belief. Theoria, 83(3):225–248, September 2017. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Thorell:2017:RPS

[1506] Amanda Thorell. Reviews: Philosophy of Science for Scientists — By Lars–Göran Johansson. Theoria, 83(3):249–253, September 2017. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Palm:2017:RRE

[1507] Elin Palm. Reviews: Refugees and the Ethics of Forced Displacement — By Serena Parekh. Theoria, 83(3):254–261, September 2017. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bulow:2017:RDT

[1508] William Bülow. Reviews: Democratic Theory and Mass Incarceration — Edited by Albert W. Dzur, Ian Loader, and Richard Sparks. Theoria, 83(3):262–267, September 2017. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2017:TBS

[1509] Sven Ove Hansson. A Theoria bibliography: Swedish theses in philosophy 2016. Theoria, 83(3):268–272, September 2017. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2017:TBI

[1510] Sven Ove Hansson. A Theoria bibliography: Index of Swedish theses in philosophy 2007–2016. Theoria, 83(3):273–293, September 2017. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2017:EMP

[1511] Sven Ove Hansson. Editorial: Mistakes in philosophy. Theoria, 83(4): 295–297, December 2017. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Lipski:2017:DIM

[1512] Joachim Lipski. Demotivating intentional mentalism. Theoria, 83(4): 298–318, December 2017. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Loschke:2017:FVI

[1513] Jörg Löschke. Friendship, value and interpretation. Theoria, 83(4):319– 340, December 2017. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Elson:2017:IVB

[1514] Luke Elson. Incommensurability as vagueness: a burden-shifting argument. Theoria, 83(4):341–363, December 2017. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Demircioglu:2017:DNE

[1515] Erhan Demircioğlu. Dretske on non-epistemic seeing. Theoria, 83(4): 364–393, December 2017. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Dicher:2017:GCP

[1516] Bogdan Dicher. On a generality condition in proof-theoretic semantics. Theoria, 83(4):394–418, December 2017. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Besong:2017:DC

[1517] Brian Besong. Disagreeing with confidence. Theoria, 83(4):419–439, December 2017. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Pascucci:2017:ARD

[1518] Matteo Pascucci. Anderson’s restriction of Deontic modalities to contingent propositions. Theoria, 83(4):440–470, December 2017. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Zamani:2017:TSA

[1519] Mohsen Zamani. The totality of states of affairs and the minimal truthmaker. Theoria, 83(4):471–483, December 2017. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Olsson:2017:REC

[1520] Erik J. Olsson. Review: Epistemic Contextualism: A Defense — By Peter Baumann. Theoria, 83(4):484–488, December 2017. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2018:II

[1521] Anonymous. Issue information. Theoria, 84(1):??, February 2018. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2018:EWP

[1522] Sven Ove Hansson. Editorial: Why publish at all? Theoria, 84(1): 1–3, February 2018. CODEN THRAA5. ISSN 0040-5825 (print), 1755-

2567 (electronic). URL https://onlinelibrary.wiley.com/doi/abs/
10.1111/theo.12141.

Sawyer:2018:SE

[1523] Sarah Sawyer. Subjective externalism. Theoria, 84(1):4–22, February 2018. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/abs/10.1111/
theo.12137.

Hauser:2018:PIR

[1524] Kai Hauser and Tahsİn Öner. Perception, intuition, and reliability. Theoria, 84(1):23–59, February 2018. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic). URL https://onlinelibrary.wiley.

com/doi/abs/10.1111/theo.12140.

Cohen:2018:MGM

[1525] Shlomo Cohen. The moral gradation of media of deception. Theoria, 84 (1):60–82, February 2018. CODEN THRAA5. ISSN 0040-5825 (print),

1755-2567 (electronic). URL https://onlinelibrary.wiley.com/doi/
abs/10.1111/theo.12139.

Hutt:2018:RDD

[1526] Donald Bello Hutt. Republicanism, deliberative democracy, and equality of access and deliberation. Theoria, 84(1):83–111, February 2018. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic). URL

https://onlinelibrary.wiley.com/doi/abs/10.1111/theo.12138.

Hansson:2018:APC

[1527] Sven Ove Hansson. Anonymous philosophical communication. Theoria, 84(2):113–119, May 2018. CODEN THRAA5. ISSN 0040-5825 (print),

1755-2567 (electronic). URL https://onlinelibrary.wiley.com/doi/
abs/10.1111/theo.12149.

Rast:2018:PD

[1528] Erich H. Rast. Perspectival disagreement. Theoria, 84(2):120–139, May 2018. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/abs/10.1111/
theo.12143.

Archer:2018:TCD

[1529] Avery Archer. Trying cognitivism: A defence of the strong belief thesis. Theoria, 84(2):140–156, May 2018. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic). URL https://onlinelibrary.wiley.

com/doi/abs/10.1111/theo.12144.

McEvoy:2018:ANS

[1530] Mark McEvoy. Apriority, necessity and the subordinate role of empirical warrant in mathematical knowledge. Theoria, 84(2):157–178, May 2018. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic). URL

https://onlinelibrary.wiley.com/doi/abs/10.1111/theo.12145.

Rattan:2018:DCU

[1531] Gurpreet Rattan. Disagreement and conceptual understanding. Theoria, 84(2):179–210, May 2018. CODEN THRAA5. ISSN 0040-5825 (print),

1755-2567 (electronic). URL https://onlinelibrary.wiley.com/doi/
abs/10.1111/theo.12150.

Korngut:2018:BRE

[1532] Michael Korngut. Book review: Erik Nis Ostenfeld, Human Wisdom. Studies in Ancient Greek Philosophy (Academia Philosophical Studies Vol. 53). Academia Verlag, Sankt Augustin, 2016. 382 pp. ISBN 978-3-89665-694-0. EUR39.00. Theoria, 84(2):211–216, May 2018. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic). URL https:

//onlinelibrary.wiley.com/doi/abs/10.1111/theo.12148.

Colomina:2018:HCT

[1533] Juan J. Colomina. Book review: Herman Cappelen, Tamar S. Gendler and John Hawthorne (eds), The Oxford Handbook of Philosophical Methodology. Oxford University Press, Oxford, 2016. xvi + 752 pp. ISBN 978-0-19-966877-9. $150.00 (hardback). Theoria, 84(2):217–219, May 2018. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/abs/10.1111/
theo.12147.

Hansson:2018:WPR

[1534] Sven Ove Hansson. What is philosophy, really? Theoria, 84(3):221–227, September 2018. CODEN THRAA5. ISSN 0040-5825 (print), 1755-

2567 (electronic). URL https://onlinelibrary.wiley.com/doi/abs/
10.1111/theo.12163.

Hansson:2018:IRM

[1535] Sven Ove Hansson. Introduction: Ruth Millikan — Rolf Schock Prize Laureate. Theoria, 84(3):228, September 2018. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic). URL https:

//onlinelibrary.wiley.com/doi/abs/10.1111/theo.12162.

Millikan:2018:BWD

[1536] Ruth Garrett Millikan. Biosemantics and words that don’t represent. Theoria, 84(3):229–241, September 2018. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic). URL https://onlinelibrary.

wiley.com/doi/abs/10.1111/theo.12146.

Borg:2018:MMM

[1537] Emma Borg. Millikan, meaning, and minimalism. Theoria, 84(3):242– 258, September 2018. CODEN THRAA5. ISSN 0040-5825 (print), 1755-

2567 (electronic). URL https://onlinelibrary.wiley.com/doi/abs/
10.1111/theo.12164.

Bjornsson:2018:SCR

[1538] Gunnar Björnsson. Strategic content: Representations of epistemic modality in biosemantics (and success semantics). Theoria, 84(3):259– 277, September 2018. CODEN THRAA5. ISSN 0040-5825 (print), 1755-

2567 (electronic). URL https://onlinelibrary.wiley.com/doi/abs/
10.1111/theo.12165.

Hansson:2018:STP

[1539] Sven Ove Hansson. Swedish theses in philosophy 2017. Theoria, 84(3): 278–280, September 2018. CODEN THRAA5. ISSN 0040-5825 (print),

1755-2567 (electronic). URL https://onlinelibrary.wiley.com/doi/
abs/10.1111/theo.12161.

Hansson:2018:PAA

[1540] Sven Ove Hansson. Philosophy after the aprioristic parenthesis? Theoria, 84(4):281–283, December 2018. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Pollock:2018:PFD

[1541] Henry Pollock. Parfit’s fission dilemma: Why relation R doesn’t matter. Theoria, 84(4):284–294, December 2018. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Wrigley:2018:SOI

[1542] Wesley Wrigley. Sider’s ontologese introduction instructions. Theoria, 84(4):295–308, December 2018. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Chan:2018:ECS

[1543] Lok-Chi Chan. Emergentism and the contingent solubility of salt. Theoria, 84(4):309–324, December 2018. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Andersson:2018:MPC

[1544] Henrik Andersson and Anders Herlitz. A more plausible collapsing principle. Theoria, 84(4):325–336, December 2018. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bognar:2018:FPD

[1545] Greg Bognar. Fairness and the puzzle of disability. Theoria, 84(4): 337–355, December 2018. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2019:IIa

[1546] Anonymous. Issue information. Theoria, 85(1):1–2, February 2019. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2019:CT

[1547] Sven Ove Hansson. Critical thinking. Theoria, 85(1):3–7, February 2019. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Lajevardi:2019:ATS

[1548] Kaave Lajevardi and Saeed Salehi. On the arithmetical truth of selfreferential sentences. Theoria, 85(1):8–17, February 2019. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bradley:2019:CTI

[1549] Seamus Bradley. A counterexample to three imprecise decision theories. Theoria, 85(1):18–30, February 2019. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Balletta:2019:MW

[1550] Sandro Balletta. The metaphysics of words. Theoria, 85(1):31–48, February 2019. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Palm:2019:KEM

[1551] Elin Palm. Book review: Kevin Macnish, The Ethics of Surveillance — An Introduction. Routledge, New York, 2017. 216 pp. ISBN 978-1-138-64378-9. Theoria, 85(1):49–57, February 2019. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2019:IIb

[1552] Anonymous. Issue information. Theoria, 85(2):59–60, April 2019. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2019:PPU

[1553] Sven Ove Hansson. Philosophical plagiarism under the spotlight. Theoria, 85(2):61–68, April 2019. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Johansson:2019:EVL

[1554] Lars-Göran Johansson. An empiricist view on laws, quantities and physical necessity. Theoria, 85(2):69–101, April 2019. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tallant:2019:YCT

[1555] Jonathan Tallant. You can trust the ladder, but you shouldn’t. Theoria, 85(2):102–118, April 2019. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Demircioglu:2019:RGP

[1556] Erhan Demircioglu. Reliabilism, the generality problem, and the basing relation. Theoria, 85(2):119–144, April 2019. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Livadas:2019:PIP

[1557] Stathis Livadas. The plausible impact of phenomenology on Gödel’s thoughts. Theoria, 85(2):145–170, April 2019. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Whiston:2019:HAA

[1558] Amna Whiston. Book review: Hanno Sauer, Moral Judgments as Educated Intuitions. Cambridge, MA, MIT Press, 2017. 312 pp. ISBN 978-0-262-03560-6. $50.00 (hbk). Theoria, 85(2):171–178, April 2019. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2019:IIc

[1559] Anonymous. Issue information. Theoria, 85(3):179–180, June 2019. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2019:SAB

[1560] Sven Ove Hansson. Stop the assault on Brazilian philosophy and sociology! Theoria, 85(3):181–184, June 2019. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Torrengo:2019:FF

[1561] Giuliano Torrengo and Samuele Iaquinto. Flow fragmentalism. Theoria, 85(3):185–201, June 2019. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Belkoniene:2019:NUP

[1562] Miloud Belkoniene. Normalcy, understanding and the problem of statistical evidence. Theoria, 85(3):202–218, June 2019. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Dougherty:2019:CPC

[1563] M. V. Dougherty. The corruption of philosophical communication by translation plagiarism. Theoria, 85(3):219–246, June 2019. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic). See comments [1573].

Hansson:2019:STP

[1564] Sven Ove Hansson. Swedish theses in philosophy 2018. Theoria, 85(3): 247–250, June 2019. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2019:IId

[1565] Anonymous. Issue information. Theoria, 85(4):251–252, August 2019. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2019:AWB

[1566] Sven Ove Hansson. Academic whistle-blowing. Theoria, 85(4):253–257, August 2019. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Haack:2019:PPT

[1567] Susan Haack. Post “post-truth”: Are we there yet? Theoria, 85(4): 258–275, August 2019. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Assadian:2019:IFG

[1568] Bahram Assadian and Jonathan Nassim. Indeterminacy and failure of grounding. Theoria, 85(4):276–288, August 2019. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Barnes:2019:ALD

[1569] Eric Christian Barnes. An argument for the law of desire. Theoria, 85 (4):289–311, August 2019. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Valente:2019:CDD

[1570] Matheus Valente. Communicating and disagreeing with distinct concepts: A defense of semantic internalism. Theoria, 85(4):312–336, August 2019. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2019:IIe

[1571] Anonymous. Issue information. Theoria, 85(5):337–338, October 2019. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2019:ANA

[1572] Sven Ove Hansson. Academic and non-academic freedom of speech. Theoria, 85(5):339–343, October 2019. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2019:FDA

[1573] Sven Ove Hansson. Further developments after Professor Dougherty’s recent article in Theoria. Theoria, 85(5):344, October 2019. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic). See [1563].

Frandberg:2019:LAM

[1574] Åke Frändberg, Risto Hilpinen, and Lars Lindahl. Lennart Åqvist in memoriam. Theoria, 85(5):345–349, October 2019. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bach:2019:DAE

[1575] Theodore Bach. In defence of armchair expertise. Theoria, 85(5):350–382, October 2019. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Travaglini:2019:IKM

[1576] Graziella Travaglini. Imagination and knowledge in the metaphorology of Paul Ricoeur. Theoria, 85(5):383–401, October 2019. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sauchelli:2019:PIT

[1577] Andrea Sauchelli. Personal identity and trivial survival. Theoria, 85(5): 402–411, October 2019. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2019:IIf

[1578] Anonymous. Issue information. Theoria, 85(6):413–414, December 2019. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2019:CS

[1579] Sven Ove Hansson. Cheating students. Theoria, 85(6):415–419, December 2019. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hilpinen:2019:SFR

[1580] Risto Hilpinen. On the sick father, the repentant sinner, and other problems in medieval Deontic logic. Theoria, 85(6):420–434, December 2019. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Strossner:2019:NFL

[1581] Corina Strössner. The normative force of logical and probabilistic reasoning in improving beliefs. Theoria, 85(6):435–458, December 2019. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ceragioli:2019:PCH

[1582] Leonardo Ceragioli. Peano’s counterexample to harmony. Theoria, 85(6): 459–484, December 2019. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2020:IIa

[1583] Anonymous. Issue information. Theoria, 86(1):1–2, February 2020. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2020:DRE

[1584] Sven Ove Hansson. Does research ethics apply to us? Theoria, 86(1):3–8, February 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Palm:2020:EMG

[1585] Elin Palm. Externalized migration governance and the limits of sovereignty: The case of partnership agreements between EU and Libya. Theoria, 86(1):9–27, February 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Duarte:2020:ECC

[1586] Melina Duarte. The ethical consequences of criminalizing solidarity in the EU. Theoria, 86(1):28–53, February 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Malfatti:2020:CTT

[1587] Federica I. Malfatti. Can testimony transmit understanding? Theoria, 86(1):54–72, February 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic). See reply [1614].

Enflo:2020:MS

[1588] Karin Enflo. Measures of similarity. Theoria, 86(1):73–99, February 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Vollet:2020:RI

[1589] Jacques-Henri Vollet. Refined invariantism. Theoria, 86(1):100–127, February 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Jerkert:2020:BRB

[1590] Jesper Jerkert. Book review: Broadbent, Alex, Philosophy of Medicine. Oxford University Press, New York, 2019. xx + 274 pp. ISBN 978-0-19-061214-6 (pbk). Theoria, 86(1):128–136, February 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2020:IIb

[1591] Anonymous. Issue information. Theoria, 86(2):137–138, April 2020. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2020:PE

[1592] Sven Ove Hansson. Philosophical expertise. Theoria, 86(2):139–144, April 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Vogel:2020:RLF

[1593] Christopher A. Vogel. Realism and lexical flexibility. Theoria, 86(2):145– 186, April 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Lando:2020:EMU

[1594] Giorgio Lando. Expansionism and mereological universalism. Theoria, 86(2):187–219, April 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Calle:2020:PPP

[1595] Olga Ramirez Calle. Patterns, patterns, patterns: Art and meaning at the crossroads between two opposing forces. Theoria, 86(2):220–244, April 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Giananti:2020:PCG

[1596] Andrea Giananti. Perception, content, generality. Theoria, 86(2):245– 267, April 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Taylor:2020:ITP

[1597] Jamie Taylor. An inconsistent triad: Priority pluralism, perdurantism and (the possibility of) gunky time. Theoria, 86(2):268–285, April 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2020:IIc

[1598] Anonymous. Issue information. Theoria, 86(3):287–288, June 2020. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2020:WCW

[1599] Sven Ove Hansson. What can we demand of a referee report? Theoria, 86(3):289–292, June 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tilley:2020:MEM

[1600] John J. Tilley. Moral explanations of moral beliefs: Inappropriate to demand them? Theoria, 86(3):293–308, June 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Pozza:2020:PLE

[1601] Carlo Dalla Pozza, Claudio Garola, Antonio Negro, and Davide Sergio. A pragmatic logic for expressivism. Theoria, 86(3):309–340, June 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kemp:2020:UAP

[1602] Gary Kemp. A unified account: Pictorial, photographic and sculptural seeing as spectral seeing. Theoria, 86(3):341–358, June 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Francen:2020:RME

[1603] Ragnar Francén. Reconsidering the meta-ethical implications of motivational internalism and externalism. Theoria, 86(3):359–388, June 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ori:2020:WRE

[1604] Meshi Ori. Why not road ethics? Theoria, 86(3):389–412, June 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Penco:2020:BRJ

[1605] Carlo Penco. Book review: John Perry, Frege’s Detour. An Essay on Meaning, Reference, and Truth. Oxford: Oxford University Press, 2019, pp. 148. ISBN 978-0-19-881282-1. Theoria, 86(3):413–424, June 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2020:STP

[1606] Sven Ove Hansson. Swedish theses in philosophy 2019. Theoria, 86(3): 425–428, June 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Fairley:2020:RSA

[1607] Neil Hamilton Fairley. Retracted: De Se attitudes and computation. Theoria, 86(3):429, June 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2020:IId

[1608] Anonymous. Issue information. Theoria, 86(4):431–432, August 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2020:DMP

[1609] Sven Ove Hansson. Do moral philosophers have to be moral? Theoria, 86 (4):433–438, August 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ciecierski:2020:USU

[1610] Tadeusz Ciecierski. Utterances, sub-utterances and token-reflexivity. Theoria, 86(4):439–462, August 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Herlitz:2020:SAP

[1611] Anders Herlitz. Spectrum arguments, parity and persistency. Theoria, 86 (4):463–481, August 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Fan:2020:FKC

[1612] Jie Fan. A family of Kripke contingency logics. Theoria, 86(4):482–499, August 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Fassio:2020:BCS

[1613] Davide Fassio and Jie Gao. Belief, credence and statistical evidence. Theoria, 86(4):500–527, August 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Gilbertson:2020:UTR

[1614] Eric Gilbertson. Understanding by testimony: A reply to Malfatti. Theoria, 86(4):528–534, August 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic). See [1587].

Anonymous:2020:IIe

[1615] Anonymous. Issue information. Theoria, 86(5):535–536, October 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2020:PBL

[1616] Sven Ove Hansson. The philosophy of Black Lives Matter. Theoria, 86 (5):537–542, October 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Rota:2020:DS

[1617] Michael Rota. On the definition of slavery. Theoria, 86(5):543–564, October 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Lin:2020:UAI

[1618] Szu-Yen Lin. The unavailability of authorial intent. Theoria, 86(5): 565–582, October 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ehring:2020:WPC

[1619] Douglas Ehring. Why Parfit can rebut Johnston’s Reductio. Theoria, 86 (5):583–594, October 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Moberger:2020:BPP

[1620] Victor Moberger. Bullshit, pseudoscience and pseudophilosophy. Theoria, 86(5):595–611, October 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Pils:2020:QSR

[1621] Raimund Pils. Quine’s scientific realism revisited. Theoria, 86(5):612– 642, October 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Zahnoun:2020:TA

[1622] Farid Zahnoun. Truth or accuracy? Theoria, 86(5):643–650, October 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Puncochar:2020:ICO

[1623] Vít Puncochár and Christopher Gauker. Indicative conditionals in objective contexts. Theoria, 86(5):651–687, October 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Gordon:2020:BRJ

[1624] John-Stewart Gordon. Book review: Joschka Haltaufderheide, Johanna Hovemann and Jochen Vollmann (eds.), Ageing Between Participation and Simulation. De Gruyter, Berlin, 2020, 244 pp. ISBN 978-3-11-067740-9. Theoria, 86(5):688–691, October 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2020:IIf

[1625] Anonymous. Issue information. Theoria, 86(6):693–694, December 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2020:DP

[1626] Sven Ove Hansson. Disguised plagiarism. Theoria, 86(6):695–703, December 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Mindus:2020:WPV

[1627] Patricia Mindus and Elena Prats. Women in philosophy: Voices from Scandinavia — an introduction. Theoria, 86(6):704–708, December 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Campeggiani:2020:NCS

[1628] Pia Campeggiani. Nec cogitare sed facere: The paradox of fiction at the tribunal of ancient poetics. Theoria, 86(6):709–726, December 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Rasmussen:2020:IBD

[1629] Katharina Berndt Rasmussen. Implicit bias and discrimination. Theoria, 86(6):727–748, December 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bjornberg:2020:WIA

[1630] Karin Edvardsson Björnberg. What, if anything, is wrong with offsetting nature? Theoria, 86(6):749–768, December 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sampson:2020:CTR

[1631] Kristin Sampson. Conceptions of temporality: Reconsidering time in an age of impending emergency. Theoria, 86(6):769–782, December 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Evers:2020:CBB

[1632] Kathinka Evers. The culture-bound brain: Epigenetic proaction revisited. Theoria, 86(6):783–800, December 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Pepp:2020:DAA

[1633] Jessica Pepp. Is Dickie’s account of aboutness-fixing explanatory? Theoria, 86(6):801–820, December 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bagnoli:2020:PKE

[1634] Carla Bagnoli. Practical knowledge, equal standing, and proper reliance on others. Theoria, 86(6):821–842, December 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Schellekens:2020:EAM

[1635] Elisabeth Schellekens. Evaluating art morally. Theoria, 86(6):843–858, December 2020. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2021:IIa

[1636] Anonymous. Issue information. Theoria, 87(1):1–2, February 2021. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2021:PAR

[1637] Sven Ove Hansson. Philosophy and alternative realities. Theoria, 87 (1):3–6, February 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ciecierski:2021:ICC

[1638] Tadeusz Ciecierski and Paweł Grabarczyk. An introduction: Concepts, contexts and communication. Theoria, 87(1):7–12, February 2021. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Zalta:2021:UTN

[1639] Edward N. Zalta. Unifying three notions of concepts. Theoria, 87(1): 13–30, February 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bloch-Mullins:2021:SRI

[1640] Corinne L. Bloch-Mullins. Similarity reimagined (with implications for a theory of concepts). Theoria, 87(1):31–68, February 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

McAllister:2021:CCA

[1641] Blake McAllister. Conceptualism and concept acquisition. Theoria, 87 (1):69–86, February 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Trajkovski:2021:CAO

[1642] Miroslava Trajkovski. The constitution of abstract objects. Theoria, 87 (1):87–108, February 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Perrin:2021:RDR

[1643] Denis Perrin. Reference by deference: The real semiotic profile of indexicals and their context. Theoria, 87(1):109–135, February 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Georgi:2021:CCU

[1644] Geoff Georgi. Contexts and constraints on use. Theoria, 87(1):136–151, February 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Odrowaz-Sypniewska:2021:FGD

[1645] Joanna Odrowaz-Sypniewska. Faultless and genuine disagreement over vague predicates. Theoria, 87(1):152–166, February 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Valente:2021:SCI

[1646] Matheus Valente. On successful communication, intentions and false beliefs. Theoria, 87(1):167–186, February 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Rudnicki:2021:CPW

[1647] Jakub Rudnicki. Centred propositions, what is asserted, and communication. Theoria, 87(1):187–206, February 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Fairley:2021:SAC

[1648] Neil Hamilton Fairley. De Se attitudes and computation. Theoria, 87(1): 207–229, February 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Poller:2021:VRN

[1649] Olga Poller. Variability, rigidity and the nesting problem. Theoria, 87 (1):230–246, February 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Stjernberg:2021:BRM

[1650] Fredrik Stjernberg. Book review: Misak, Cheryl, Frank Ramsey: A Sheer Excess of Powers. Oxford, Oxford University Press, 2020. 544 pp. Hardback, £25.00. Theoria, 87(1):247–251, February 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2021:IIb

[1651] Anonymous. Issue information. Theoria, 87(2):253–254, April 2021. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2021:PP

[1652] Sven Ove Hansson. Plagiarize or perish? Theoria, 87(2):255–258, April 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Shelah:2021:DCD

[1653] Saharon Shelah. Divide and conquer: Dividing lines and universality. Theoria, 87(2):259–348, April 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Vaananen:2021:OSS

[1654] Jouko Väänänen. An overview of Saharon Shelah’s contributions to mathematical logic, in particular to model theory. Theoria, 87(2):349– 360, April 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Baldwin:2021:DLM

[1655] John T. Baldwin. The dividing line methodology: Model theory motivating set theory. Theoria, 87(2):361–393, April 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Istvan:2021:RDD

[1656] Michael A. Istvan, Jr. A rationalist defence of determinism. Theoria, 87(2):394–434, April 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Skrzypulec:2021:TUV

[1657] Błazej Skrzypulec. Tropes, universals and visual phenomenology. Theoria, 87(2):435–456, April 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Azzano:2021:IUT

[1658] Lorenzo Azzano and Massimiliano Carrara. Are identities unexplainable? Towards a non-causal contrastive explanation of identities. Theoria, 87 (2):457–482, April 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ishida:2021:WMD

[1659] Shu Ishida. What makes discrimination morally wrong? A harmbased view reconsidered. Theoria, 87(2):483–499, April 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2021:IIc

[1660] Anonymous. Issue information. Theoria, 87(3):501–502, June 2021. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2021:HEC

[1661] Sven Ove Hansson. How exact can philosophy be? Theoria, 87(3):503– 505, June 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Mosdell:2021:WRL

[1662] Matthew Mosdell. Work’s role in learning how. Theoria, 87(3):506– 519, June 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Daoust:2021:FBM

[1663] Marc-Kevin Daoust. False beliefs and misleading evidence. Theoria, 87 (3):520–541, June 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Altehenger:2021:DAS

[1664] Hannah Altehenger. Deliberative agency, self-control, and the divided mind. Theoria, 87(3):542–558, June 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Klausen:2021:WTK

[1665] Søren Harnow Klausen and Esben Nedenskov Petersen. What is theoretical knowledge? Theoria, 87(3):559–577, June 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Giannotti:2021:FYG

[1666] Joaquim Giannotti. Fundamental yet grounded. Theoria, 87(3):578– 599, June 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Parvizian:2021:LIP

[1667] Saja Parvizian. Leibniz on intellectual pleasure, perception of perfection, and power. Theoria, 87(3):600–627, June 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bonardi:2021:CCP

[1668] Paolo Bonardi. The “Cicero”/“Cicero” puzzling case. Theoria, 87(3): 628–642, June 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Volacu:2021:JSV

[1669] Alexandru Volacu. Justice, symmetry, and voting rights ceilings. Theoria, 87(3):643–658, June 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Freschi:2021:DCT

[1670] Elisa Freschi and Matteo Pascucci. Deontic concepts and their clash in Mimam .sa: Towards an interpretation. Theoria, 87(3):659–703, June 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

vanRooij:2021:WBR

[1671] Robert van Rooij and Katrin Schulz. Why those biscuits are relevant and on the sideboard. Theoria, 87(3):704–712, June 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Rasanen:2021:SAP

[1672] Joona Räsänen and Pekka Louhiala. Should acknowledgments in published academic articles include gratitude for reviewers who reviewed for journals that rejected those articles? Theoria, 87(3):713–728, June 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Beni:2021:ACE

[1673] Majid D. Beni. Agency, causation, and empiricism. Theoria, 87(3):729– 751, June 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Lanius:2021:WVV

[1674] David Lanius. What is the value of vagueness? Theoria, 87(3):752– 780, June 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Laasik:2021:CC

[1675] Kristjan Laasik. Constancy and constitution. Theoria, 87(3):781–798, June 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Koscholke:2021:NAG

[1676] Jakob Koscholke. A new argument for Goldman and Olsson’s solution to the extra-value-of-knowledge problem. Theoria, 87(3):799–812, June 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Paoletti:2021:MIF

[1677] Michele Paolini Paoletti. Masks, interferers, finks, and mimickers: a novel approach. Theoria, 87(3):813–836, June 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Pietroiusti:2021:DCH

[1678] Giulio Pietroiusti. Disagreement and conflict: How moral and taste judgements do Not differ. Theoria, 87(3):837–846, June 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Vignolo:2021:LTA

[1679] Massimiliano Vignolo. The lying test, ambiguity, and determination of content: Lying and determination of content. Theoria, 87(3):847–857, June 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2021:IId

[1680] Anonymous. Issue information. Theoria, 87(4):859–860, August 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2021:ERG

[1681] Sven Ove Hansson. Ethical reasoning: Guidance or just rationalization? Theoria, 87(4):861–865, August 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2021:STP

[1682] Sven Ove Hansson. Swedish theses in philosophy 2020. Theoria, 87 (4):866–869, August 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Cortes:2021:MTE

[1683] María Clara Cortés, Juliette Kennedy, and Andrés Villaveces. Mapping traces: Editorial introduction. Theoria, 87(4):870–873, August 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kennedy:2021:GIS

[1684] Juliette Kennedy. Gödel and the integrated self, or: On the philosopher’s second sailing. Theoria, 87(4):874–884, August 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Zalamea:2021:MAL

[1685] Fernando Zalamea. Mathematical, artistic and literary traces in reason/ heart (razón/corazón) transitions. Theoria, 87(4):885–896, August 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Zwicky:2021:SDT

[1686] Jan Zwicky. Show, don’t tell. Theoria, 87(4):897–912, August 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Baldwin:2021:IM

[1687] John T. Baldwin. Images in mathematics. Theoria, 87(4):913–936, August 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Malagon-kurka:2021:ISA

[1688] María Margarita Malagón-kurka, Clemencia Echeverri, and Beatriz Eugenia Vallejo Franco. Indexical signs and artistic, political and historical complexity. Theoria, 87(4):937–958, August 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kossak:2021:LSA

[1689] Roman Kossak and Wanda Siedlecka. Logic and structure: an art project. Theoria, 87(4):959–970, August 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Morales:2021:ALP

[1690] John A. Cruz Morales, Andrés Villaveces, and Boris Zilber. Around logical perfection. Theoria, 87(4):971–985, August 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Vaananen:2021:TIC

[1691] Jouko Väänänen. Tracing internal categoricity. Theoria, 87(4):986–1000, August 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Villaveces:2021:CEI

[1692] Andrés Villaveces. Categoricity, external and internal: an excerpt from a conversation with Saharon Shelah. Theoria, 87(4):1001–1012, August 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2021:IIe

[1693] Anonymous. Issue information. Theoria, 87(5):1013–1014, October 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2021:PO

[1694] Sven Ove Hansson. Philosophical obscurity. Theoria, 87(5):1015–1019, October 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Lee:2021:CPL

[1695] Jeonggyu Lee. The contingent A Priori, linguistic stipulation, and singular thought. Theoria, 87(5):1020–1037, October 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tang:2021:WSC

[1696] Zhiheng Tang. We should not be a counterpart theorist of events if we want to be a counterfactual theorist of causation. Theoria, 87(5): 1038–1049, October 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Zinke:2021:RS

[1697] Alexandra Zinke. Rational suspension. Theoria, 87(5):1050–1066, October 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Fatoorchi:2021:SSI

[1698] Pirooz Fatoorchi. Soul-switching and the immateriality of human nature: On an argument reported by Razi. Theoria, 87(5):1067–1082, October 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sanni:2021:UOR

[1699] John Sodiq Sanni. Ubuntu and the ontology of radical escape. Theoria, 87(5):1083–1098, October 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Deonna:2021:WAF

[1700] Julien A. Deonna and Fabrice Teroni. Which attitudes for the fitting attitude analysis of value? Theoria, 87(5):1099–1122, October 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bricker:2021:CEV

[1701] Adam Michael Bricker. Close error, visual perception, and neural phase: a critique of the modal approach to knowledge. Theoria, 87(5):1123–1152, October 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Egan:2021:SIL

[1702] Magnus Skytterholm Egan. Structural injustice and labour migration — from individual responsibility to collective action. Theoria, 87(5): 1153–1174, October 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Francez:2021:LGC

[1703] Nissim Francez. Logical grounding: The case of “if-then-else”. Theoria, 87(5):1175–1192, October 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Raatikainen:2021:CS

[1704] Panu Raatikainen. Chalmers and semantics. Theoria, 87(5):1193–1221, October 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Volpe:2021:CHE

[1705] Giorgio Volpe. Contrastive hinge epistemology. Theoria, 87(5):1222– 1249, October 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Lee:2021:PDR

[1706] Kok Yong Lee. On the problem of deviant realizations. Theoria, 87(5): 1250–1269, October 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Paasch:2021:DSI

[1707] Jt Paasch. Duns Scotus on identities — I mean, mereological fusions. Theoria, 87(5):1270–1306, October 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Fratantonio:2021:NII

[1708] Giada Fratantonio. No infelicity for the infallibilist. Theoria, 87(5): 1307–1321, October 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hincu:2021:WAA

[1709] Mihai Hîncu and Dan Zeman. On Wyatt’s absolutist account of faultless disagreement in matters of personal taste. Theoria, 87(5):1322–1341, October 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2021:IIf

[1710] Anonymous. Issue information. Theoria, 87(6):1343–1344, December 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2021:H

[1711] Sven Ove Hansson. Holism. Theoria, 87(6):1345–1348, December 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Norton:2021:EIQ

[1712] Joshua Norton. Empirically incoherent quantum gravity. Theoria, 87(6): 1349–1379, December 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tevdoradze:2021:CCL

[1713] Nino Tevdoradze. The complexity of the concept of literary autonomy. Theoria, 87(6):1380–1396, December 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kim:2021:FCL

[1714] Junyeol Kim. Frege’s conception of logic: Truth, the true, and assertion. Theoria, 87(6):1397–1417, December 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tuzet:2021:PEW

[1715] Giovanni Tuzet. The policy of evidence: When epistemic and practical interests conflict, the case of FRE 407. Theoria, 87(6):1418–1443, December 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Campdelacreu:2021:CAG

[1716] Marta Campdelacreu. Colocationist answers to the grounding problem. Theoria, 87(6):1444–1467, December 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Noller:2021:LIK

[1717] Jörg Noller. The logic of illusion: Kant on the reasons of error. Theoria, 87(6):1468–1480, December 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Smilansky:2021:CRM

[1718] Saul Smilansky. Contribution, replaceability and the meaning of our lives. Theoria, 87(6):1481–1496, December 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bien:2021:HOD

[1719] Eric Nenkia Bien. How obscurantism differs from bullshit: a proposal. Theoria, 87(6):1497–1526, December 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sander:2021:MEI

[1720] Thorsten Sander. The myth of epistemic implicata. Theoria, 87(6): 1527–1547, December 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Perrine:2021:CMP

[1721] Timothy Perrine. Cleaning up, and moving past, simple swamping. Theoria, 87(6):1548–1561, December 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Verdejo:2021:TD

[1722] Víctor M. Verdejo and Xavier de Donato-Rodríguez. Thinking disagreement. Theoria, 87(6):1562–1584, December 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Nefdt:2021:SMP

[1723] Ryan M. Nefdt. Scientific models and political theory: The ideal theory debate revisited. Theoria, 87(6):1585–1608, December 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Zaar:2021:WCK

[1724] Balder Edmund Ask Zaar. Has Williamson’s claim that knowledge is the most general factive mental state been disproved? Theoria, 87(6): 1609–1634, December 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hausmann:2021:UNT

[1725] Marco Hausmann. The uniqueness of necessary truth and the status of S4 and S5. Theoria, 87(6):1635–1650, December 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Zimmermann:2021:PTS

[1726] Ernst Zimmermann. Proof-theoretic semantics of natural deduction based on inversion. Theoria, 87(6):1651–1670, December 2021. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2022:IIa

[1727] Anonymous. Issue information. Theoria, 88(1):1–2, February 2022. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2022:PF

[1728] Sven Ove Hansson. Has philosophy failed? Theoria, 88(1):3–7, February 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Lagerlund:2022:MS

[1729] Henrik Lagerlund. Medieval scepticism. Theoria, 88(1):8–12, February 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Mihirig:2022:TSP

[1730] Abdurrahman Ali Mihirig. Typologies of scepticism in the philosophical tradition of Kalam. Theoria, 88(1):13–48, February 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Eichner:2022:SFI

[1731] Heidrun Eichner. Scepticisms in the formation of Islamic rational theology: Abu al-Qasim al-Balkhi and Ibn al-Malah.imi providing a window on the transmission of arguments from Late Antiquity. Theoria, 88(1): 49–71, February 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Benevich:2022:MDS

[1732] Fedor Benevich. Meaning and definition: Scepticism and semantics in Twelfth-Century Arabic philosophy. Theoria, 88(1):72–108, February 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hadisi:2022:GTA

[1733] Reza Hadisi. Ghazali’s transformative answer to scepticism. Theoria, 88 (1):109–142, February 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Halper:2022:DMS

[1734] Yehuda Halper. Dialectic and metaphysical skepticism in Jacob Anatoli. Theoria, 88(1):143–164, February 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Greig:2022:RRS

[1735] Jonathan Greig. Reason, revelation, and sceptical argumentation in 12thto 14th-Century Byzantium. Theoria, 88(1):165–201, February 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Dasdemir:2022:ASI

[1736] Yusuf Dasdemir. Akbarian scepticism in Islam: Qunawi’s sceptical arguments from relativity and disagreement. Theoria, 88(1):202–225, February 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Schmitt:2022:SSE

[1737] Jens Ole Schmitt. On some sceptical elements in Barhebraeus. Theoria, 88(1):226–243, February 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Perler:2022:CWK

[1738] Dominik Perler. Can we know substances? Suárez on a sceptical puzzle. Theoria, 88(1):244–269, February 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2022:IIb

[1739] Anonymous. Issue information. Theoria, 88(2):271–272, April 2022. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2022:DDS

[1740] Sven Ove Hansson. Defining disciplines and subdisciplines. Theoria, 88 (2):273–275, April 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Almog:2022:IVN

[1741] Joseph Almog. Introduction to the volume “Naming and Necessity: A 40th-Year Anniversary”. Theoria, 88(2):276–277, April 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bianchi:2022:BGA

[1742] Andrea Bianchi. Back to the Golden Age: Saul Kripke’s Naming and Necessity and Twenty-First Century Philosophy. Theoria, 88(2):278– 295, April 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kamp:2022:LCC

[1743] Hans Kamp. The links of causal chains. Theoria, 88(2):296–325, April 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sinokki:2022:WES

[1744] Jani Sinokki. “What on Earth is Smenkhkare?” WH-questions, truthmakers, and causal-informational account of reference. Theoria, 88(2): 326–347, April 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Pepp:2022:TSB

[1745] Jessica Pepp. Towards a sensible bifurcationism (concerning what grounds thought about particulars). Theoria, 88(2):348–364, April 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Almog:2022:FYL

[1746] Joseph Almog. Forty years later: Naming without necessity, necessity without naming. Theoria, 88(2):365–402, April 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Fine:2022:SRR

[1747] Kit Fine. Some remarks on the role of essence in Kripke’s “Naming and Necessity”. Theoria, 88(2):403–405, April 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Devitt:2022:RNN

[1748] Michael Devitt. Reflections on Naming and Necessity. Theoria, 88(2): 406–433, April 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Salmon:2022:LNN

[1749] Nathan Salmón. The legacy of naming and necessity. Theoria, 88(2):434– 437, April 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Stalnaker:2022:RFC

[1750] Robert Stalnaker. Reference fixing and the contingent a priori. Theoria, 88(2):438–452, April 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Schwartz:2022:BP

[1751] Stephen P. Schwartz. “A better picture … ”. Theoria, 88(2):453–463, April 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Korte:2022:FAK

[1752] Tapio Korte. Frege’s answer to Kripke. Theoria, 88(2):464–479, April 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2022:IIc

[1753] Anonymous. Issue information. Theoria, 88(3):481–482, June 2022. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2022:LCF

[1754] Anonymous. Learning from Condorcet on formal philosophy. Theoria, 88(3):483–486, June 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Rabinowicz:2022:SLM

[1755] Wlodek Rabinowicz. Sten Lindström in memoriam. Theoria, 88(3):487– 490, June 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Rott:2022:ISL

[1756] Hans Rott. Introduction to Sten Lindström’s “A semantic approach to nonmonotonic reasoning: Inference operations and choice”. Theoria, 88 (3):491–493, June 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Lindstrom:2022:SAN

[1757] Sten Lindström. A semantic approach to nonmonotonic reasoning: Inference operations and choice. Theoria, 88(3):494–528, June 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2022:STP

[1758] Sven Ove Hansson. Swedish theses in philosophy 2021. Theoria, 88(3): 529–531, June 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Iaquinto:2022:PNS

[1759] Samuele Iaquinto and Giuseppe Spolaore. A philosophically neutral semantics for perception sentences. Theoria, 88(3):532–544, June 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anaya:2022:LMD

[1760] Alfonso Anaya. “Let me double-check that”: a challenge for conciliationism. Theoria, 88(3):545–557, June 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

PhD:2022:UAP

[1761] Felipe Morales Carbonell PhD. Understanding attribution: Problems, options, and a proposal. Theoria, 88(3):558–583, June 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Taylor:2022:FOL

[1762] R. Gregory Taylor. First-order logics over fixed domain. Theoria, 88 (3):584–606, June 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Fernandez:2022:SEJ

[1763] Ricardo Mejía Fernández and Javier Romero. Social evolution in Jürgen Habermas: Towards a weak anthropological naturalism between Kant and Darwin. Theoria, 88(3):607–628, June 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Giovannini:2022:MGB

[1764] Eduardo N. Giovannini and Abel Lassalle-Casanave. From magnitudes to geometry and back: De Zolt’s postulate. Theoria, 88(3):629–652, June 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Mosquera:2022:DDR

[1765] Julia Mosquera. Defining disability and the role of the disability and the medical communities. Theoria, 88(3):653–665, June 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Jecker:2022:AER

[1766] Nancy S. Jecker. African ethics, respect for persons, and moral dissent. Theoria, 88(3):666–678, June 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Berghofer:2022:HSM

[1767] Philipp Berghofer. How to supplement mentalist evidentialism: What are the fundamental epistemological principles? Theoria, 88(3):679– 700, June 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hirvonen:2022:DD

[1768] Ilmari Hirvonen and Janne Karisto. Demarcation without dogmas. Theoria, 88(3):701–720, June 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2022:IId

[1769] Anonymous. Issue information. Theoria, 88(4):721–722, August 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2022:PSA

[1770] Sven Ove Hansson. Protect science against frivolous litigation! Theoria, 88(4):723–726, August 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Norton:2022:I

[1771] Joshua Norton. Introduction. Theoria, 88(4):727–732, August 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Buonocore:2022:SIR

[1772] Giovanni Buonocore. A structuralist interpretation of the relational interpretation. Theoria, 88(4):733–742, August 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hemmo:2022:FP

[1773] Meir Hemmo and Orly Shenker. Flat physicalism. Theoria, 88(4):743– 764, August 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Shirtz:2022:PB

[1774] Aviv Spector Shirtz. Pain begins. Theoria, 88(4):765–781, August 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Diez:2022:GPS

[1775] José A. Díez and C. Ulises Moulines. Guiding principles and special laws. Theoria, 88(4):782–798, August 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Haji:2022:SI

[1776] Ishtiyaque Haji. Semicompatibilism imperiled. Theoria, 88(4):799–811, August 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Aliyev:2022:PCS

[1777] Alexey Aliyev. Paraphrasability and the causal status of types. Theoria, 88(4):812–828, August 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Ceragioli:2022:NPT

[1778] Leonardo Ceragioli. New problems for Tennant’s definition of harmony. Theoria, 88(4):829–849, August 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Goldman:2022:RAM

[1779] Alan H. Goldman. The right argument from moral disagreement. Theoria, 88(4):850–867, August 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sauchelli:2022:CWP

[1780] Andrea Sauchelli. Centred worlds, personal identity and imagination. Theoria, 88(4):868–880, August 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2022:IIe

[1781] Anonymous. Issue information. Theoria, 88(5):881–882, October 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2022:PC

[1782] Sven Ove Hansson. Philosophical conversations. Theoria, 88(5):883–886, October 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

vanRiel:2022:RP

[1783] Raphael van Riel. Representationalist physicalism. Theoria, 88(5):887– 898, October 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Honkasalo:2022:SCD

[1784] Aleksi Honkasalo. Is semantic correctness descriptive? Theoria, 88(5): 899–907, October 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Reshenin:2022:HES

[1785] Semyon Reshenin. Harmony without equality: Schiller’s theory of virtue. Theoria, 88(5):908–926, October 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kalpokas:2022:NCM

[1786] Daniel E. Kalpokas. A new case of the Myth of the Given? Theoria, 88 (5):927–942, October 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Shech:2022:SUA

[1787] Elay Shech. Scientific understanding in the Aharonov–Bohm effect. Theoria, 88(5):943–971, October 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Lau:2022:AA

[1788] Tony Tsz Fung Lau. An analysis of absurdity. Theoria, 88(5):972–981, October 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Belleri:2022:ODR

[1789] Delia Belleri. Ontological disagreements, reference, and charity: a challenge for Hirsch’s deflationism. Theoria, 88(5):982–996, October 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Schipper:2022:NCP

[1790] Arthur Schipper. Necessitation and the changing past. Theoria, 88(5): 997–1013, October 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

dAragona:2022:NEP

[1791] Antonio Piccolomini d’Aragona. A note on Etchemendy’s and Prawitz’s reduction principles for the Tarskian and model-theoretic concept of consequence. Theoria, 88(5):1014–1036, October 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Valtonen:2022:GIR

[1792] Pasi Valtonen. Generic inferential rules for slurs and contrasting senses. Theoria, 88(5):1037–1052, October 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Bennet:2022:NTU

[1793] Christian Bennet and Rasmus Blanck. Never trust an unsound theory. Theoria, 88(5):1053–1056, October 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2022:IIf

[1794] Anonymous. Issue information. Theoria, 88(6):1057–1058, December 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2022:EOU

[1795] Sven Ove Hansson. Equal opportunity at the university. Theoria, 88(6): 1059–1062, December 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Marceta:2022:PAP

[1796] Jesper Ahlin Marceta and Niklas Juth. Personal autonomy: From practice to theory. Theoria, 88(6):1063–1065, December 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Gidron:2022:PPA

[1797] Tamar Gidron and Elad Schild. Protection of patient autonomy via consumer protection litigation: The Israeli Eltroxin Class Action as a case study. Theoria, 88(6):1066–1085, December 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Rueda:2022:SDO

[1798] Jon Rueda. From self-determination to offspring-determination? reproductive autonomy, procrustean parenting, and genetic enhancement. Theoria, 88(6):1086–1110, December 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

deAlmeida:2022:MPG

[1799] Claudio de Almeida. Moore’s paradox generalized. Theoria, 88(6):1111– 1127, December 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sun:2022:RDN

[1800] Yifan Sun. Response-dependence and normativity. Theoria, 88(6):1128– 1143, December 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Intropi:2022:FVG

[1801] Pietro Intropi. Freedom’s values: The good and the right. Theoria, 88(6): 1144–1162, December 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

De:2022:HCT

[1802] Michael De. Haecceitism and counterpart theory. Theoria, 88(6):1163– 1179, December 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Andersson:2022:DJS

[1803] Emil Andersson. Distributive justice, social cooperation, and the basis of equality. Theoria, 88(6):1180–1195, December 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Fives:2022:MOC

[1804] Allyn Fives. Membership, obligation, and the communitarian thesis. Theoria, 88(6):1196–1210, December 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Arenhart:2022:LAE

[1805] Jonas R. Becker Arenhart. Logical anti-exceptionalism meets the “logicas-models” approach. Theoria, 88(6):1211–1227, December 2022. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

vanAken:2022:GGC

[1806] Erik van Aken. Getting a grip: On causation, agency, and the meaning of “manipulation”. Theoria, 88(6):1228–1247, December 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Mizumoto:2022:PEC

[1807] Masaharu Mizumoto. A prolegomenon to the empirical cross-linguistic study of truth. Theoria, 88(6):1248–1273, December 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Orilia:2022:NBP

[1808] Francesco Orilia. A new bridge principle for the normativity of logic. Theoria, 88(6):1274–1292, December 2022. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic). See [1827].

Anonymous:2023:IIa

[1809] Anonymous. Issue information. Theoria, 89(1):1–2, February 2023. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2023:PO

[1810] Sven Ove Hansson. Philosophical originality. Theoria, 89(1):3–6, February 2023. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Stigall:2023:ETC

[1811] John Stigall. Not every truth could have a truthmaker. Theoria, 89 (1):7–13, February 2023. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tilley:2023:DEE

[1812] John J. Tilley. On deducing ethical egoism from psychological egoism. Theoria, 89(1):14–30, February 2023. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

DeFlorio:2023:CMA

[1813] Ciro De Florio and Aldo Frigerio. A critique of Malpass’s argument against Supervaluationism. Theoria, 89(1):31–41, February 2023. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hamdo:2023:GSC

[1814] Manhal Hamdo. On Galen Strawson’s central approach to the self. Theoria, 89(1):42–56, February 2023. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Capuano:2023:NA

[1815] Antonio Capuano. Naming and analyticity. Theoria, 89(1):57–72, February 2023. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Tana:2023:CUP

[1816] Guido Tana. Closure, underdetermination and the peculiarity of Sceptical scenarios. Theoria, 89(1):73–97, February 2023. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Akiba:2023:APT

[1817] Takeshi Akiba. Alethic pluralism and truthmaker theory. Theoria, 89 (1):98–113, February 2023. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Miller:2023:MVD

[1818] J. T. M. Miller. Merely verbal disputes and common ground. Theoria, 89 (1):114–123, February 2023. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Humberstone:2023:TVR

[1819] Lloyd Humberstone. Truth-value relations and logical relations. Theoria, 89(1):124–147, February 2023. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2023:IIb

[1820] Anonymous. Issue information. Theoria, 89(2):149–150, April 2023. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2023:HDP

[1821] Sven Ove Hansson. How not to deal with plagiarism. Theoria, 89(2):151– 155, April 2023. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sudarshan:2023:RDM

[1822] Saranga Sudarshan. Reasonable disagreement and metalinguistic negotiation. Theoria, 89(2):156–175, April 2023. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Folland:2023:FNI

[1823] Anna Folland. Feit on the normative importance of harm. Theoria, 89 (2):176–187, April 2023. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Beni:2023:CIB

[1824] Majid D. Beni. Casting inference to the best explanation’s lot with active inference. Theoria, 89(2):188–203, April 2023. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Pawlowski:2023:IPD

[1825] Pawel Pawlowski and Rafal Urbaniak. Informal provability and dialetheism. Theoria, 89(2):204–215, April 2023. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Jagnow:2023:PPE

[1826] René Jagnow. The particularity of photographic experience. Theoria, 89(2):216–231, April 2023. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2023:CNB

[1827] Anonymous. Correction to ‘A new bridge principle for the normativity of logic’. Theoria, 89(2):232, April 2023. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic). See [1808].

Anonymous:2023:IIc

[1828] Anonymous. Issue information. Theoria, 89(3):233–234, June 2023. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2023:PPR

[1829] Sven Ove Hansson. Philosophy without peer review? Theoria, 89(3): 235–238, June 2023. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Carrara:2023:TOO

[1830] Massimiliano Carrara and Luca Zanetti. Thin objects: an overview. Theoria, 89(3):239–246, June 2023. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Linnebo:2023:P

[1831] Øystein Linnebo. Précis. Theoria, 89(3):247–255, June 2023. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Carrara:2023:TOT

[1832] Massimiliano Carrara, Ciro De Florio, and Francesca Poggiolesi. From thin objects to thin concepts? Theoria, 89(3):256–265, June 2023. CO-DEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Donaldson:2023:MPN

[1833] Thomas Donaldson. A metaphysical puzzle for neo-fregean abstractionists. Theoria, 89(3):266–279, June 2023. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Sereni:2023:MTA

[1834] Andrea Sereni and Luca Zanetti. Minimalism, trivialism, aristotelianism. Theoria, 89(3):280–297, June 2023. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Pantsar:2023:WGD

[1835] Markus Pantsar. On what ground do thin objects exist? in search of the cognitive foundation of number concepts. Theoria, 89(3):298–313, June 2023. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Plebani:2023:TOT

[1836] Matteo Plebani, Luca San Mauro, and Giorgio Venturi. Thin objects are not transparent. Theoria, 89(3):314–325, June 2023. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Pearce:2023:NAR

[1837] Gareth Rhys Pearce. A nominalist alternative to reference by abstraction. Theoria, 89(3):326–337, June 2023. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Lando:2023:TMS

[1838] Giorgio Lando. Thin mereological sums, abstraction, and interpretational modalities. Theoria, 89(3):338–355, June 2023. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Eklund:2023:TE

[1839] Matti Eklund. Thin entities. Theoria, 89(3):356–365, June 2023. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Studd:2023:LAB

[1840] J. P. Studd. Linnebo’s abstractionism and the bad company problem. Theoria, 89(3):366–392, June 2023. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Linnebo:2023:R

[1841] Øystein Linnebo. Replies. Theoria, 89(3):393–406, June 2023. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Anonymous:2023:IId

[1842] Anonymous. Issue information. Theoria, 89(4):407–408, August 2023. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Hansson:2023:SWP

[1843] Sven Ove Hansson. Should we be politically correct? Theoria, 89(4): 409–413, August 2023. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Omori:2023:CLC

[1844] Hitoshi Omori and Jonas R. B. Arenhart. Change of logic, without change of meaning. Theoria, 89(4):414–431, August 2023. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Schwengerer:2023:AJA

[1845] Lukas Schwengerer. An anchored joint acceptance account of group justification. Theoria, 89(4):432–450, August 2023. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kallestrup:2023:MTL

[1846] Jesper Kallestrup. The myth of true lies. Theoria, 89(4):451–466, August 2023. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Gustafsson:2023:DPM

[1847] Johan E. Gustafsson. Dennett’s prime-mammal objection to the consequence argument. Theoria, 89(4):467–470, August 2023. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Kaluzinski:2023:RPA

[1848] Bartosz Kaluziński. Rules, practices, and assessment of linguistic behaviour. Theoria, 89(4):471–482, August 2023. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Moore:2023:WBD

[1849] Andrew Garford Moore and George Botterill. Why beliefs are not dispositional stereotypes. Theoria, 89(4):483–494, August 2023. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Muhlebach:2023:MDS

[1850] Deborah Mühlebach. Meaning in derogatory social practices. Theoria, 89 (4):495–515, August 2023. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Grindrod:2023:DCA

[1851] Jumbly Grindrod. A defence of conceptual analysis as a linguistic endeavour. Theoria, 89(4):516–534, August 2023. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Lemanek:2023:ISD

[1852] Kamil Lemanek. Inferentialism and social delusion. Theoria, 89(4):535– 547, August 2023. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).

Rosenthal:2023:CAO

[1853] Jacob Rosenthal. Conditional analyses of options for action: a partial defence. Theoria, 89(4):548–563, August 2023. CODEN THRAA5. ISSN 0040-5825 (print), 1755-2567 (electronic).